### v2

# Forecasting from Clinical Textual Time Series: Adaptations of the Encoder and Decoder Language Model Families

Abstract
Clinical case reports encode rich, temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data. In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings–extracted via an LLM-assisted annotation pipeline–serve as the primary input for prediction. We systematically evaluate a diverse suite of models, including fine-tuned decoder-based large language models and encoder-based transformers, on tasks of event occurrence prediction, temporal ordering, and survival analysis. Our experiments reveal that encoder-based models consistently achieve higher F1 scores and superior temporal concordance for short- and long-horizon event forecasting, while fine-tuned masking approaches enhance ranking performance. In contrast, instruction-tuned decoder models demonstrate a relative advantage in survival analysis, especially in early prognosis settings. Our sensitivity analyses further demonstrate the importance of time ordering, which requires clinical time series construction, as compared to text ordering, the format of the text inputs that LLMs are classically trained on. This highlights the additional benefit that can be ascertained from time-ordered corpora, with implications for temporal tasks in the era of widespread LLM use.
**footnotetext: These authors contributed equally to this work
1 Introduction
Clinical risk forecasting is a cornerstone problem in health care decision making that guides eligibility and reasoning about interventions. Because risk prediction performs best with access to rich data streams, great lengths are taken to capture the relevant health factors that affect decision making. However, while many forecasting tasks are performed on structured tabular data, evidence suggests that, despite widespread efforts to collect large amount of data in structured format, only a small fraction of structured data overlaps with that documented in text (Seinen et al., 2025). Furthermore, empirically we observe that predictions are considerably better when text is integrated (Kline et al., 2022).
A complementary data stream in healthcare comes from the textual data that providers use in documenting their encounters. Likely the most comprehensive part are the case reports and discharge summaries, which are a sort of “post-mortem”, retrospective discussion of the relevant factors for decision making and observed progression. However, without further effort, we are unable to leverage this data stream for forecasting because the times that are documented are interwoven in the narrative passage.
To construct and learn from these patient trajectories, this work investigates the development of a forecasting pipeline for textual time series from retrospective analysis of clinical trajectories from case reports. Using the framework in Wang and Weiss (2025), we leverage extracted temporally located clinical findings, namely lists of (event, time) tuples that represent the patient encounter. The mapping from case report to time-localized events allows us to make anytime forecasts by masking those events that occur after the requested forecast time. Without this annotation process, it would be unclear which findings could be used so as to avoid “causal leakage” from events occurring in the future.
Our approach differs from other efforts in the NLP and clinical informatics community and is driven by our particular emphasis on high-performance forecasting and temporally-aware risk characterization. For example, our definition of clinical finding is related to clinical entity and clinical concept definitions from the field of named entity recognition. However, we avoid mapping to a discrete set of entities to increase the availability of contextual information for forecasting, and thus require an alternative annotation guideline (Uzuner et al., 2011). With respect to our temporal annotations, our approach differs from the classical temporal relations task, which, while capturing relative ordering and overlap of events, makes it difficult to perform risk calculations and multimodal alignments that are functions of time (Kohane, 1987; Zhou and Hripcsak, 2007; Sun et al., 2013).
Figure 1: Example case report (left) with temporal representations as timelines (middle) and text-ordered, event-time tuples (right).
We use LLMs to assist us in this task, so it is natural to question whether we need this extraction/annotation step at all, since a multitude of BERT-based and LLM models have consumed these, similar post-mortem reports, and more during training (Peng et al., 2020; Weber et al., 2024). From an architectural perspective, both encoder (BERT-style) and decoder (LLM-style) models will not perform the temporal masking necessary to encode leakage-free embeddings. For BERT models, which use full attention modules with masking, the probability that the random masking recovers a time-ordered representation is vanishingly small (tested empirically on the example timeline in Figure 1 (right) with default BERT masking probability of 0.15; probability of a time-ordered masking: 2/100,000). For decoder models, their causal attention mask means that embeddings are processed in text-order rather than time-order. Given that the concordance, the fraction of the time pairwise temporal ordering are correct, between text-ordered and time-ordered events is lower than 1 in most reports (c=0.23 in the example in Figure 1), we would expect suboptimal temporal reasoning capabilities.
As we will see, we empirically find that these models do not forecast well without further training. If we can address these textual-temporal discrepancies, we expect to be able to substantially improve LLM capabilities at clinical forecasting, which could lead to downstream improvements in clinical temporal reasoning, risk estimation, causal reasoning, and decision making.
1.1 Contributions
  1. 1.
    Annotation and Extraction Pipeline: Building on methods of Noroozizadeh and Weiss (2025) for extracting information from the PubMed Open Access corpus, we develop a pipeline that converts case reports into textual time series using both regex-based filtering and LLM-assisted extraction. The pipeline produces temporally localized (event, time) tuples, mitigating potential causal leakage by ensuring that only pre-cutoff events inform forecasts.
  2. 2.
    Comprehensive Model Comparison: We rigorously evaluate a broad suite of models across multiple forecasting tasks. Our study compares instruction-tuned decoder-based models (both via fine-tuning with MLP heads and prompting strategies) against encoder-based models with task-specific fine-tuning. Experiments on event occurrence prediction, temporal ordering, and survival analysis reveal complementary strengths: no single modeling approach consistently outperforms across all tasks, highlighting the importance of aligning model choice with the specific forecasting objective.
  3. 3.
    Analysis of Temporal versus Text Ordering: We perform a detailed investigation into the impact of ordering on performance by contrasting training on time-ordered versus text-ordered annotations. Our experiments demonstrate that preserving the clinical narrative’s natural order can enhance generalization on external datasets, while explicit time ordering can improve ranking performance under specific conditions.
  4. 4.
    Sensitivity Analysis of Temporal Masking: Through systematic dropout experiments, we analyze the effects of randomly masking portions of the clinical history on forecasting accuracy and event ordering. The findings reveal that while classification performance (F1 score) deteriorates with higher levels of masking, the concordance index remains largely stable—highlighting differing sensitivities of binary prediction and ranking tasks to historical context in our textual time series setup.
Generalizable Insights about Machine Learning in the Context of Healthcare
Extensive research has been in the clinical NLP community towards temporal understanding of clinical narratives (Kohane, 1987; Leeuwenberg and Moens, 2020; Cheng and Weiss, 2023). While much has focused on temporal relations, relative-time event tuples facilitate the use of longitudinal and survival models so long as they can use free-text inputs and outcomes (possibly via mappings to an ontology). Whereas the primary focus in this space is on the entity and temporal extraction tasks, downstream application studies remain infrequent. Our study seeks to close thie gap by applying forecasting methodologies to these data streams to characterize the performance of leading machine learning and NLP approaches. A better empirical understanding of the strength and weakness of these approaches will enable future work to advance improved textual time series methods. Our investigation also highlights that risk estimation and event anticipation is not optimally captured in LLMs (as we demonstrate in contrasts between fine-tuning and prompt-based strategies), suggesting the need for careful investigation of LLMs when applied to assessment of risk.
2 Related Work
Temporal Information Extraction from Clinical Text:
Extracting timelines from clinical narratives has been a longstanding challenge in biomedical NLP. Early efforts such as the i2b2 2012 challenge provided annotated datasets for temporal relation extraction, focusing on identifying and ordering events in discharge summaries (Uzuner et al., 2011). Follow-up work constructed timelines from these narratives by linking clinical events to timestamps or temporal expressions (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024). These methods often relied on pre-defined event spans and excerpts, rather than full-document processing. In contrast, following Noroozizadeh and Weiss (2025), we assign timepoints to clinical findings from full-length case reports, allowing for finer temporal resolution. While some works integrate structured EHR data to anchor timelines, our approach focuses on timeline construction and forecasting using text alone, which is essential for corpora such as the PubMed Open-Access Subset, where no structured metadata is available. Our annotation strategy enables direct supervision of event-time alignment, overcoming the limitations of earlier span-only annotations. This aligns with the broader need in sepsis and critical care research to incorporate temporality into clinical phenotyping (Johnson et al., 2018; Kamran et al., 2024; Seinen et al., 2025), especially given the high missingness in structured data (Johnson et al., 2016, 2023).
Predictive Modeling with Clinical Text:
Incorporating unstructured text (clinical notes, narratives) into predictive models has been explored in various contexts, primarily for outcomes like mortality or readmission. Notably, ClinicalBERT and related transformer models pre-trained on medical notes have achieved state-of-the-art results in predicting hospital readmission and mortality (Huang et al., 2019; Gu et al., 2021). These works highlight that text contains complementary information to structured data, for example, documenting patient symptoms or social factors not in coded form. However, directly modeling text is challenging due to noise and length; researchers often rely on the full text or bag-of-words representations, which lose temporal granularity. Our approach differs by preserving the chronological sequence of events mentioned in text, effectively treating narrative data as a time series. This connects to recent studies on re-purposing LLMs for traditional time-series forecasting tasks: techniques like Time-LLM have shown that language models can be “re-programmed” to handle numeric sequences by mapping time points to token sequences (Jin et al., 2023). In the clinical domain, our work is the first to apply language models to text-derived time series, bridging free-text analysis with time-aware prediction. We also relate to prior sepsis prediction models. Traditional models use vital sign trends or clinical scores (e.g., SOFA, SAPS) to predict sepsis onset or mortality (Hou et al., 2020; Noroozizadeh et al., 2023). Unlike these, our method relies solely on narrative descriptions rather than real-time vitals, and is thus complementary: it demonstrates how knowledge embedded in case reports (which often include the clinicians’ interpretations and context) can be harnessed for prediction. This could be especially useful in settings where detailed narratives are available (such as case report literature or perhaps patient histories) but structured data is sparse.
Large Language Models in Healthcare:
The advent of powerful LLMs like GPT-3 and GPT-4 has sparked interest in various healthcare applications, from medical question-answering to clinical note summarization. These models, trained on massive text corpora, contain a wealth of medical knowledge and the ability to perform reasoning to some extent. For instance, GPT-4 has been tested on medical board exam questions with impressive results, and pilot studies show LLMs can abstract patient records for quality reporting (Boussina et al., 2024). Nevertheless, using LLMs for forecasting in healthcare is relatively uncharted territory. A limitation of out-of-the-box LLMs is their tendency to generate plausible-sounding but unverified content (hallucinations) and their lack of alignment with specific clinical data distributions. Our benchmark explicitly evaluates LLMs through prompting (zero-shot and few-shot) to see if they can ”forecast” events when given a patient’s story so far. We also compare that against fine-tuned models that are directly trained on the task, to quantify the gap between general-purpose and task-specific modeling in this context. This comparison adds to the growing body of evidence that, while large models have broad knowledge, domain-specific fine-tuning or structuring of input is often necessary for reliable performance in clinical tasks (Huang et al., 2019; Gu et al., 2021). Some recent analyses suggest that medical-domain fine-tuning may not consistently improve performance over foundation models (Jeong et al., 2024), further motivating the need for task- and input-specific evaluation setups like ours.
3 Methods
 [^1] We will be releasing our code upon publication.
We outline the task specification, define key terms, describe the extraction and annotation process, and present the evaluation criteria (Figure 2).
Figure 2: Flowchart of the T2S2 forecasting analysis pipeline
3.1 Dataset extraction
We utilized the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis. Building on the method described in Wang and Weiss (2025), we extract the body of each case by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus. To maintain relevance, we include only those documents that contain case-insensitive matches to the regular expressions (case report|case presenta) and year-? ?old. Prior work has shown that this regex-based approach demonstrated notably higher specificity compared to PubMed metadata on a random sample of 20 case reports Noroozizadeh and Weiss (2025).
As demonstrated in prior work Noroozizadeh and Weiss (2025), we applied additional filtering to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care. We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender. For this step, both GPT-4o and Llama-3.1-8B-Instruct were employed. Reports describing more than one case were excluded. A case report was retained if either model classified it as involving a Sepsis-3 diagnosis.
This multi-step process yielded a total of 2319 Sepsis-3 case reports. From this set, we randomly selected two subsets for evaluation: a group of 10 reports (sepsis-10), which underwent expert clinical annotation to serve as a gold standard, and a larger group of 100 reports (sepsis-100) used for broader testing (Figure 2).
3.2 Textual time-series annotation
As established in previous work Noroozizadeh and Weiss (2025), we aimed to construct a textual time series case report corpus of Sepsis-3 patients from the PMOA sepsis case reports. A textual time series refers to a sequence of clinical findings, each paired with a timestamp (either absolute or relative to the case presentation time), corresponding to an individual patient. Here, a clinical finding is defined as a free-text expression describing an entity that pertains to or may impact a person’s health.
A textual time-series is different from conventional approaches to clinical concept annotation (e.g., Sun et al. (2013); Uzuner et al. (2011)) by extending the span of extracted text to better capture the specificity and contextual meaning of each clinical finding. In our work, the interpretation of a clinical finding differs from the i2b2 concept guidelines in the following respects : (1) Clinical findings are not restricted to single prepositional phrases following a markable entity. For instance, instead of splitting “pain in chest that radiates substernally” into less informative parts, we retain the full phrase as a single, complete finding to preserve its meaning and context. (2) To improve clarity, coordinated phrases are broken into individual findings. For example, “metastases in the liver and pancreas” becomes “metastasis in the liver” and “metastasis in the pancreas.”
We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports. We refer to this as Textual Time-Series for Sepis (T2S2) The exact LLM extraction prompt can be found in Appendix E. Our dataset was split into training (n = 1795), testing T2S2-test (n = 244) and the 2 external validation sets sepsis 10 (n = 10) and sepsis 100 (n = 90). The training set was further split into 80% for model training and 20% for model validation (e.g. early stopping).
3.3 Task Definition
Using the timelines extracted from the annotations, we defined two primary forecasting tasks and a survival analysis task:
Event Occurrence Prediction:
Given a prefix of the clinical timeline (all events up to a certain time $t$), the model is tasked with predicting whether each of the immediate next $k$ events occurs within a specified time horizon. This setup is repeated across multiple time cut-offs to simulate an “online” forecasting scenario, where the model must output a binary label for each of the next $k$ events: does this event occur within $h$ hours after $t$? Time horizons used include 1 hour, 24 hours (1 day), and 168 hours (1 week). The task is framed as a series of binary classification problems, with one binary decision per event. Evaluation is based on precision, recall, and F1 score, averaged across event positions for each time horizon.
Temporal Ordering Prediction:
This task assesses the model’s ability to reconstruct the correct sequence of future events. At each time cut-off $t$, we extract the next $k$ events from the timeline and remove their timestamps. The model must correctly output a permutation over these events that matches their true chronological order. This is framed as a ranking task, evaluated by computing the pairwise concordance between the predicted and true orderings (e.g., proportion of correctly ordered pairs). This tests whether the model can infer temporal progression from unordered event content alone.
Survival Analysis for Mortality Time:
We include a classical survival analysis task to model time until death. Many case reports specify whether the patient died and, if so, when (e.g., “the patient died on hospital day 10”), which enables us to define a time-to-event outcome. For this task, we evaluate models at predefined cut-off times – specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week) – and use the extracted event history up to each cut-off as input. A survival model is trained to predict the probability of survival over time beyond each cut-off. We evaluate performance using the time-dependent concordance index to assess how well the predicted survival times align with actual outcomes.
3.4 Modeling Approaches
To evaluate performance on the event forecasting and survival prediction tasks, we implement three modeling paradigms: (i) large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), and (iii) BERT-style encoder models with task-specific fine-tuned heads. In Appendix C and  D, we provide detailed explanations of how we approach the event occurrence prediction, temporal ordering prediction, and survival analysis tasks using encoder-only models.
Fine-tuned LLMs:
We apply instruction-tuned decoder-only models from the LLaMA and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants. Each model is paired with a lightweight multilayer perceptron (MLP) head trained for classification or ranking, depending on the task. The input consists of a text-formatted prefix of events (e.g., a clinical timeline up to time $t$), optionally accompanied by an instruction template. The output layer produces task-specific predictions: binary labels for event occurrence or a permutation over $k$ events for ordering. Training is conducted using cross-entropy loss for classification and pairwise ranking loss for ordering.
Prompted LLMs:
In the zero- or few-shot setting, we use the same LLM architectures as above, but without any fine-tuning. Instead, we supply structured prompts at inference time to guide the model toward task-specific outputs. Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format. Prompt designs are customized for each task and are detailed in Appendix F. Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
Fine-tuned Encoder-Only Models:
We also evaluate a range of encoder-based models trained end-to-end on each task. Architectures include BERT-base-uncased, RoBERTa-base, DeBERTa-v3-small, ModernBERT-small, and ModernBERT-large. For each model, we append a task-specific MLP head for event occurrence or ordering prediction. The model input is the same flattened prefix of the event sequence used in other settings, tokenized and formatted according to the respective architecture. These models are trained using standard supervised learning objectives and evaluated on the same metrics as other methods: F1 score for event occurrence, and pairwise concordance for temporal ordering. For survival analysis, the final hidden states are passed into a time-to-event prediction head and evaluated using the time-dependent concordance index.
4 Results
Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1. Bold values indicate best in category amongst all models.
T2S2 (DeepSeek R1) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
Prompting
L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452
Encoder-MLP head
ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
Encoder-masking-fine-tuned
ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
Encoder-masking-zeroshot
ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
Table 2: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from Llama 3.3-70B. Bold values indicate best in category amongst all models.
T2S2 (Llama 3.3) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
Prompting
L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
Encoder-MLP head
ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
Encoder-masking-fine-tuned
ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
Encoder-masking-zeroshot
ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
4.1 Event forecast within next 24 hours: F1 performance
Our results demonstrate a clear advantage of encoder-based models over LLMs in event forecasting, for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2) and particularly in short-term and medium-term prediction windows. The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A. Encoder models, particularly those with a fine-tuned MLP head, achieve substantially higher F1 scores than the LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
Among encoder models, different training strategies yield distinct performance trends. The MLP head fine-tuning approach consistently outperforms both fine-tuned masking and zero-shot masking models, particularly in long-horizonforecasting. ModernBERT-base achieves the highest overall F1 scores across all forecasting horizons, particularly at 168 hours, where it outperforms all other models. ModernBERT-large follows closely but does not consistently surpass its base variant, suggesting diminishing returns with increased model size. Meanwhile, zero-shot masking models, particularly standard BERT and RoBERTa, fail to make meaningful predictions, often yielding near-zero F1 scores at the 1-hour mark. However, DeBERTa-small and ModernBERT-base exhibit relatively better zero-shot performance, indicating that certain architectures generalize better without task-specific training.
Performance trends also vary across different datasets. Across all models, F1 scores improve as the forecasting window increases, with 1-hour predictions being the most challenging and 168-hour predictions yielding the highest scores. This suggests that longer-term trends are more predictable, whereas immediate event forecasting remains difficult, possibly due to the sparsity and volatility of short-term event occurrences. Comparing internal and external datasets, models perform worse on external validation sets (sepsis-10 and sepsis-100) than on the internal T2S2 test set, with a particularly noticeable drop in F1 at the 1-day horizon. This indicates potential generalization challenges, as models trained on internal data struggle to adapt to distributional shifts in clinical settings. Notably, ModernBERT models maintain a relative performance edge even in external datasets, reinforcing their robustness across different domains.
4.2 Temporal ordering of forecasted events: Concordance
In addition to event occurrence prediction, models are evaluated on their ability to correctly rank the order of upcoming events, measured through concordance (c-index) for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2). The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A.
Similar to F1 trends, encoder-based models significantly outperform LLMs, with LLMs failing to achieve reliable ordinal rankings even when fine-tuned or prompted. Within encoder models, different training approaches show distinct performance characteristics: MLP-head fine-tuned models dominate in event occurrence classification (F1) but do not always yield the best concordance values. Instead, fine-tuned masking models, particularly ModernBERT-large, achieve the highest concordance scores, suggesting that this approach is better suited for ordering rather than binary classification.
Performance differences across datasets also highlight the challenges of generalization. While internal (T2S2) test results show relatively strong concordance for fine-tuned models, c-index values drop considerably on external datasets, particularly for models that are not ModernBERT-based. Among the best-performing models, ModernBERT-large (fine-tuned masking) achieves the highest c-index across both internal and external datasets, reinforcing its strength in capturing event order relationships. However, zero-shot masking models continue to struggle, mirroring their poor F1 performance. Additionally, while MLP-head fine-tuned models maintain high F1 scores, they exhibit slightly lower concordance than fine-tuned masking models, particularly at the 168-hour horizon. This suggests that different optimization strategies may be required depending on whether the task prioritizes strict classification accuracy or relative event ordering.
Overall, the results confirm that ModernBERT architectures are the strongest performers across both F1 and concordance evaluations, with fine-tuned MLP-head models excelling in binary prediction tasks and fine-tuned masking models achieving superior ranking performance. However, all models face challenges in generalizing to external datasets, emphasizing the need for domain adaptation techniques to improve real-world forecasting robustness.
4.3 Survival analysis
Model Performance across Timepoints and Cohorts.
Table 3 presents the time-dependent concordance index (concordance_td) for all models across three evaluation cohorts—t2s2, Sepsis10, and Sepsis100—at three observation windows (0h, 24h, and 168h). For each cohort and timepoint, we report performance from the best-fitting survival model (selected based on validation concordance).
Across cohorts, we observe that providing additional observation time (e.g., from 0h to 168h) does not consistently improve predictive performance. In the t2s2 cohort, several models achieve their peak concordance at the earliest or intermediate timepoints (e.g., DeepSeek-R1-Llama-70B peaks at 0h, while Llama-3.3-70B-Instruct peaks at 24h), with some degradation at 168h. In Sepsis100, certain models benefit from longer observation (e.g., DeepSeek-R1-Distill-Llama-8B improves steadily across timepoints), but others decline or plateau. The Sepsis10 cohort exhibits high and uniform performance across models, particularly after 24h, suggesting potential ceiling effects or limited evaluation variability in that subset.
Instruction-tuned LLMs such as Llama-3.3-70B-Instruct, DeepSeek-R1-Distill- Llama-70B, and Llama-3.1-8B-Instruct consistently outperform conventional transformer baselines like bert-base-uncased or roberta-base. This trend is particularly notable in the t2s2 and Sepsis100 cohorts, where LLMs demonstrate stronger early signal detection at 0h and 24h. These results suggest that large-scale pretraining and instruction tuning confer advantages in extracting actionable clinical representations, even with limited temporal context.
Taken together, these findings highlight that strong early performance does not require extended observation windows and that pretrained instruction-following models can offer substantial benefits for survival prediction in real-world, noisy clinical text scenarios.
Model t2s2 Sepsis-10 Sepsis-100
0h 24h 168h 0h 24h 168h 0h 24h 168h
bert-base-uncased 0.60 0.61 0.53 1.00 1.00 1.00 0.58 0.62 0.58
roberta-base 0.55 0.60 0.60 0.75 1.00 1.00 0.55 0.58 0.57
deberta-v3-small 0.56 0.57 0.57 0.75 0.00 0.75 0.52 0.49 0.51
ModernBERT-base 0.52 0.58 0.58 1.00 1.00 1.00 0.54 0.51 0.51
ModernBERT-large 0.53 0.59 0.58 0.50 1.00 1.00 0.59 0.55 0.49
DeepSeek-R1-Llama-70B 0.64 0.63 0.59 0.50 1.00 1.00 0.61 0.53 0.55
Llama-3.3-70B-Instruct 0.62 0.63 0.58 0.50 1.00 0.75 0.64 0.57 0.53
DeepSeek-R1-Llama-8B 0.60 0.58 0.58 1.00 1.00 1.00 0.53 0.66 0.65
Llama-3.1-8B-Instruct 0.62 0.61 0.61 1.00 1.00 1.00 0.60 0.59 0.61
Table 3: Concordance index (concordance_td) for each model across three test sets (t2s2, Sepsis10, Sepsis100), evaluated at three observation time thresholds (0h, 24h, 168h).
4.4 Sensitivity analysis
To better understand the robustness and design implications of our forecasting models, we conduct two complementary sensitivity analyses focused on the temporal structure and completeness of input sequences. First, we compare two strategies for ordering clinical events within textual time-series (TTS) inputs: text-ordered, where events follow their extraction sequence from the original case report narratives, and time-ordered, where events are chronologically sorted. This comparison assesses the relative benefits of preserving narrative flow versus enforcing temporal alignment. Second, we evaluate the impact of input sparsity by randomly masking portions of the clinical history, simulating scenarios with incomplete documentation or missing data. Together, these analyses shed light on the importance of chronological ordering and data completeness for both classification accuracy and temporal event ranking.
4.4.1 Time-ordered vs text-ordered training
We evaluated the impact of training on text-ordered versus time-ordered textual time-series (TTS) annotations. In the text-ordered condition, events in the textual time-series order follow the order in which they were extracted from the case reports using LLMs (DeepSeek-R1 and Llama-3.3-70B), while in the time-ordered condition, they are sorted chronologically in time. This comparison aimed to assess the impact of chronological versus narrative ordering of events on forecasting performance
As shown in Table 4, time-ordered training generally leads to higher concordance (C-index), particularly for DeepSeek-R1 annotatiobs across both internal (T2S2) and external test datasets (sepsis-10 and sepsis-100). However, F1 scores are less consistently affected by the ordering strategy. In several cases—especially with Llama-3.3-70B annotations on external datasets, text-ordered training results in equal or better F1 scores, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order. These results highlight a trade-off between chronological alignment and narrative structure in optimizing forecasting performance.
Table 4: Comparison of text-ordered and time-ordered training for Deep Seek R1 and Llama 3.3 annotations on F1 scores and C-index values.
Dataset Metric DeepSeek-R1 Llama 3.3-70B
Text-ordered Time-ordered Text-ordered Time-ordered
T2S2 Test F1 (1hr) 0.290 0.306 0.311 0.257
F1 (1day) 0.569 0.576 0.648 0.645
F1 (7days) 0.876 0.877 0.898 0.903
C-index 0.659 0.672 0.586 0.632
Sepsis10 F1 (1day) 0.361 0.428 0.423 0.382
C-index 0.598 0.576 0.526 0.595
Sepsis100 F1 (1day) 0.615 0.607 0.662 0.616
C-index 0.609 0.613 0.541 0.612
4.4.2 Impact of Randomly Masking History on F1 and Concordance
Further, we systematically evaluate how randomly omitting portions of clinical history impacts both forecasting accuracy and the model’s ability to correctly order clinical events. Towards this end, we simulate incomplete patient histories by randomly masking input time steps at varying rates. We refer to this as the timestep drop rate (TDR), ranging from 0% (no masking) to 90% (severely truncated history). We assess the effect of this masking on both classification performance (F1 score at 1 day) and ranking performance (concordance index) across three datasets—T2S2-DS, Sepsis-10, and Sepsis-100—using both DeepSeek-R1 (Table 5) and Llama-3.3-70B annotations (Table 8 in Appendix B).
F1 scores decline progressively with increasing timestep dropout rate (TDR), with sharp reductions beyond 60% (Table 5). At TDR = 0%, the F1 scores for T2S2-DS, Sepsis-10, and Sepsis-100 are 0.576, 0.428, and 0.607, respectively. By TDR = 90%, these values fall to 0.423, 0.200, and 0.510, with Sepsis10 showing the steepest decline. Between 30% and 60% dropout, performance remains relatively stable for T2S2-DS and Sepsis100, while Sepsis10 exhibits a more pronounced drop at TDR = 60%. Similar trends are observed with Llama-3.3 annotations (Table 8 in Appendix B).
In contrast, concordance (c-index) remains largely consistent across dropout levels. T2S2-DS (Table 5) reaches its highest concordance (0.678) at TDR = 60%, while Sepsis-10 and Sepsis-100 vary only marginally. This suggests that while masking temporal history substantially reduces classification accuracy, it has limited impact on event ranking performance. Comparable patterns are observed for the Llama-3.3 models (Table 8).
Overall, DeepSeek models show notable degradation in F1 scores when more than 30% of history is removed—particularly for datasets sensitive to longer-term dependencies—whereas concordance scores remain relatively stable (Table 5 and Table 8).
Table 5: Effect of Randomly Masking History (DeepSeek R1 annotations) on F1 and Concordance. TDR : Timestep drop rate
TDR F1 (1 day) Concordance (c-index)
T2S2-DS Sepsis-10 Sepsis-100 T2S2-DS Sepsis-10 Sepsis-100
0% 0.576 0.428 0.607 0.672 0.576 0.613
15% 0.572 0.415 0.604 0.667 0.564 0.616
30% 0.548 0.357 0.591 0.672 0.542 0.619
60% 0.542 0.310 0.588 0.678 0.578 0.616
90% 0.423 0.200 0.510 0.661 0.561 0.614
5 Discussion and Conclusion
Our findings highlight the superiority of encoder-based models over decoder-based LLMs for event forecasting, emphasizing the limitations of autoregressive models in structured prediction tasks. Despite prompting and fine-tuning, LLMs struggle with short- and medium-term forecasting, reinforcing the need for models optimized for sequence encoding rather than generative token prediction.
Among encoder models, ModernBERT-base and ModernBERT-large consistently perform best, with MLP-head fine-tuning excelling in F1 scores and fine-tuned masking models achieving higher concordance (c-index), particularly in long-term forecasting. This suggests that classification and ranking tasks benefit from distinct optimization strategies, and a hybrid training approach may further enhance performance.
The gradual improvement of F1 scores over longer time horizons indicates that event patterns become more predictable over time, whereas short-term forecasting remains challenging due to higher variability. This limitation suggests that incorporating temporal modeling techniques, such as recurrent or convolutional structures, could refine short-term predictions while maintaining strong long-range forecasting.
Performance drops on external validation datasets (sepsis-10 and sepsis-100) highlight generalization challenges, with models struggling to adapt to shifts in event distributions. While ModernBERT models exhibit relative robustness, domain-specific fine-tuning or adaptive learning strategies are needed to bridge this gap.
Our survival analysis results complement the forecasting findings by revealing that strong predictive performance is often achievable with limited observation windows. Several models reach peak concordance at early (0h) or intermediate (24h) timepoints rather than at the longest window (168h), particularly in the T2S2 and Sepsis100 cohorts. Instruction-tuned LLMs such as Llama-3.3-70B-Instruct and DeepSeek-R1-Distill-Llama-70B outperform traditional transformer baselines in these survival settings, suggesting that instruction tuning and large-scale pretraining confer meaningful advantages for modeling time-to-event outcomes from clinical narratives in the textual time series setup.
Sensitivity analyses further highlight the importance of temporal structure and data completeness. When comparing time-ordered and text-ordered training regimes, we find that time-ordered training generally improves concordance, particularly for DeepSeek-R1 annotations, while text-ordering can yield competitive or better F1 scores on external datasets. This underscores a trade-off between temporal alignment and narrative coherence, with implications for designing learning objectives based on downstream task priorities. Moreover, robustness experiments simulating documentation sparsity reveal that F1 scores degrade significantly as more clinical history is masked, particularly beyond a timestep dropout rate of 60%. In contrast, concordance remains relatively stable under these conditions, indicating that event ranking is less sensitive to partial history than event classification.
Overall, our results demonstrate that encoder-based models, particularly ModernBERT, are well-suited for event forecasting but require enhancements in short-term prediction and cross-domain generalization. In contrast, for survival analysis tasks, we find that instruction-tuned LLMs have more predictive power. Future work should explore hybrid training objectives that balance classification and ranking performance, incorporate temporal modeling architectures, and investigate domain adaptation and pretraining strategies that explicitly account for data sparsity and real-world variability that we observe in our clinical textual time series.
Acknowledgments
This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health. This work utilized the computational resources of the NIH HPC Biowulf cluster.
Appendix A Forecasting performance on DeepSeek-R1 and Llama-3.3-70B textual time-series annotations
Tables 6 and 7 present the forecasting performance of our models on textual time-series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively. We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
Table 6: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1. Bold indicates best in that particular category amongst all models.
T2S2 (DeepSeek R1) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
L3.3 70B 0.000 0.295 0.764 0.631 0.312 0.569 0.404 0.597
DS-L3.1 8B 0.000 0.492 0.650 0.630 0.381 0.566 0.601 0.600
L3.1 8B 0.000 0.355 0.650 0.630 0.114 0.575 0.442 0.604
Prompting
L3.3 70B 0-shot 0.085 0.368 0.560 0.236 0.452
L3.3 70B 0-shot ordinal 0.082 0.352 0.646 0.289 0.483
L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452
L3.3 70B few-shot ordinal 0.063 0.356 0.682 0.313 0.480
Encoder-MLP head
BERT 0.264 0.550 0.871 0.501 0.441 0.498 0.614 0.501
RoBERTa 0.250 0.562 0.866 0.500 0.405 0.500 0.614 0.500
DeBERTa-small 0.262 0.570 0.871 0.601 0.382 0.498 0.617 0.556
ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
ModernBERT-large 0.277 0.568 0.874 0.598 0.391 0.539 0.602 0.553
Encoder-masking-fine-tuned
BERT 0.166 0.432 0.690 0.648 0.323 0.554 0.482 0.600
RoBERTa 0.169 0.416 0.688 0.646 0.313 0.539 0.457 0.609
DeBERTa-small 0.175 0.435 0.695 0.671 0.325 0.579 0.492 0.617
ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
ModernBERT-large 0.177 0.460 0.683 0.663 0.342 0.615 0.496 0.632
Encoder-masking-zeroshot
BERT 0.000 0.000 0.000 0.495 0.000 0.513 0.000 0.494
RoBERTa 0.049 0.207 0.375 0.500 0.137 0.500 0.318 0.500
DeBERTa-small 0.170 0.460 0.323 0.497 0.268 0.512 0.376 0.509
ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
Table 7: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from Llama-3.3-70B.
T2S2 (Llama 3.3) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
L3.3 70B 0.265 0.466 0.807 0.619 0.160 0.567 0.466 0.597
DS-L3.1 8B 0.264 0.584 0.837 0.624 0.258 0.576 0.605 0.598
L3.1 8B 0.338 0.565 0.823 0.619 0.368 0.569 0.536 0.596
Prompting
L3.3 70B 0-shot 0.085 0.443 0.616 0.236 0.452
L3.3 70B 0-shot ordinal 0.063 0.454 0.707 0.289 0.483
L3.3 70B few-shot 0.099 0.457 0.703 0.360 0.452
L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
Encoder-MLP head
BERT 0.224 0.622 0.899 0.500 0.400 0.500 0.612 0.500
RoBERTa 0.180 0.608 0.900 0.500 0.364 0.500 0.597 0.500
DeBERTa-small 0.258 0.641 0.898 0.603 0.427 0.585 0.616 0.592
ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
ModernBERT-large 0.312 0.622 0.897 0.500 0.404 0.500 0.623 0.500
Encoder-masking-fine-tuned
BERT 0.131 0.471 0.741 0.622 0.320 0.580 0.467 0.622
RoBERTa 0.129 0.484 0.659 0.625 0.329 0.587 0.472 0.614
DeBERTa-small 0.135 0.468 0.693 0.648 0.318 0.602 0.462 0.623
ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
ModernBERT-large 0.124 0.470 0.712 0.646 0.318 0.586 0.468 0.631
Encoder-masking-zeroshot
BERT 0.000 0.000 0.000 0.504 0.000 0.515 0.000 0.506
RoBERTa 0.074 0.279 0.445 0.500 0.083 0.500 0.319 0.500
DeBERTa-small 0.008 0.439 0.839 0.562 0.296 0.518 0.518 0.501
ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
ModernBERT-large 0.109 0.413 0.622 0.499 0.326 0.509 0.339 0.518
Appendix B Masking history: Llama-3.3-70B textual time-series annotations
Table 8 examines the impact of randomly masking portions of the input history on model performance when using Llama-3.3-70B-generated annotations. We evaluate performance across three datasets (T2S2-L33, Sepsis-10, and Sepsis-100) using F1 scores for prediction of next $k$ events happening with the next day and concordance index (c-index) correctly ordering the next $k$ events. As the proportion of masked history increases, we observe a general decline in F1 performance, particularly at higher masking rates (e.g., 90%), suggesting that access to historical context is important for accurate forecasting. Interestingly, the concordance index remains relatively stable, indicating the model’s robustness in ranking the next $k$ events with partial input histories.
Table 8: Effect of Randomly Masking History on F1 and Concordance (Llama-3.3 Annotations)
TDR F1 (1 day) Concordance (c-index)
T2S2-L33 Sepsis-10 Sepsis-100 T2S2-L33 Sepsis-10 Sepsis-100
0% 0.645 0.431 0.626 0.632 0.595 0.612
15% 0.628 0.422 0.624 0.635 0.592 0.610
30% 0.634 0.423 0.621 0.634 0.586 0.598
60% 0.604 0.391 0.605 0.634 0.577 0.613
90% 0.531 0.328 0.534 0.622 0.546 0.607
Appendix C Textual Time Series Forecasting Framework of Encoder Models
We develop a general framework for predictive modeling over textual time series of the encoder-only models, designed to accommodate the irregular, event-based structure common to clinical case reports. As highlighted Section 3.2, each data instance consists of a sequence of clinical events, where each event is a free-text string describing a clinical finding, intervention, or diagnosis, paired with a timestamp in hours relative to the time of admission. The format resembles a two-column table, where the first column contains textual clinical events and the second column contains integer-valued timestamps.
Input Preprocessing and Timeseries Construction.
From each case report, we construct multiple training examples for sequence forecasting. For a fixed forecast horizon parameter $K$, we generate overlapping windows of historical events and forecast targets. Specifically, for each case report with $L$ rows, we construct a sequence of up to $L-K$ forecasting examples. Each example is structured as follows:
  • A history segment $\mathcal{H}_{t}=\{(e_{i},t_{i})\}_{i=1}^{n}$, consisting of clinical events $e_{i}$ occurring at timestamps $t_{i}\leq t$.
  • A forecast target $\mathcal{F}_{t}=\{(f_{j},\tau_{j})\}_{j=1}^{K}$, consisting of the next $K$ events after $t$.
To preserve the temporal coherence of events, we treat all events with the same timestamp as an atomic unit. That is, if multiple events share the same timestamp, they are included together in the history or forecast block, and the next forecasting window is only advanced after the full timestamp block has been consumed.
Model Input Representation.
Each training example is serialized into a transformer-compatible input by linearizing the historical events and forecast events. Importantly, we preserve both temporal and textual content by encoding each event in the form “t:e” where t is the timestamp and e is the event string. The input to the language model takes the form:
[…]
This is followed by additional target-specific components, as described below.
Language Models.
To test the generality of our framework, we experiment with a variety of pre-trained encoder-only language models, including:
  • bert-base-uncased (BERT; Devlin et al., 2018)
  • roberta-base (Liu et al., 2019)
  • microsoft/deberta-v3-small (He et al., 2021)
  • answerdotai/ModernBERT-base and answerdotai/ModernBERT-large (pretrained on modern clinical narratives)
Each model is used with its corresponding tokenizer. For models with SentencePiece tokenization (e.g., DeBERTa), we use the appropriate fast tokenizer such as DebertaV2Tokenizer to ensure compatibility. All inputs are truncated from the left (i.e., keeping the most recent tokens) if they exceed the model’s context window (typically 512 tokens).
Prediction Tasks.
We define two primary prediction tasks over $\mathcal{F}_{t}$, each formulated as a binary classification problem.
1. Event Ordering (Concordance) Task.
The goal in this task is to predict the correct temporal ordering between future events. For each example, we generate all comparable pairs of events in $\mathcal{F}_{t}$, defined as pairs $(f_{i},f_{j})$ where $\tau_{i}\neq\tau_{j}$. For each pair, we form a binary classification example as follows:
  • Input: The serialized history segment followed by:
[…]
  • Label: $y=1$ if $\tau_{i}<\tau_{j}$, else $y=0$
The model output is taken from the [CLS] token, passed through a linear classification head, and softmaxed to produce a binary probability. We optimize cross-entropy loss across all pairwise comparisons. At evaluation time, we report the concordance index (c-index), which measures the proportion of correctly predicted pairs among all comparable pairs.
2. Time-Window Classification Task.
This task evaluates whether each forecast event occurs within a fixed prediction window $H$ hours of the last historical event. For each $f_{j}\in\mathcal{F}_{t}$, we generate an input:
  • Input: The serialized history segment followed by:
[…]
  • Label: $y=1$ if $\tau_{j}-t_{n}\leq H$, where $t_{n}$ is the timestamp of the last historical event.
This is again treated as a binary classification task over the [CLS] embedding, using a linear layer and softmax. We report macro-averaged F1 scores over all prediction examples.
Training and Evaluation.
Each task is trained separately using AdamW with linear learning rate decay and early stopping based on validation loss. We save the model checkpoint with the best validation performance. During inference, we reload this checkpoint and evaluate on a held-out test set (which is obtained from textual time series of unseen case reports).
For both tasks, we extract softmax probabilities rather than hard labels to allow for threshold calibration and ROC-based analyses. All training is conducted using PyTorch and HuggingFace Transformers.
Appendix D Survival Modeling with Language Model Embeddings
To evaluate the prognostic value of textual information encoded in large language models (LLMs), we adopt a two-stage framework: (1) extraction of fixed-dimensional sequence embeddings from various pre-trained LLMs, and (2) downstream survival modeling using these embeddings as covariates.
In the first stage, we process each textual time series (which is converted to a textual context of ”(time) clinical event [SEP] … [SEP] (time) clinical event [SEP]”) using a suite of LLMs to obtain dense vector representations that summarize the content of the input sequence. For models belonging to the encoder family, including bert-base-uncased, roberta-base, deberta-v3-small, ModernBERT-base, and ModernBERT-large, we extract the final hidden state corresponding to the [CLS] token, which is conventionally used to represent the entire sequence in classification tasks. This token-specific embedding serves as a compact, sequence-level representation.
For decoder-based models that do not utilize a dedicated [CLS] token, including DeepSeek-R1-Distill-Llama-70B, Llama-3.3-70B-Instruct, DeepSeek-R1-Distill-Llama-8B, and Llama-3.1-8B-Instruct, we compute the mean-pooled embedding over the last hidden states of all non-padding tokens. This pooling strategy yields a fixed-length vector that captures the overall semantic content of the input while mitigating the impact of padding artifacts.
In the second stage, these embedding vectors are used as input covariates to three survival models, each designed to capture time-to-event dynamics in different ways:
  • Random Survival Forest (RSF): A nonparametric ensemble method based on decision trees, capable of handling complex interactions and high-dimensional inputs. The model is tuned over the number of trees (n_estimators), the minimum number of samples required to split an internal node (min_samples_split), and the minimum number of samples at a leaf node (min_samples_leaf).
  • DeepSurv: A neural-network generalization of the Cox proportional hazards model. The architecture comprises fully connected layers with varying widths (num_nodes) and dropout rates, with a fixed number of training epochs.
  • DeepHit: A multi-task neural network that jointly models the discrete hazard and survival functions. As with DeepSurv, we perform a grid search over hidden layer sizes and dropout rates while keeping the number of epochs fixed.
Hyperparameter tuning for each survival model is performed using a validation split from the training set. For each combination of language model and survival model, we conduct grid search over the respective hyperparameter space, selecting the configuration that maximizes the time-dependent concordance index on the validation set. Final evaluation is conducted on held-out test sets (t2s2_test, sepsis10, sepsis100), and results are aggregated across survival metrics to assess model performance.
This framework allows for a systematic evaluation of how different pre-trained LLMs, and their associated embedding strategies, contribute to downstream survival prediction tasks.
Appendix E LLM prompt to generate clinical textual time-series from PMOA case reports
You are a physician. Extract the clinical events and the related time stamp from the case report. The admission event has timestamp 0. If the event is not available, we treat the event, e.g. current main clinical diagnosis or treatment with timestamp 0. The events that happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time. The timestamp are in hours. The unit will be omitted when output the result. If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation. We want to predict the future events given the events happened in history. For example, here is the case report.
An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash. Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks. With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome. The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present. One day later the patient was discharged.
Let’s find the locations of event in the case report, it shows that four weeks ago of fever and rash, four weeks ago, he was diagnosed with acne and receive treatment. So the event of fever and rash happened four weeks ago, 672 hours, it is before admitted to the hospital, so the time stamp is -672. diffuse erythematous or maculopapular eruption with pruritus was documented on the admission exam, so the timestamp is 0 hours, since it happens right at admission. DRESS syndrome has no specific time, but it should happen soon after admission to the hospital, so we use our clinical judgment to give the diagnosis of DRESS syndrome the timestamp 0. then the output should look like: 18 years old | 0
male | 0
admitted to the hospital | 0 fever | -72
rash | -72
acne | -672
minocycline | -672
increased WBC count | 0
eosinophilia| 0
systemic involvement| 0
diffuse erythematous or maculopapular eruption| 0
pruritis | 0
DRESS syndrome | 0
fever persisted | 0
rash persisted | 0
discharged | 24
Separate conjunctive phrases into its component events and assign them the same timestamp (for example, separation of ‘fever and rash’ into 2 events: ‘fever’ and ‘rash’). If the event has duration, assign the event time as the start of the time interval. Attempt to use the text span without modifications except ‘history of’ where applicable. Include all patient events, even if they appear in the discussion; do not omit any events; include termination/discontinuation events; include the pertinent negative findings, like ‘no shortness of breath’ and ‘denies chest pain’. Show the events and timestamps in rows, each row has two columns: one column for the event, the other column for the timestamp. The time is a numeric value in hour unit. The two columns are separated by a pipe ‘|’ as a bar-separated file. Skip the title of the table. Reply with the table only.
Appendix F Prompting strategies for forecasting tasks
Zero-shot:
You are an expert physician.
Reply to the prompt with structured predictions in a k-item, bar-separated row. For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: 0 | 1 | 0 would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Zero-shot ordinal:
You are an expert physician.
Reply to the prompt with structured predictions in a k-item, bar-separated row. Assess whether the event happens in the time window using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen. For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: $1|5|1$ would be the correct response. Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Few-shot:
You are an expert physician.
Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window. Example 1. If there k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 | 1 | 0 would be the correct response.
Example 2. Suppose the time series is:
admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
Then, if the prediction is at 5 hours, the time preceding ‘‘--- PREDICTION WINDOW ---’’, the events admitted, history of diabetes, and elevated troponin have occurred. Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor, and discharge happen. So the correct response for the example is: 1 | 1 | 0 | 0 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
Reply with the answer only. Do NOT provide any other text. Do NOT explain.
Few-shot ordinal:
You are an expert physician.
Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window, using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen.
Example 1. If there k=3 events (A, B, C) and only B occurs in the time window, then: 1 | 5 | 1 would be the correct response.
Example 2. Suppose the time series is:
admitted | 0
history of diabetes | 0
elevated troponin | 5
--- PREDICTION WINDOW ---
cardiac catheterization | 6
admission to icu | 24
transfer to floor | 72
discharge | 144
office visit | 1440
Then, if the prediction is at 5 hours, the time preceding ‘‘--- PREDICTION WINDOW ---’’, the events admitted, history of diabetes, and elevated troponin have occurred. Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor and discharge happen. So the an excellent response for the example is: 5 | 4 | 1 | 1 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
Reply with the answer only. Do NOT provide any other text. Do NOT explain.

### v3

@@
-Forecasting from Clinical Textual Time Series: Adaptations of the Encoder and Decoder Language Model Families
+Forecasting Clinical Risk from Textual Time Series: Structuring Narratives for Temporal AI in Healthcare
 Abstract
-Clinical case reports encode rich, temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data.
+Clinical case reports encode temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data.
 In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings–extracted via an LLM-assisted annotation pipeline–serve as the primary input for prediction.
@@
 This highlights the additional benefit that can be ascertained from time-ordered corpora, with implications for temporal tasks in the era of widespread LLM use.
-**footnotetext: These authors contributed equally to this work
-1 Introduction
-Clinical risk forecasting is a cornerstone problem in health care decision making that guides eligibility and reasoning about interventions.
-Because risk prediction performs best with access to rich data streams, great lengths are taken to capture the relevant health factors that affect decision making.
-However, while many forecasting tasks are performed on structured tabular data, evidence suggests that, despite widespread efforts to collect large amount of data in structured format, only a small fraction of structured data overlaps with that documented in text (Seinen et al., 2025).
-Furthermore, empirically we observe that predictions are considerably better when text is integrated (Kline et al., 2022).
-A complementary data stream in healthcare comes from the textual data that providers use in documenting their encounters.
-Likely the most comprehensive part are the case reports and discharge summaries, which are a sort of “post-mortem”, retrospective discussion of the relevant factors for decision making and observed progression.
-However, without further effort, we are unable to leverage this data stream for forecasting because the times that are documented are interwoven in the narrative passage.
-To construct and learn from these patient trajectories, this work investigates the development of a forecasting pipeline for textual time series from retrospective analysis of clinical trajectories from case reports.
-Using the framework in Wang and Weiss (2025), we leverage extracted temporally located clinical findings, namely lists of (event, time) tuples that represent the patient encounter.
-The mapping from case report to time-localized events allows us to make anytime forecasts by masking those events that occur after the requested forecast time.
-Without this annotation process, it would be unclear which findings could be used so as to avoid “causal leakage” from events occurring in the future.
-Our approach differs from other efforts in the NLP and clinical informatics community and is driven by our particular emphasis on high-performance forecasting and temporally-aware risk characterization.
-For example, our definition of clinical finding is related to clinical entity and clinical concept definitions from the field of named entity recognition.
-However, we avoid mapping to a discrete set of entities to increase the availability of contextual information for forecasting, and thus require an alternative annotation guideline (Uzuner et al., 2011).
-With respect to our temporal annotations, our approach differs from the classical temporal relations task, which, while capturing relative ordering and overlap of events, makes it difficult to perform risk calculations and multimodal alignments that are functions of time (Kohane, 1987; Zhou and Hripcsak, 2007; Sun et al., 2013).
+Introduction
+Healthcare disparities persist globally, with preventable deaths from conditions like sepsis disproportionately affecting underserved populations who often present to under-resourced facilities with limited access to specialized expertise.
+In these settings, much of the critical diagnostic and prognostic information exists only in unstructured clinical narratives—case reports, discharge summaries, and progress notes—as comprehensive structured data infrastructure may be unavailable (Anzalone et al.
+2025; Seinen et al.
+2025).
+While machine learning approaches have shown promise on structured data, with recent work demonstrating that incorporating text alongside structured inputs significantly improves predictive performance (Kline et al.
+2022), large language models struggle with clinical risk estimation, especially in patient-facing scenarios (Wong et al.
+2025).
+This gap highlights the need for automated risk forecasting from textual clinical records that could democratize access to expert-level clinical reasoning, particularly in resource-constrained environments where timely specialist consultation is not available.
+The challenge of LLM risk estimation motivates a structured treatment of risk forecasting and survival modeling from unstructured narrative sources.
+Among textual sources, retrospective case reports serve as a rich training ground for developing temporally-aware AI systems that can eventually be deployed on real-time clinical notes.
+Case reports provide holistic accounts of clinical trajectories with explicit temporal reasoning—exactly the type of structured thinking needed for real-time clinical decision support.
+However, their narrative structure interleaves temporally unordered information, making them difficult to apply in forecasting tasks.
+Without identifying the time of occurrence for each event, models are prone to causal leakage—using information not available at the time of prediction.
+Solving this temporal reasoning challenge in case reports establishes the foundation for deployment in live clinical environments where such capabilities could improve patient outcomes.
 Figure 1: Example case report (left) with temporal representations as timelines (middle) and text-ordered, event-time tuples (right).
-We use LLMs to assist us in this task, so it is natural to question whether we need this extraction/annotation step at all, since a multitude of BERT-based and LLM models have consumed these, similar post-mortem reports, and more during training (Peng et al., 2020; Weber et al., 2024).
-From an architectural perspective, both encoder (BERT-style) and decoder (LLM-style) models will not perform the temporal masking necessary to encode leakage-free embeddings.
-For BERT models, which use full attention modules with masking, the probability that the random masking recovers a time-ordered representation is vanishingly small (tested empirically on the example timeline in Figure 1 (right) with default BERT masking probability of 0.15; probability of a time-ordered masking: 2/100,000).
-For decoder models, their causal attention mask means that embeddings are processed in text-order rather than time-order.
-Given that the concordance, the fraction of the time pairwise temporal ordering are correct, between text-ordered and time-ordered events is lower than 1 in most reports (c=0.23 in the example in Figure 1), we would expect suboptimal temporal reasoning capabilities.
-As we will see, we empirically find that these models do not forecast well without further training.
-If we can address these textual-temporal discrepancies, we expect to be able to substantially improve LLM capabilities at clinical forecasting, which could lead to downstream improvements in clinical temporal reasoning, risk estimation, causal reasoning, and decision making.
-1.1 Contributions
-  1. 1.
-    Annotation and Extraction Pipeline: Building on methods of Noroozizadeh and Weiss (2025) for extracting information from the PubMed Open Access corpus, we develop a pipeline that converts case reports into textual time series using both regex-based filtering and LLM-assisted extraction.
-The pipeline produces temporally localized (event, time) tuples, mitigating potential causal leakage by ensuring that only pre-cutoff events inform forecasts.
-  2. 2.
-    Comprehensive Model Comparison: We rigorously evaluate a broad suite of models across multiple forecasting tasks.
-Our study compares instruction-tuned decoder-based models (both via fine-tuning with MLP heads and prompting strategies) against encoder-based models with task-specific fine-tuning.
-Experiments on event occurrence prediction, temporal ordering, and survival analysis reveal complementary strengths: no single modeling approach consistently outperforms across all tasks, highlighting the importance of aligning model choice with the specific forecasting objective.
-  3. 3.
-    Analysis of Temporal versus Text Ordering: We perform a detailed investigation into the impact of ordering on performance by contrasting training on time-ordered versus text-ordered annotations.
-Our experiments demonstrate that preserving the clinical narrative’s natural order can enhance generalization on external datasets, while explicit time ordering can improve ranking performance under specific conditions.
-  4. 4.
-    Sensitivity Analysis of Temporal Masking: Through systematic dropout experiments, we analyze the effects of randomly masking portions of the clinical history on forecasting accuracy and event ordering.
-The findings reveal that while classification performance (F1 score) deteriorates with higher levels of masking, the concordance index remains largely stable—highlighting differing sensitivities of binary prediction and ranking tasks to historical context in our textual time series setup.
-Generalizable Insights about Machine Learning in the Context of Healthcare
-Extensive research has been in the clinical NLP community towards temporal understanding of clinical narratives (Kohane, 1987; Leeuwenberg and Moens, 2020; Cheng and Weiss, 2023).
-While much has focused on temporal relations, relative-time event tuples facilitate the use of longitudinal and survival models so long as they can use free-text inputs and outcomes (possibly via mappings to an ontology).
-Whereas the primary focus in this space is on the entity and temporal extraction tasks, downstream application studies remain infrequent.
-Our study seeks to close thie gap by applying forecasting methodologies to these data streams to characterize the performance of leading machine learning and NLP approaches.
-A better empirical understanding of the strength and weakness of these approaches will enable future work to advance improved textual time series methods.
-Our investigation also highlights that risk estimation and event anticipation is not optimally captured in LLMs (as we demonstrate in contrasts between fine-tuning and prompt-based strategies), suggesting the need for careful investigation of LLMs when applied to assessment of risk.
-2 Related Work
+One might ask whether large pretrained language models can already perform forecasting from clinical narratives, given their exposure to biomedical texts (Peng, Chen, and Lu 2020; Weber et al.
+2024).
+However, these models have architectural limitations for temporal reasoning tasks.
+Encoder-based models apply random masking that rarely yields temporally coherent representations, while decoder-style models mask text in linear sequence, modeling events in text order rather than time order.
+Our experiments demonstrate these limitations quantitatively–for example, using the timeline in Figure 1, the probability of recovering a valid time-ordered mask is $0.002\%$, and decoder models achieve low temporal concordance (e.g., $c=0.23$), leading to suboptimal forecasting accuracy without task-specific restructuring.
+To address these temporal reasoning challenges, we develop a forecasting framework that transforms free-text narratives into textual time series—sequences of (event, time) tuples and makes explicit the temporal structure of patient trajectories.
+Building on recent annotation work (Wang and Weiss 2025), we propose a hybrid pipeline combining rule-based heuristics and LLM-assisted extraction to recover temporally localized clinical findings, validated on real clinical notes with expert verification.
+Our work builds on extensive research in clinical NLP and temporal reasoning (Kohane 1987; Leeuwenberg and Moens 2020; Cheng and Weiss 2023), but shifts the emphasis from extraction to downstream clinical forecasting.
+While prior efforts have focused on entity recognition and temporal relation extraction (Zhou and Hripcsak 2007; Uzuner et al.
+2011; Sun, Rumshisky, and Uzuner 2013), few have systematically assessed how such representations influence downstream tasks such as risk prediction and survival modeling.
+We bridge this gap by formalizing textual time series representations and evaluating their effectiveness for temporally grounded clinical prediction.
+We empirically assess this framework across multiple tasks and modeling strategies, benchmarking encoder and decoder models—both fine-tuned and prompted versions—on three forecasting tasks: event prediction, survival analysis, and temporal ordering of future clinical events.
+Our results show no single model dominates across all tasks, highlighting the importance of aligning model architecture with forecasting objectives.
+We further examine how input ordering (time vs.
+narrative) impacts generalization and analyze the sensitivity of models to missing historical context through systematic masking experiments.
+In summary, this paper makes the following contributions: (i) Annotation and Extraction Pipeline: Building on prior methods for extracting information from the PubMed Open Access corpus (Noroozizadeh and Weiss 2025), this work introduces a pipeline that transforms case reports into textual time series via regular expression filtering and LLM-assisted extraction, yielding temporally anchored (event, time) tuples while minimizing causal leakage by restricting input to past events; (ii) Comprehensive Model Comparison: We conduct a comprehensive evaluation of encoder- and instruction-tuned decoder-based models—using both fine-tuned MLP heads and prompting—across event prediction, temporal ordering, and survival analysis tasks.
+Results show no single model excels universally, underscoring the need to align model choice with the specific forecasting objective; (iii) Temporal- versus Text- Ordering: We investigate the impact of annotation order by comparing training on time-ordered versus text-ordered data.
+Results show that preserving the narrative’s natural order can improve generalization to external datasets, while time-based ordering can enhance ranking performance; (iv) Sensitivity Analysis of Temporal Masking: We conduct systematic dropout experiments by randomly masking parts of the clinical history to assess their impact on forecasting and event ordering.
+While higher masking levels reduce classification performance (F1), the concordance index remains largely stable, highlighting differing sensitivities of binary prediction and ranking tasks to historical context in textual time series; (v) Methodological Framework for Temporal Clinical Forecasting: Our approach demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks, providing a replicable methodology that could be adapted to other clinical and non-clinical text sources.
+Related Works
 Temporal Information Extraction from Clinical Text:
-Extracting timelines from clinical narratives has been a longstanding challenge in biomedical NLP.
-Early efforts such as the i2b2 2012 challenge provided annotated datasets for temporal relation extraction, focusing on identifying and ordering events in discharge summaries (Uzuner et al., 2011).
-Follow-up work constructed timelines from these narratives by linking clinical events to timestamps or temporal expressions (Leeuwenberg and Moens, 2020; Frattallone-Llado et al., 2024).
-These methods often relied on pre-defined event spans and excerpts, rather than full-document processing.
-In contrast, following Noroozizadeh and Weiss (2025), we assign timepoints to clinical findings from full-length case reports, allowing for finer temporal resolution.
-While some works integrate structured EHR data to anchor timelines, our approach focuses on timeline construction and forecasting using text alone, which is essential for corpora such as the PubMed Open-Access Subset, where no structured metadata is available.
-Our annotation strategy enables direct supervision of event-time alignment, overcoming the limitations of earlier span-only annotations.
-This aligns with the broader need in sepsis and critical care research to incorporate temporality into clinical phenotyping (Johnson et al., 2018; Kamran et al., 2024; Seinen et al., 2025), especially given the high missingness in structured data (Johnson et al., 2016, 2023).
+Extracting timelines from clinical narratives is a challenging biomedical NLP task.
+The i2b2 2012 challenge introduced annotations datasets for temporal relation extraction from discharge summaries (Uzuner et al.
+2011).
+Subsequent methods linked clinical events to timestamps or temporal expressions (Leeuwenberg and Moens 2020; Frattallone-Llado et al.
+2024), typically using pre-defined event spans.
+We adopt the approach of Noroozizadeh and Weiss (2025) to assign timepoints to findings from full-length case reports, enabling finer temporal resolution.
+Unlike methods using structured EHR data, we focus solely on text—crucial for sources like PubMed that lack structured metadata.
+By directly supervising event-time alignment, we overcome limitations of span-based annotations (Rosenbloom et al.
+2011).
+This aligns with growing emphasis on temporality in sepsis and critical care phenotyping (Johnson et al.
+2018; Kamran et al.
+2024), particularly given high missingness rates in structured data (Johnson et al.
+2023; Seinen et al.
+2025).
 Predictive Modeling with Clinical Text:
-Incorporating unstructured text (clinical notes, narratives) into predictive models has been explored in various contexts, primarily for outcomes like mortality or readmission.
-Notably, ClinicalBERT and related transformer models pre-trained on medical notes have achieved state-of-the-art results in predicting hospital readmission and mortality (Huang et al., 2019; Gu et al., 2021).
-These works highlight that text contains complementary information to structured data, for example, documenting patient symptoms or social factors not in coded form.
-However, directly modeling text is challenging due to noise and length; researchers often rely on the full text or bag-of-words representations, which lose temporal granularity.
-Our approach differs by preserving the chronological sequence of events mentioned in text, effectively treating narrative data as a time series.
-This connects to recent studies on re-purposing LLMs for traditional time-series forecasting tasks: techniques like Time-LLM have shown that language models can be “re-programmed” to handle numeric sequences by mapping time points to token sequences (Jin et al., 2023).
-In the clinical domain, our work is the first to apply language models to text-derived time series, bridging free-text analysis with time-aware prediction.
-We also relate to prior sepsis prediction models.
-Traditional models use vital sign trends or clinical scores (e.g., SOFA, SAPS) to predict sepsis onset or mortality (Hou et al., 2020; Noroozizadeh et al., 2023).
-Unlike these, our method relies solely on narrative descriptions rather than real-time vitals, and is thus complementary: it demonstrates how knowledge embedded in case reports (which often include the clinicians’ interpretations and context) can be harnessed for prediction.
-This could be especially useful in settings where detailed narratives are available (such as case report literature or perhaps patient histories) but structured data is sparse.
+Clinical text has been used for outcome prediction tasks like mortality and readmission, with models like ClinicalBERT achieving high levels of performance on EHR notes (Huang, Altosaar, and Ranganath 2019; Gu et al.
+2021).
+Text captures complementary information—symptoms and social factors—not found in structured codes.
+However, traditional approaches using full-text or bag-of-words representations obscure temporal dynamics.
+Our approach preserves event order within narratives, treating data as time series.
+This aligns with recent efforts adapting LLMs for time series tasks, such as Time-LLM mapping numeric sequences to tokens (Jin et al.
+2023).
+To our knowledge, this is the first work applying LLMs to text-derived time series in clinical domains.
+While prior sepsis prediction models rely on vitals or scores like SOFA and SAPS (Hou et al.
+2020; Noroozizadeh, Weiss, and Chen 2023), our method is complementary—leveraging narrative descriptions that include clinician interpretations and context.
+This is valuable when case reports or patient histories are available but structured data is sparse or unavailable.
 Large Language Models in Healthcare:
-The advent of powerful LLMs like GPT-3 and GPT-4 has sparked interest in various healthcare applications, from medical question-answering to clinical note summarization.
-These models, trained on massive text corpora, contain a wealth of medical knowledge and the ability to perform reasoning to some extent.
-For instance, GPT-4 has been tested on medical board exam questions with impressive results, and pilot studies show LLMs can abstract patient records for quality reporting (Boussina et al., 2024).
-Nevertheless, using LLMs for forecasting in healthcare is relatively uncharted territory.
-A limitation of out-of-the-box LLMs is their tendency to generate plausible-sounding but unverified content (hallucinations) and their lack of alignment with specific clinical data distributions.
-Our benchmark explicitly evaluates LLMs through prompting (zero-shot and few-shot) to see if they can ”forecast” events when given a patient’s story so far.
-We also compare that against fine-tuned models that are directly trained on the task, to quantify the gap between general-purpose and task-specific modeling in this context.
-This comparison adds to the growing body of evidence that, while large models have broad knowledge, domain-specific fine-tuning or structuring of input is often necessary for reliable performance in clinical tasks (Huang et al., 2019; Gu et al., 2021).
-Some recent analyses suggest that medical-domain fine-tuning may not consistently improve performance over foundation models (Jeong et al., 2024), further motivating the need for task- and input-specific evaluation setups like ours.
-3 Methods
- [^1] We will be releasing our code upon publication.
-We outline the task specification, define key terms, describe the extraction and annotation process, and present the evaluation criteria (Figure 2).
-Figure 2: Flowchart of the T2S2 forecasting analysis pipeline
-3.1 Dataset extraction
+Recent LLMs such as GPT-3 and GPT-4 have enabled applications from medical QA to note summarization.
+These models encode substantial medical knowledge and reasoning capacity—GPT-4 demonstrates strong performance on board exams and can support patient record abstraction for quality reporting (Boussina et al.
+2024).
+However, their use in clinical forecasting remains largely unexplored.
+Our benchmark tests LLMs via zero- and few-shot prompting to assess clinical event forecasting from patient narratives, comparing results to fine-tuned models to quantify the gap between general-purpose and task-specific modeling.
+This contributes evidence that while LLMs hold broad medical knowledge, fine-tuning and structured inputs are often required for clinical reliability (Huang, Altosaar, and Ranganath 2019; Gu et al.
+2021).
+Some studies suggest domain-specific tuning does not always improve over foundation models (Jeong et al.
+2024), reinforcing the need for task-specific evaluations like ours.
+Methods [^1] Our code is available at https://github.com/Shahriarnz14/Textual-Time-Series-Forecasting
+We next describe our dataset construction and annotation pipeline, define the forecasting and survival prediction tasks, detail the modeling approaches evaluated, and present the sensitivity analyses used to assess robustness (Figure 2).
+Figure 2: T2S2 forecasting analysis pipeline
+Dataset Extraction
 We utilized the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis.
-Building on the method described in Wang and Weiss (2025), we extract the body of each case by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus.
-To maintain relevance, we include only those documents that contain case-insensitive matches to the regular expressions (case report|case presenta) and year-?
-?old.
-Prior work has shown that this regex-based approach demonstrated notably higher specificity compared to PubMed metadata on a random sample of 20 case reports Noroozizadeh and Weiss (2025).
-As demonstrated in prior work Noroozizadeh and Weiss (2025), we applied additional filtering to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care.
+Building on the method described in prior work (Wang and Weiss 2025), the body of each case was extracted by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus.
+To maintain relevance, we include only documents containing case-insensitive matches to (case report|case presenta) and year-?
+?old, an approach shown to have higher specificity than PubMed metadata (Noroozizadeh and Weiss 2025).
+As demonstrated in prior work of (Noroozizadeh and Weiss 2025), additional filtering was applied to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care.
 We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender.
@@
 This multi-step process yielded a total of 2319 Sepsis-3 case reports.
-From this set, we randomly selected two subsets for evaluation: a group of 10 reports (sepsis-10), which underwent expert clinical annotation to serve as a gold standard, and a larger group of 100 reports (sepsis-100) used for broader testing (Figure 2).
-3.2 Textual time-series annotation
-As established in previous work Noroozizadeh and Weiss (2025), we aimed to construct a textual time series case report corpus of Sepsis-3 patients from the PMOA sepsis case reports.
+From this set, we randomly selected two subsets for evaluation: a group of 10 reports (sepsis-10), which underwent expert clinical annotation to serve as a gold standard, and a larger group of 100 reports (sepsis-100) as a bronze standard for broader testing based on an alternative annotator (O1-preview; Figure 2).
+Textual Time Series Annotation
+Following Noroozizadeh and Weiss (2025), a textual time series case report corpus of Sepsis-3 patients was constructed from the PMOA sepsis reports.
 A textual time series refers to a sequence of clinical findings, each paired with a timestamp (either absolute or relative to the case presentation time), corresponding to an individual patient.
 Here, a clinical finding is defined as a free-text expression describing an entity that pertains to or may impact a person’s health.
-A textual time-series is different from conventional approaches to clinical concept annotation (e.g., Sun et al.
-(2013); Uzuner et al.
-(2011)) by extending the span of extracted text to better capture the specificity and contextual meaning of each clinical finding.
+A textual time series differs from conventional approaches to clinical concept annotation (e.g., (Sun, Rumshisky, and Uzuner 2013; Uzuner et al.
+2011)) in its extension of the text span to better capture the specificity and contextual meaning of each clinical finding.
 In our work, the interpretation of a clinical finding differs from the i2b2 concept guidelines in the following respects : (1) Clinical findings are not restricted to single prepositional phrases following a markable entity.
@@
 We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports.
-We refer to this as Textual Time-Series for Sepis (T2S2) The exact LLM extraction prompt can be found in Appendix E.
-Our dataset was split into training (n = 1795), testing T2S2-test (n = 244) and the 2 external validation sets sepsis 10 (n = 10) and sepsis 100 (n = 90).
-The training set was further split into 80% for model training and 20% for model validation (e.g.
-early stopping).
-3.3 Task Definition
-Using the timelines extracted from the annotations, we defined two primary forecasting tasks and a survival analysis task:
-Event Occurrence Prediction:
+We refer to this as Textual Time-Series for Sepsis (T2S2).
+The exact LLM extraction prompt can be found in Appendix E.
+Our dataset was split into training (n = 1,795; 80% train, 20% validation), testing T2S2-test (n = 244), and two external validation sets: sepsis-10 (n = 10) and sepsis-100 (n = 90).
+Forecasting Tasks
+Using the timelines extracted from the annotations, we defined two primary forecasting tasks and a survival analysis task as follows:
+Event Occurrence Prediction.
 Given a prefix of the clinical timeline (all events up to a certain time $t$), the model is tasked with predicting whether each of the immediate next $k$ events occurs within a specified time horizon.
@@
 Evaluation is based on precision, recall, and F1 score, averaged across event positions for each time horizon.
-Temporal Ordering Prediction:
+Temporal Ordering Prediction.
 This task assesses the model’s ability to reconstruct the correct sequence of future events.
@@
 This tests whether the model can infer temporal progression from unordered event content alone.
-Survival Analysis for Mortality Time:
+Survival Analysis for Mortality Time.
 We include a classical survival analysis task to model time until death.
-Many case reports specify whether the patient died and, if so, when (e.g., “the patient died on hospital day 10”), which enables us to define a time-to-event outcome.
+Many case reports specify whether the patient died and, if so, when (e.g., ”the patient died on hospital day 10”), which enables us to define a time-to-event outcome.
 For this task, we evaluate models at predefined cut-off times – specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week) – and use the extracted event history up to each cut-off as input.
 A survival model is trained to predict the probability of survival over time beyond each cut-off.
-We evaluate performance using the time-dependent concordance index to assess how well the predicted survival times align with actual outcomes.
-3.4 Modeling Approaches
-To evaluate performance on the event forecasting and survival prediction tasks, we implement three modeling paradigms: (i) large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), and (iii) BERT-style encoder models with task-specific fine-tuned heads.
-In Appendix C and  D, we provide detailed explanations of how we approach the event occurrence prediction, temporal ordering prediction, and survival analysis tasks using encoder-only models.
-Fine-tuned LLMs:
-We apply instruction-tuned decoder-only models from the LLaMA and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants.
+We evaluate using the time-dependent concordance index to measure alignment of the predicted survival times with actual outcomes.
+Modeling Approaches
+To evaluate performance on the event forecasting and survival prediction tasks, we implement five modeling paradigms: (i) decoder-only large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), (iii) encoder-only models with task-specific fine-tuned heads, (iv) encoder-masking models with fine-tuning, and (v) encoder-masking models in zero-shot settings.
+In Appendix C and  D, we provide detailed explanations of implementation details.
+(i) Fine-tuned LLMs.
+We apply instruction-tuned decoder-only models from the LLaMA and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants (Grattafiori et al.
+2024; Liu et al.
+2024).
+We also include open-source models with documented training corpora (OLMO-32B-Instruct (OLMo et al.
+2024), RedPajama-INCITE-7B-Instruct (Weber et al.
+2024)) to address potential data leakage concerns–notably, RedPajama explicitly excludes PubMed from its training data.
+Additionally, we evaluate a medically fine-tuned decoder model (MediPhi-PubMed (Corbeil et al.
+2025)) to assess domain-specific benefits.
 Each model is paired with a lightweight multilayer perceptron (MLP) head trained for classification or ranking, depending on the task.
@@
 Training is conducted using cross-entropy loss for classification and pairwise ranking loss for ordering.
-Prompted LLMs:
+(ii) Prompted LLMs.
 In the zero- or few-shot setting, we use the same LLM architectures as above, but without any fine-tuning.
@@
 Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format.
-Prompt designs are customized for each task and are detailed in Appendix F.
+Prompt designs are customized for each task and are detailed in Appendix E.
 Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
-Fine-tuned Encoder-Only Models:
+(iii) Fine-tuned Encoder-Only Models.
 We also evaluate a range of encoder-based models trained end-to-end on each task.
-Architectures include BERT-base-uncased, RoBERTa-base, DeBERTa-v3-small, ModernBERT-small, and ModernBERT-large.
+Architectures include general-purpose models (BERT-base-uncased (Devlin et al.
+2019), RoBERTa-base (Liu et al.
+2019), DeBERTa-v3-small (He, Gao, and Chen 2021), ModernBERT-base/large (Warner et al.
+2024)) and biomedically-pretrained variants (BioClinical-ModernBERT-base/large (Sounack et al.
+2025)).
 For each model, we append a task-specific MLP head for event occurrence or ordering prediction.
@@
 These models are trained using standard supervised learning objectives and evaluated on the same metrics as other methods: F1 score for event occurrence, and pairwise concordance for temporal ordering.
-For survival analysis, the final hidden states are passed into a time-to-event prediction head and evaluated using the time-dependent concordance index.
-4 Results
+For survival analysis, the final hidden states are passed into a time-to-event prediction head and evaluated using the time-dependent concordance index (Antolini, Boracchi, and Biganzoli 2005).
+(iv, v) Encoder-Masking Models.
+We adapt masked language modeling (MLM) for our temporal reasoning tasks.
+For event occurrence, models predict masked tokens (“yes”/“no”) in prompts such as “Will [event] happen within 24 hours?
+[MASK]”.
+For temporal ordering, models predict “before”/“after” tokens in prompts comparing event pairs.
+We evaluate both fine-tuned variants (trained on task-specific objectives) and zero-shot variants (using pretrained MLM capabilities without additional training).
+This approach leverages the bidirectional context of encoder models while maintaining interpretable predictions through constrained vocabulary.
+Sensitivity Analyses
+We conduct sensitivity analyses to examine how the temporal structure and completeness of input sequences affect forecasting performance.
+These analyses probe the robustness of models to changes in event ordering and to varying levels of missing historical information:
+Time-ordering Strategies.
+We evaluate two strategies for ordering clinical events within textual time series inputs.
+In the text-ordered setting, events are presented in the order they appear in the original case report narratives, preserving the narrative structure produced during extraction.
+In the time-ordered setting, events are sorted chronologically by their documented occurrence time, enforcing strict temporal alignment.
+This comparison isolates the effect of narrative sequencing versus explicit temporal ordering on model performance.
+For text-ordered inputs, to prevent causal leakage, events occurring after the forecast cutoff time $t$ are masked for encoder-based models and omitted for decoder-based models.
+Timestep Dropout.
+To assess robustness to incomplete patient histories, we introduce a timestep dropout rate (TDR), defined as the proportion of input timesteps randomly removed prior to inference.
+This procedure simulates varying degrees of missing clinical documentation that may arise in deployment settings.
+We vary TDR from 0% (full history) to 90% (severely truncated history) in increments, masking events independently and uniformly at random.
+This setup enables controlled evaluation of model performance under partial information, probing the extent to which predictive accuracy and event ordering depend on the completeness of historical context.
+Together, these analyses characterize model robustness to disruptions in temporal coherence and input completeness – two key challenges in real-world deployment scenarios.
 Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
-All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1.
-Bold values indicate best in category amongst all models.
-T2S2 (DeepSeek R1) sepsis-10 sepsis-100
-F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
+All models are trained/fine-tuned on time-ordered annotations from either DeepSeek-R1 or Llama 3.3-70B.
+Bold values indicate best in category within each column group (refer to Tables A1 and A2 in Appendix A for detailed performance statistics on all models).
+Performance statistics for all models are presented in A1 and A2, respectively.
+Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively.
+c10/100: c-index for sepsis-10 and sepsis-100 respectively.
+T2S2 (DeepSeek-R1) T2S2 (Llama 3.3-70B)
+F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100 F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
 LLM-MLP head
-DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
+DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
+OLMO 32B Instruct 0.000 0.411 0.688 0.621 0.273 0.561 0.432 0.593 0.190 0.315 0.764 0.604 0.238 0.551 0.315 0.596
+RedPajama-INCITE 7B Instruct 0.000 0.352 0.651 0.618 0.282 0.572 0.471 0.594
+MediPhi-PubMed 0.000 0.268 0.751 0.626 0.169 0.561 0.210 0.597 0.000 0.507 0.801 0.623 0.277 0.564 0.534 0.600
 Prompting
-L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452
+L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452 0.064 0.460 0.729 0.313 0.480
+OLMO 32B few-shot ordinal 0.013 0.315 0.651 0.247 0.422 0.046 0.413 0.647 0.300 0.424
 Encoder-MLP head
-ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
+ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
+Bioclinical ModernBERT-base 0.264 0.559 0.879 0.677 0.395 0.598 0.635 0.610 0.290 0.653 0.902 0.650 0.449 0.618 0.662 0.627
 Encoder-masking-fine-tuned
-ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
+ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
+Bioclinical ModernBERT-base 0.166 0.450 0.692 0.676 0.336 0.604 0.499 0.655 0.129 0.481 0.733 0.653 0.331 0.615 0.475 0.658
 Encoder-masking-zeroshot
-ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
-ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
-Table 2: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
-All models are trained/fine-tuned on time-ordered annotations from Llama 3.3-70B.
-Bold values indicate best in category amongst all models.
-T2S2 (Llama 3.3) sepsis-10 sepsis-100
-F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
-LLM-MLP head
-DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
-Prompting
-L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
-Encoder-MLP head
-ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
-Encoder-masking-fine-tuned
-ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
-Encoder-masking-zeroshot
-ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
-4.1 Event forecast within next 24 hours: F1 performance
-Our results demonstrate a clear advantage of encoder-based models over LLMs in event forecasting, for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2) and particularly in short-term and medium-term prediction windows.
-The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A.
-Encoder models, particularly those with a fine-tuned MLP head, achieve substantially higher F1 scores than the LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
-Among encoder models, different training strategies yield distinct performance trends.
+ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
+Bioclinical ModernBERT-base 0.246 0.458 0.614 0.498 0.301 0.509 0.457 0.498 0.197 0.460 0.533 0.499 0.299 0.520 0.465 0.499
+Results
+In this section, we present results for our three main evaluation settings: event forecast within a subsequent time window, temporal ordering of forecasted events, and survival analysis.
+We then provide sensitivity analyses on the forecasting tasks, examining the effects of temporal ordering and historical context availability.
+Forecasting Tasks
+Event forecast within next 24 hours: F1 performance.
+Our results show that encoder-based models outperform decoder-based LLMs in event forecasting across both DeepSeek-R1 and Llama-3.3-70B annotations (Table 1).
+Forecasting performance across all models is shown in Tables A1 and A2 (Appendix A).
+Encoder models, especially those with a fine-tuned MLP head, achieve substantially higher F1 scores than decoder LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
+Among decoder models, we observe varied performance patterns.
+RedPajama-INCITE-7B-Instruct, with documented training that excludes PubMed, achieves 0.352 F1 at 24h—better than some decoder models like MediPhi-PubMed (0.268) and within the range of decoder performance, though below top performers like DeepSeek-Llama-3.3-70B (0.482).
+OLMO-32B-Instruct performs competitively (0.411), suggesting that open models without biomedical pretraining can still achieve reasonable results for temporal forecasting tasks.
+Among encoder models, both general-purpose and biomedically-pretrained variants show strong performance.
+While ModernBERT-base achieves slightly higher F1 scores on internal test sets (0.576 vs 0.559), BioClinical-ModernBERT-base demonstrates superior concordance (0.677 vs 0.646 for DeepSeek-R1) and notably better generalization to external datasets, with F1 scores of 0.635 and 0.662 on sepsis-100 compared to 0.607 and 0.626 for the general-purpose variant.
+This suggests that while both ModernBERT architectures are effective, biomedical pretraining particularly enhances temporal reasoning and cross-dataset generalization.
+Different training strategies yield distinct performance patterns.
 The MLP head fine-tuning approach consistently outperforms both fine-tuned masking and zero-shot masking models, particularly in long-horizon forecasting.
-ModernBERT-base achieves the highest overall F1 scores across all forecasting horizons, particularly at 168 hours, where it outperforms all other models.
-ModernBERT-large follows closely but does not consistently surpass its base variant, suggesting diminishing returns with increased model size.
-Meanwhile, zero-shot masking models, particularly standard BERT and RoBERTa, fail to make meaningful predictions, often yielding near-zero F1 scores at the 1-hour mark.
-However, DeBERTa-small and ModernBERT-base exhibit relatively better zero-shot performance, indicating that certain architectures generalize better without task-specific training.
-Performance trends also vary across different datasets.
-Across all models, F1 scores improve as the forecasting window increases, with 1-hour predictions being the most challenging and 168-hour predictions yielding the highest scores.
-This suggests that longer-term trends are more predictable, whereas immediate event forecasting remains difficult, possibly due to the sparsity and volatility of short-term event occurrences.
-Comparing internal and external datasets, models perform worse on external validation sets (sepsis-10 and sepsis-100) than on the internal T2S2 test set, with a particularly noticeable drop in F1 at the 1-day horizon.
-This indicates potential generalization challenges, as models trained on internal data struggle to adapt to distributional shifts in clinical settings.
-Notably, ModernBERT models maintain a relative performance edge even in external datasets, reinforcing their robustness across different domains.
-4.2 Temporal ordering of forecasted events: Concordance
-In addition to event occurrence prediction, models are evaluated on their ability to correctly rank the order of upcoming events, measured through concordance (c-index) for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2).
-The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A.
-Similar to F1 trends, encoder-based models significantly outperform LLMs, with LLMs failing to achieve reliable ordinal rankings even when fine-tuned or prompted.
-Within encoder models, different training approaches show distinct performance characteristics: MLP-head fine-tuned models dominate in event occurrence classification (F1) but do not always yield the best concordance values.
-Instead, fine-tuned masking models, particularly ModernBERT-large, achieve the highest concordance scores, suggesting that this approach is better suited for ordering rather than binary classification.
-Performance differences across datasets also highlight the challenges of generalization.
-While internal (T2S2) test results show relatively strong concordance for fine-tuned models, c-index values drop considerably on external datasets, particularly for models that are not ModernBERT-based.
-Among the best-performing models, ModernBERT-large (fine-tuned masking) achieves the highest c-index across both internal and external datasets, reinforcing its strength in capturing event order relationships.
-However, zero-shot masking models continue to struggle, mirroring their poor F1 performance.
-Additionally, while MLP-head fine-tuned models maintain high F1 scores, they exhibit slightly lower concordance than fine-tuned masking models, particularly at the 168-hour horizon.
-This suggests that different optimization strategies may be required depending on whether the task prioritizes strict classification accuracy or relative event ordering.
-Overall, the results confirm that ModernBERT architectures are the strongest performers across both F1 and concordance evaluations, with fine-tuned MLP-head models excelling in binary prediction tasks and fine-tuned masking models achieving superior ranking performance.
-However, all models face challenges in generalizing to external datasets, emphasizing the need for domain adaptation techniques to improve real-world forecasting robustness.
-4.3 Survival analysis
-Model Performance across Timepoints and Cohorts.
-Table 3 presents the time-dependent concordance index (concordance_td) for all models across three evaluation cohorts—t2s2, Sepsis10, and Sepsis100—at three observation windows (0h, 24h, and 168h).
-For each cohort and timepoint, we report performance from the best-fitting survival model (selected based on validation concordance).
-Across cohorts, we observe that providing additional observation time (e.g., from 0h to 168h) does not consistently improve predictive performance.
-In the t2s2 cohort, several models achieve their peak concordance at the earliest or intermediate timepoints (e.g., DeepSeek-R1-Llama-70B peaks at 0h, while Llama-3.3-70B-Instruct peaks at 24h), with some degradation at 168h.
-In Sepsis100, certain models benefit from longer observation (e.g., DeepSeek-R1-Distill-Llama-8B improves steadily across timepoints), but others decline or plateau.
-The Sepsis10 cohort exhibits high and uniform performance across models, particularly after 24h, suggesting potential ceiling effects or limited evaluation variability in that subset.
-Instruction-tuned LLMs such as Llama-3.3-70B-Instruct, DeepSeek-R1-Distill- Llama-70B, and Llama-3.1-8B-Instruct consistently outperform conventional transformer baselines like bert-base-uncased or roberta-base.
-This trend is particularly notable in the t2s2 and Sepsis100 cohorts, where LLMs demonstrate stronger early signal detection at 0h and 24h.
-These results suggest that large-scale pretraining and instruction tuning confer advantages in extracting actionable clinical representations, even with limited temporal context.
-Taken together, these findings highlight that strong early performance does not require extended observation windows and that pretrained instruction-following models can offer substantial benefits for survival prediction in real-world, noisy clinical text scenarios.
+While ModernBERT-base achieves strong F1 scores, BioClinical-ModernBERT-base demonstrates the best overall performance when considering both accuracy and generalization.
+Zero-shot masking models, particularly standard BERT and RoBERTa, fail to make meaningful predictions at the 1-hour mark.
+However, ModernBERT variants exhibit relatively better zero-shot performance, with BioClinical-ModernBERT-base achieving notably higher scores in zero-shot settings (0.246 F1 at 1h), indicating that biomedical pretraining improves out-of-the-box temporal reasoning.
+Performance trends remain consistent across forecasting windows.
+F1 scores improve as the window increases, with 1-hour predictions being most challenging and 168-hour predictions yielding highest scores.
+Models perform worse on external validation sets than internal T2S2 test sets, though BioClinical-ModernBERT models show the smallest performance degradation, maintaining their advantage in cross-dataset generalization.
+Temporal ordering of forecasted events: concordance.
+Encoder models consistently outperform decoders in correctly ranking the order of upcoming events, as measured by concordance (c-index; Table 1; complete results in Tables A1 and A2 for DeepSeek-R1 and Llama-3.3-70B annotations, respectively, with full details in Appendix A.
+Decoder models with potential biomedical exposure achieve modest concordance (0.618-0.632), while RedPajama-INCITE-7B-Instruct without PubMed data in its training corpus performs similarly, suggesting domain-specific knowledge is less critical here than for event prediction.
+The comparable performance of RedPajama-INCITE-7B-Instruct to other general models (with access to PubMed in their pretraining) also provides evidence that the T2S2 prediction task is distinct from raw PubMed text and assuages causal leakage concerns.
+Biomedical pretraining in encoder models also demonstrates clear advantages, with BioClinical-ModernBERT-base attaining the highest concordance (0.677 with MLP head, 0.676 fine-tuned masking) and robust performance across datasets.
+MLP-head training yields superior F1 scores, while fine-tuned masking excels in concordance.
+BioClinical-ModernBERT-base maintains strong generalization on external datasets (c-index >0.60), suggesting that biomedical pretraining provides stability in temporal reasoning across dataset shifts.
+Survival analysis: model performance across timepoints and cohorts.
+Table 2 presents time-dependent concordance across evaluation cohorts (T2S2, sepsis-10, sepsis-100) at multiple observation windows (0h, 24h, 168h), selecting models based on validation concordance.
+General models, such as RedPajama-INCITE-7B-Instruct, show strong survival prediction capabilities even without biomedical exposure (0.76 concordance at 168h, sepsis-100).
+Meanwhile, biomedical encoders like BioClinical-ModernBERT-base consistently perform well, particularly early (0h).
+Predictive performance does not universally improve with more observation time.
+T2S2 cohort models often peak at earlier windows, while sepsis-10 experiences ceiling effects.
+Instruction-tuned decoders excel overall, yet biomedical encoders remain competitive, particularly at admission (0h).
+Survival prediction thus benefits from both larger decoder models and domain-specific pretraining, each excelling in different temporal contexts.
+Table 2: Survival analyses evaluated using time-dependent concordance index for each model across DeepSeek-R1 test set annotations (T2S2, Sepsis-10, Sepsis-100), evaluated at three observation time thresholds (0h, 24h, and 168h).
 Model T2S2 sepsis-10 sepsis-100
-0h 24h 168h 0h 24h 168h 0h 24h 168h
+Time Threshold 0h 24h 168h 0h 24h 168h 0h 24h 168h
 bert-base-uncased 0.60 0.61 0.53 1.00 1.00 1.00 0.58 0.62 0.58
@@
 ModernBERT-large 0.53 0.59 0.58 0.50 1.00 1.00 0.59 0.55 0.49
+Bioclinical ModernBERT-base 0.57 0.53 0.53 1.00 0.75 0.75 0.63 0.59 0.56
+Bioclinical ModernBERT-large 0.54 0.55 0.54 0.75 0.75 0.75 0.61 0.57 0.49
 DeepSeek-R1-Llama-70B 0.64 0.63 0.59 0.50 1.00 1.00 0.61 0.53 0.55
@@
 Llama-3.1-8B-Instruct 0.62 0.61 0.61 1.00 1.00 1.00 0.60 0.59 0.61
-Table 3: Concordance index (concordance_td) for each model across three test sets (t2s2, Sepsis10, Sepsis100), evaluated at three observation time thresholds (0h, 24h, 168h).
-4.4 Sensitivity analysis
-To better understand the robustness and design implications of our forecasting models, we conduct two complementary sensitivity analyses focused on the temporal structure and completeness of input sequences.
-First, we compare two strategies for ordering clinical events within textual time-series (TTS) inputs: text-ordered, where events follow their extraction sequence from the original case report narratives, and time-ordered, where events are chronologically sorted.
-This comparison assesses the relative benefits of preserving narrative flow versus enforcing temporal alignment.
-Second, we evaluate the impact of input sparsity by randomly masking portions of the clinical history, simulating scenarios with incomplete documentation or missing data.
-Together, these analyses shed light on the importance of chronological ordering and data completeness for both classification accuracy and temporal event ranking.
-4.4.1 Time-ordered vs text-ordered training
-We evaluated the impact of training on text-ordered versus time-ordered textual time-series (TTS) annotations.
-In the text-ordered condition, events in the textual time-series order follow the order in which they were extracted from the case reports using LLMs (DeepSeek-R1 and Llama-3.3-70B), while in the time-ordered condition, they are sorted chronologically in time.
-This comparison aimed to assess the impact of chronological versus narrative ordering of events on forecasting performance
-As shown in Table 4, time-ordered training generally leads to higher concordance (C-index), particularly for DeepSeek-R1 annotatiobs across both internal (T2S2) and external test datasets (sepsis-10 and sepsis-100).
-However, F1 scores are less consistently affected by the ordering strategy.
-In several cases—especially with Llama-3.3-70B annotations on external datasets, text-ordered training results in equal or better F1 scores, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order.
-These results highlight a trade-off between chronological alignment and narrative structure in optimizing forecasting performance.
-Table 4: Comparison of text-ordered and time-ordered training for Deep Seek R1 and Llama 3.3 annotations on F1 scores and C-index values.
-Dataset Metric DeepSeek-R1 Llama 3.3-70B
-Text-ordered Time-ordered Text-ordered Time-ordered
-T2S2 Test F1 (1hr) 0.290 0.306 0.311 0.257
-F1 (1day) 0.569 0.576 0.648 0.645
-F1 (7days) 0.876 0.877 0.898 0.903
-C-index 0.659 0.672 0.586 0.632
-Sepsis10 F1 (1day) 0.361 0.428 0.423 0.382
-C-index 0.598 0.576 0.526 0.595
-Sepsis100 F1 (1day) 0.615 0.607 0.662 0.616
-C-index 0.609 0.613 0.541 0.612
-4.4.2 Impact of Randomly Masking History on F1 and Concordance
-Further, we systematically evaluate how randomly omitting portions of clinical history impacts both forecasting accuracy and the model’s ability to correctly order clinical events.
-Towards this end, we simulate incomplete patient histories by randomly masking input time steps at varying rates.
-We refer to this as the timestep drop rate (TDR), ranging from 0% (no masking) to 90% (severely truncated history).
-We assess the effect of this masking on both classification performance (F1 score at 1 day) and ranking performance (concordance index) across three datasets—T2S2-DS, Sepsis-10, and Sepsis-100—using both DeepSeek-R1 (Table 5) and Llama-3.3-70B annotations (Table 8 in Appendix B).
-F1 scores decline progressively with increasing timestep dropout rate (TDR), with sharp reductions beyond 60% (Table 5).
-At TDR = 0%, the F1 scores for T2S2-DS, Sepsis-10, and Sepsis-100 are 0.576, 0.428, and 0.607, respectively.
-By TDR = 90%, these values fall to 0.423, 0.200, and 0.510, with Sepsis10 showing the steepest decline.
-Between 30% and 60% dropout, performance remains relatively stable for T2S2-DS and Sepsis100, while Sepsis10 exhibits a more pronounced drop at TDR = 60%.
-Similar trends are observed with Llama-3.3 annotations (Table 8 in Appendix B).
-In contrast, concordance (c-index) remains largely consistent across dropout levels.
-T2S2-DS (Table 5) reaches its highest concordance (0.678) at TDR = 60%, while Sepsis-10 and Sepsis-100 vary only marginally.
-This suggests that while masking temporal history substantially reduces classification accuracy, it has limited impact on event ranking performance.
-Comparable patterns are observed for the Llama-3.3 models (Table 8).
-Overall, DeepSeek models show notable degradation in F1 scores when more than 30% of history is removed—particularly for datasets sensitive to longer-term dependencies—whereas concordance scores remain relatively stable (Table 5 and Table 8).
-Table 5: Effect of Randomly Masking History (DeepSeek R1 annotations) on F1 and Concordance. TDR : Timestep drop rate
-TDR F1 (1 day) Concordance (c-index)
-T2S2-DS Sepsis-10 Sepsis-100 T2S2-DS Sepsis-10 Sepsis-100
-0% 0.576 0.428 0.607 0.672 0.576 0.613
-15% 0.572 0.415 0.604 0.667 0.564 0.616
-30% 0.548 0.357 0.591 0.672 0.542 0.619
-60% 0.542 0.310 0.588 0.678 0.578 0.616
-90% 0.423 0.200 0.510 0.661 0.561 0.614
-5 Discussion and Conclusion
+OLMO 32B Instruct 0.60 0.62 0.59 0.75 0.75 0.75 0.66 0.58 0.56
+RedPajama-INCITE 7B Instruct 0.56 0.60 0.61 0.50 1.00 1.00 0.61 0.53 0.76
+MediPhi-PubMed 0.54 0.62 0.55 1.00 1.00 1.00 0.62 0.58 0.61
+Sensitivity analysis
+Time-Ordered versus Text-Ordered Training.
+Table 3 shows time-ordering generally improves concordance, particularly for DeepSeek-R1 annotations, while F1 scores show mixed results.
+Text-ordering sometimes yields better external dataset performance with Llama-3.3-70B, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order.
+These results highlight a trade-off between temporal alignment and narrative coherence for different tasks.
+Table 3: Comparison of text-ordered and time-ordered training for DeepSeek-R1 and Llama 3.3-70B annotations on F1 scores and c-index values.
+F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
+T2S2 (DeepSeek-R1)
+Text-ordered 0.290 0.569 0.876 0.659 0.361 0.598 0.615 0.609
+Time-ordered 0.306 0.576 0.877 0.672 0.428 0.576 0.607 0.613
+T2S2 (Llama 3.3-70B)
+Text-ordered 0.311 0.648 0.898 0.586 0.423 0.526 0.662 0.541
+Time-ordered 0.257 0.645 0.903 0.632 0.382 0.595 0.616 0.612
+Impact of timestep dropout on F1 and concordance.
+Tables 4 and A3 (Appendix B) report the effect of timestep dropout on classification and ranking performance.
+F1 scores decrease monotonically with increasing TDR, with relatively mild degradation up to 30–60% dropout and sharper declines thereafter.
+For example, in the DeepSeek-R1 annotations, T2S2/sepsis-10/sepsis-100 drop from 0.576/0.428/0.607 at TDR = 0% to 0.423/0.200/0.510 at TDR = 90%, with sepsis-10 showing the largest relative decline.
+In contrast, concordance (c-index) remains stable across dropout levels (0.661–0.678), indicating that ranking-based evaluations are less sensitive to history truncation.
+These trends are consistent for Llama-3.3 annotations (Appendix B), suggesting that ranking objectives capture temporal relationships that are more robust to missing historical information than binary classification accuracy of predicting whether an events happens within a time window.
+Table 4: Effect of Randomly Masking History (DeepSeek-R1 annotations) on F1 and Concordance.
+Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
+0% 0.576 0.672 0.428 0.576 0.607 0.613
+15% 0.572 0.667 0.415 0.564 0.604 0.616
+30% 0.548 0.672 0.357 0.542 0.591 0.619
+60% 0.542 0.678 0.310 0.578 0.588 0.616
+90% 0.423 0.661 0.200 0.561 0.510 0.614
+Discussion
 Our findings highlight the superiority of encoder-based models over decoder-based LLMs for event forecasting, emphasizing the limitations of autoregressive models in structured prediction tasks.
-Despite prompting and fine-tuning, LLMs struggle with short- and medium-term forecasting, reinforcing the need for models optimized for sequence encoding rather than generative token prediction.
-Among encoder models, ModernBERT-base and ModernBERT-large consistently perform best, with MLP-head fine-tuning excelling in F1 scores and fine-tuned masking models achieving higher concordance (c-index), particularly in long-term forecasting.
-This suggests that classification and ranking tasks benefit from distinct optimization strategies, and a hybrid training approach may further enhance performance.
+Among encoder models, BioClinical-ModernBERT-base consistently performs best, achieving superior concordance (0.677 vs 0.646) and better external generalization despite slightly lower internal F1 scores than general-purpose ModernBERT-base.
+MLP-head fine-tuning excels in F1 scores while fine-tuned masking models achieve higher concordance, particularly in long-term forecasting.
+This suggests that classification and ranking tasks benefit from distinct optimization strategies.
+The consistent advantage of biomedically-pretrained models across tasks deserves special attention.
+While general-purpose models achieve strong internal performance, biomedical pretraining provides crucial benefits for real-world deployment: BioClinical-ModernBERT variants show the smallest performance degradation on external datasets and maintain higher zero-shot capabilities (0.246 F1 at 1h vs near-zero for standard BERT/RoBERTa).
+This suggests that domain-specific pretraining enhances not just accuracy but also robustness and generalization—critical factors when considering clinical deployment where distribution shifts are common.
 The gradual improvement of F1 scores over longer time horizons indicates that event patterns become more predictable over time, whereas short-term forecasting remains challenging due to higher variability.
-This limitation suggests that incorporating temporal modeling techniques, such as recurrent or convolutional structures, could refine short-term predictions while maintaining strong long-range forecasting.
-Performance drops on external validation datasets (sepsis-10 and sepsis-100) highlight generalization challenges, with models struggling to adapt to shifts in event distributions.
-While ModernBERT models exhibit relative robustness, domain-specific fine-tuning or adaptive learning strategies are needed to bridge this gap.
-Our survival analysis results complement the forecasting findings by revealing that strong predictive performance is often achievable with limited observation windows.
-Several models reach peak concordance at early (0h) or intermediate (24h) timepoints rather than at the longest window (168h), particularly in the T2S2 and Sepsis100 cohorts.
-Instruction-tuned LLMs such as Llama-3.3-70B-Instruct and DeepSeek-R1-Distill-Llama-70B outperform traditional transformer baselines in these survival settings, suggesting that instruction tuning and large-scale pretraining confer meaningful advantages for modeling time-to-event outcomes from clinical narratives in the textual time series setup.
-Sensitivity analyses further highlight the importance of temporal structure and data completeness.
-When comparing time-ordered and text-ordered training regimes, we find that time-ordered training generally improves concordance, particularly for DeepSeek-R1 annotations, while text-ordering can yield competitive or better F1 scores on external datasets.
-This underscores a trade-off between temporal alignment and narrative coherence, with implications for designing learning objectives based on downstream task priorities.
-Moreover, robustness experiments simulating documentation sparsity reveal that F1 scores degrade significantly as more clinical history is masked, particularly beyond a timestep dropout rate of 60%.
-In contrast, concordance remains relatively stable under these conditions, indicating that event ranking is less sensitive to partial history than event classification.
-Overall, our results demonstrate that encoder-based models, particularly ModernBERT, are well-suited for event forecasting but require enhancements in short-term prediction and cross-domain generalization.
-In contrast, for survival analysis tasks, we find that instruction-tuned LLMs have more predictive power.
-Future work should explore hybrid training objectives that balance classification and ranking performance, incorporate temporal modeling architectures, and investigate domain adaptation and pretraining strategies that explicitly account for data sparsity and real-world variability that we observe in our clinical textual time series.
-Acknowledgments
-This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health.
-This work utilized the computational resources of the NIH HPC Biowulf cluster.
+Performance drops on external validation datasets highlight generalization challenges, though ModernBERT models exhibit relative robustness.
+For survival analysis, instruction-tuned LLMs like Llama-3.3-70B-Instruct outperform traditional transformer baselines, with several models achieving peak concordance at early timepoints rather than with extended observation.
+Our sensitivity analysis reveals several trade-offs: time-ordered training generally improves concordance while text-ordering can yield better F1 scores on externally-annotated datasets.
+Robustness experiments show F1 scores degrade significantly beyond 60% timestep dropout, while concordance remains stable, indicating event ranking is less sensitive to partial history than event classification.
+Methodological Contributions and Potential Impact.
+Our framework demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks.
+The ability to extract temporally structured insights from unstructured clinical text could potentially support clinical decision-making, particularly in settings with limited access to structured data or specialized expertise.
+Importantly, while consumers frequently consult LLMs about health risk via prompting in chats, this study demonstrates that the prompting approach performs substantially worse in risk prediction than alternative approaches, at least with respect to precision/recall/F1, with prompted LLMs achieving at best 0.460 F1 at 24h compared to 0.653 for fine-tuned encoders.
+Our study highlights several alternative approaches and characterizes their relative performance strengths and weaknesses.
+In addition, our finding that concordance remains stable with 60% missing context suggests robustness to incomplete documentation scenarios, which highlights the degradation pattern in performance as it relates to context availability.
+Our approach also demonstrates that our language model systems can reliably capture the temporal reasoning that clinicians use in practice based on their clinical documentation.
+However, additional work would be required to validate performance on real-time clinical data and demonstrate measurable impact on patient outcomes in deployment.
+Further discussion on societal impact of our work is in Appendix I.
+Limitations and Future Directions.
+Our study has important limitations to consider.
+First, the pipeline relies on case reports from the PubMed Open Access (PMOA) corpus, which often reflect rare or atypical presentations and differ from routine clinical notes (e.g., progress notes, discharge summaries), potentially limiting the model’s generalizability to real-world health system texts.
+Compared to routine clinical notes, case reports offer rich, temporally structured narratives with explicit clinical rationale, making them especially useful for evaluating model performance under sparse documentation.
+This study develops a general framework for textual time series forecasting, using case reports as an interpretable, temporally rich foundation for method development with potential extensions to real-world clinical corpora like MIMIC discharge summaries.
+Second, while we focused on sepsis due to its clinical relevance and prevalence in case reports, our framework is fundamentally disease-agnostic.
+Preliminary results from over 125K PMOA case reports across various diagnostic conditions show promising generalization, with future work planned on broader diagnostic categories and real-world corpora such as MIMIC-IV discharge summaries.
+Third, despite efforts to mitigate causal leakage during training and evaluation (e.g., timestamp extraction, temporal masking, and use of external test sets), pretrained language models may have been exposed to PMOA content.
+Although these models were not trained on our derived (event, time) sequences or temporally reordered data, it is possible that the original case report narratives were included in their pretraining data.
+Our pipeline applies a series of transformations—including temporal reordering and masking of future events beyond a prediction cutoff—which fundamentally alter the input representation and enforce a causal structure that is not present in the original documents.
+This creates a distinct prediction setup that requires models to reason over partial, temporally grounded inputs rather than complete narratives.
+To further assess the impact of pretraining exposure, we include results from RedPajama-INCITE-7B-Instruct, an open-source model with a fully documented, PubMed-free training corpus.
+Its comparable performance to other decoder models (0.352 F1 at 24h) and similar concordance (0.618) to models potentially exposed to PubMed provides evidence that our T2S2 prediction task is sufficiently distinct from raw PubMed text, helping to mitigate causal leakage concerns.
+A detailed discussion on the scope of using open-source models for avoiding potential data leakage has been presented in Appendix H.
+Finally, our analysis is limited by the modest size of the high-cost manual annotations collected as a gold standard to calibrate and validate our LLM extraction pipeline.
+Expanding gold-standard validation across diverse diagnoses could assess the generalizability of our approach and aid in evaluating model robustness, and is a direction for future work.
+Supplementary Material
+This supplementary material provides detailed information organized as follows:
+  • Appendix A: Complete forecasting performance tables for all models on DeepSeek-R1 and Llama-3.3-70B annotations
+  • Appendix B: Sensitivity analysis of masking history on Llama-3.3-70B annotations
+  • Appendix C: Technical details of encoder-based forecasting methods, including MLP-head and masking approaches
+  • Appendix D: Survival modeling framework using language model embeddings and the hyperparameters used
+  • Appendix E: Computing infrastructure used for experiments, including hardware specifications, memory requirements, operating system, and software environment
+  • Appendix E: LLM prompt template for extracting clinical time series from case reports
+  • Appendix E: Prompting strategies for zero-shot and few-shot forecasting
+  • Appendix H: Analysis of open-source models to address data leakage concerns
+  • Appendix I: Broader societal impact and ethical considerations
 Appendix A Forecasting performance on DeepSeek-R1 and Llama-3.3-70B textual time-series annotations
-Tables 6 and 7 present the forecasting performance of our models on textual time-series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively.
-We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
-Table 6: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
+Table A1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
+Models italicized represent the open-source and medically fine-tuned models.
 All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1.
@@
 L3.1 8B 0.000 0.355 0.650 0.630 0.114 0.575 0.442 0.604
+OLMO 32B Instruct 0.000 0.411 0.688 0.621 0.273 0.561 0.432 0.593
+RedPajama-INCITE 7B Instruct 0.000 0.352 0.651 0.618 0.282 0.572 0.471 0.594
+MediPhi-PubMed 0.000 0.268 0.751 0.626 0.169 0.561 0.210 0.597
 Prompting
@@
 L3.3 70B few-shot ordinal 0.063 0.356 0.682 0.313 0.480
+OLMO 32B 0-shot 0.018 0.333 0.652 0.322 0.430
+OLMO 32B 0-shot ordinal 0.006 0.323 0.685 0.325 0.443
+OLMO 32B few-shot 0.014 0.316 0.630 0.219 0.423
+OLMO 32B few-shot ordinal 0.013 0.315 0.651 0.247 0.422
 Encoder-MLP head
@@
 ModernBERT-large 0.277 0.568 0.874 0.598 0.391 0.539 0.602 0.553
+Bioclinical ModernBERT-base 0.264 0.559 0.879 0.677 0.395 0.598 0.635 0.610
+Bioclinical ModernBERT-large 0.313 0.562 0.868 0.671 0.399 0.491 0.618 0.607
 Encoder-masking-fine-tuned
@@
 ModernBERT-large 0.177 0.460 0.683 0.663 0.342 0.615 0.496 0.632
+Bioclinical ModernBERT-base 0.166 0.450 0.692 0.676 0.336 0.604 0.499 0.655
+Bioclinical ModernBERT-large 0.165 0.422 0.696 0.652 0.325 0.626 0.425 0.603
 Encoder-masking-zeroshot
@@
 ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
-Table 7: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
+Bioclinical ModernBERT-base 0.246 0.458 0.614 0.498 0.301 0.509 0.457 0.498
+Bioclinical ModernBERT-large 0.244 0.446 0.596 0.479 0.364 0.502 0.416 0.497
+Table A2: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
+Models italicized represent the open-source and medicallty fine-tuned models.
 All models are trained/fine-tuned on time-ordered annotations from Llama-3.3-70B.
@@
 L3.1 8B 0.338 0.565 0.823 0.619 0.368 0.569 0.536 0.596
+OLMO 32B Instruct 0.190 0.315 0.764 0.604 0.238 0.551 0.315 0.596
+RedPajama-INCITE 7B Instruct 0.127 0.459 0.705 0.620 0.257 0.586 0.491 0.594
+MediPhi-PubMed 0.000 0.507 0.801 0.623 0.277 0.564 0.534 0.600
 Prompting
@@
 L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
+OLMO 32B 0-shot 0.036 0.417 0.708 0.254 0.473
+OLMO 32B 0-shot ordinal 0.034 0.457 0.741 0.167 0.427
+OLMO 32B few-shot 0.013 0.428 0.631 0.267 0.424
+OLMO 32B few-shot ordinal 0.046 0.413 0.647 0.300 0.424
 Encoder-MLP head
@@
 ModernBERT-large 0.312 0.622 0.897 0.500 0.404 0.500 0.623 0.500
+Bioclinical ModernBERT-base 0.290 0.653 0.902 0.650 0.449 0.618 0.662 0.627
+Bioclinical ModernBERT-large 0.258 0.648 0.901 0.626 0.409 0.602 0.634 0.593
 Encoder-masking-fine-tuned
@@
 ModernBERT-large 0.124 0.470 0.712 0.646 0.318 0.586 0.468 0.631
+Bioclinical ModernBERT-base 0.129 0.481 0.733 0.653 0.331 0.615 0.475 0.658
+Bioclinical ModernBERT-large 0.131 0.480 0.840 0.658 0.318 0.587 0.478 0.641
 Encoder-masking-zeroshot
@@
 RoBERTa 0.074 0.279 0.445 0.500 0.083 0.500 0.319 0.500
-DeBERTa-small 0.008 0.439 0.839 0.562 0.296 0.518 0.518 0.501
+DeBERTa-small 0.008 0.439 0.839 0.562 0.296 0.518 0.515 0.501
 ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
-ModernBERT-large 0.109 0.413 0.622 0.499 0.326 0.509 0.339 0.518
+ModernBERT-large 0.109 0.413 0.622 0.499 0.326 0.521 0.339 0.518
+Bioclinical ModernBERT-base 0.197 0.460 0.533 0.499 0.299 0.520 0.465 0.499
+Bioclinical ModernBERT-large 0.194 0.451 0.525 0.496 0.342 0.489 0.433 0.501
+Tables A1 and A2 present the forecasting performance of our models on textual time-series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively.
+We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
 Appendix B Masking history: Llama-3.3-70B textual time-series annotations
-Table 8 examines the impact of randomly masking portions of the input history on model performance when using Llama-3.3-70B-generated annotations.
+Table A3 examines the impact of randomly masking portions of the input history on model performance when using Llama-3.3-70B-generated annotations.
 We evaluate performance across three datasets (T2S2-L33, Sepsis-10, and Sepsis-100) using F1 scores for prediction of next $k$ events happening with the next day and concordance index (c-index) correctly ordering the next $k$ events.
@@
 Interestingly, the concordance index remains relatively stable, indicating the model’s robustness in ranking the next $k$ events with partial input histories.
-Table 8: Effect of Randomly Masking History on F1 and Concordance (Llama-3.3 Annotations)
-TDR F1 (1 day) Concordance (c-index)
-T2S2-L33 Sepsis-10 Sepsis-100 T2S2-L33 Sepsis-10 Sepsis-100
-0% 0.645 0.431 0.626 0.632 0.595 0.612
-15% 0.628 0.422 0.624 0.635 0.592 0.610
-30% 0.634 0.423 0.621 0.634 0.586 0.598
-60% 0.604 0.391 0.605 0.634 0.577 0.613
-90% 0.531 0.328 0.534 0.622 0.546 0.607
+Table A3: Effect of Randomly Masking History (Llama-3.3-70B annotations) on F1 and Concordance.
+Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively.
+c10/100: c-index for sepsis-10 and sepsis-100 respectively.
+Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
+0% 0.645 0.632 0.431 0.595 0.626 0.612
+15% 0.628 0.635 0.422 0.592 0.624 0.610
+30% 0.634 0.634 0.423 0.586 0.621 0.598
+60% 0.604 0.634 0.391 0.577 0.605 0.613
+90% 0.531 0.622 0.328 0.546 0.534 0.607
 Appendix C Textual Time Series Forecasting Framework of Encoder Models
 We develop a general framework for predictive modeling over textual time series of the encoder-only models, designed to accommodate the irregular, event-based structure common to clinical case reports.
-As highlighted Section 3.2, each data instance consists of a sequence of clinical events, where each event is a free-text string describing a clinical finding, intervention, or diagnosis, paired with a timestamp in hours relative to the time of admission.
+As highlighted Section Textual Time Series Annotation, each data instance consists of a sequence of clinical events, where each event is a free-text string describing a clinical finding, intervention, or diagnosis, paired with a timestamp in hours relative to the time of admission.
 The format resembles a two-column table, where the first column contains textual clinical events and the second column contains integer-valued timestamps.
@@
 For a fixed forecast horizon parameter $K$, we generate overlapping windows of historical events and forecast targets.
-Specifically, for each case report with $L$ rows, we construct a sequence of up to $L-K$ forecasting examples.
+Specifically, for each case report with $L$ rows, we construct a sequence of up to LKL-K forecasting examples.
 Each example is structured as follows:
@@
 The input to the language model takes the form:
-[…]
+$\texttt{[CLS]}\;t_{1}\texttt{:}e_{1}\;\texttt{[SEP]}\;t_{2}\texttt{:}e_{2}\;\texttt{[SEP]}\;\dots\;\texttt{[SEP]}\;t_{n}\texttt{:}e_{n}$ (1)
 This is followed by additional target-specific components, as described below.
@@
   • microsoft/deberta-v3-small (He et al., 2021)
-  • answerdotai/ModernBERT-base and answerdotai/ModernBERT-large (pretrained on modern clinical narratives)
+  • answerdotai/ModernBERT-base and answerdotai/ModernBERT-large
+  • thomas-sounack/BioClinical-ModernBERT-base and thomas-sounack/BioClinical-ModernBERT-large
 Each model is used with its corresponding tokenizer.
@@
   • Input: The serialized history segment followed by:
-[…]
+    $\texttt{[SEP]}\;\tau_{i}\texttt{:}f_{i}\;\texttt{[SEP]}\;\tau_{j}\texttt{:}f_{j}\;\texttt{[SEP]}$ (2)
   • Label: $y=1$ if $\tau_{i}<\tau_{j}$, else $y=0$
@@
   • Input: The serialized history segment followed by:
-[…]
+    $\texttt{[SEP]}\;\tau_{j}\texttt{:}f_{j}\;\texttt{[SEP]}$ (3)
   • Label: $y=1$ if $\tau_{j}-t_{n}\leq H$, where $t_{n}$ is the timestamp of the last historical event.
 This is again treated as a binary classification task over the [CLS] embedding, using a linear layer and softmax. We report macro-averaged F1 scores over all prediction examples.
+Masked Language Modeling Approaches.
+In addition to the MLP-head approach described above, we implement masked language modeling (MLM) variants that leverage the pretrained masked token prediction capabilities of encoder models.
+These approaches reformulate our forecasting tasks as cloze-style problems:
+1. Masked Event Occurrence Prediction.
+For the time-window classification task, we construct prompts that include a [MASK] token whose prediction indicates whether an event occurs within the specified horizon:
+  • Input: History formatted as: [CLS] $h_{1}$ [SEP] $h_{2}$ $[SEP]$ Will $e_{j}$” happen within $H$ hours? [MASK] [SEP]
+  • Target: The model predicts either ”yes” or ”no” at the [MASK] position
+We constrain the output vocabulary to only these two tokens and compute cross-entropy loss over their logits.
+For fine-tuned variants, we train the entire model to optimize this masked prediction task.
+For zero-shot variants, we directly use the pretrained model’s predictions without additional training.
+2. Masked Temporal Ordering.
+For the concordance task, we predict relational tokens that indicate temporal ordering:
+  • Input: History followed by: [SEP] $e_{1}$ [MASK] $e_{2}$ [SEP]
+  • Target: The model predicts ”before” if $e_{1}$ occurs before $e_{2}$, else ”after
+Again, we restrict predictions to these two tokens. This formulation allows the model to leverage its pretrained understanding of temporal relationships encoded during MLM pretraining.
+Note that in our dataset, we randomize the relative ordering of $e_{1}$ and $e_{2}$ each time to avoid introducing positional biases that could lead the model to rely on superficial cues rather than genuine temporal reasoning.
+Implementation Details.
+For both tasks, we extract the logits at the [MASK] position and apply a linear projection to the target vocabulary size (2 tokens).
+The key advantage of this approach is that it can leverage pretrained masked language modeling capabilities, potentially requiring less task-specific training.
+We evaluate both fine-tuned versions (which update all model parameters) and zero-shot versions (which use frozen pretrained weights).
 Training and Evaluation.
@@
   • Random Survival Forest (RSF): A nonparametric ensemble method based on decision trees, capable of handling complex interactions and high-dimensional inputs.
-The model is tuned over the number of trees (n_estimators), the minimum number of samples required to split an internal node (min_samples_split), and the minimum number of samples at a leaf node (min_samples_leaf).
+The model is tuned over the number of trees (n_estimators), the minimum number of samples required to split an internal node (min_samples_split), and the minimum number of samples at a leaf node (min_samples_leaf); see Table A4 for search ranges.
   • DeepSurv: A neural-network generalization of the Cox proportional hazards model.
-The architecture comprises fully connected layers with varying widths (num_nodes) and dropout rates, with a fixed number of training epochs.
+The architecture comprises fully connected layers with varying widths (num_nodes) and dropout rates; see Table A4 for search ranges.
+The number of training epochs is fixed.
   • DeepHit: A multi-task neural network that jointly models the discrete hazard and survival functions.
-As with DeepSurv, we perform a grid search over hidden layer sizes and dropout rates while keeping the number of epochs fixed.
+As with DeepSurv, we perform a grid search over hidden layer sizes and dropout rates while keeping the number of epochs fixed; see Table A4.
 Hyperparameter tuning for each survival model is performed using a validation split from the training set.
-For each combination of language model and survival model, we conduct grid search over the respective hyperparameter space, selecting the configuration that maximizes the time-dependent concordance index on the validation set.
-Final evaluation is conducted on held-out test sets (t2s2_test, sepsis10, sepsis100), and results are aggregated across survival metrics to assess model performance.
+For each combination of language model and survival model, we conduct grid search over the respective hyperparameter space, selecting the configuration that maximizes the time-dependent concordance index (Antolini, Boracchi, and Biganzoli 2005) on the validation set.
+Final evaluation is conducted on held-out test sets (t2s2_test, sepsis-10, sepsis-100), and results are aggregated across survival metrics to assess model performance.
 This framework allows for a systematic evaluation of how different pre-trained LLMs, and their associated embedding strategies, contribute to downstream survival prediction tasks.
-Appendix E LLM prompt to generate clinical textual time-series from PMOA case reports
-You are a physician.
+Table A4: Hyperparameter search ranges for survival models.
+Model Hyperparameters RSF n_estimators $\in$ {20,50,100,200,500,1000} min_samples_split $\in$ {2,10,20,50,100} min_samples_leaf $\in$ {1,5,10,20} DeepSurv/DeepHit num_nodes $\in$ {32,64,128,256,512,768,1024,1200, 1536,1800,2048,3072,4096} dropout $\in$ {0.1,0.2,0.4,0.5}
+Appendix E Computing Resources
+All experiments were conducted on a high-performance computing cluster running AlmaLinux.
+For the largest language models (DSL33 and L33), training required three NVIDIA A100 GPUs (80 GB each; total 240 GB GPU memory), while inference was performed on two A100 GPUs.
+The encoder models were trained and evaluated on a single NVIDIA A100 GPU (80 GB).
+After getting the embedding vectors from LLMs, survival models were trained and evaluated on a workstation equipped with an NVIDIA RTX 4090 GPU.
+System memory usage was 32 GB RAM.
+Software dependencies, including specific versions of Python libraries and frameworks, are documented in environment_tts_forecasting.yml file of our anonymized repository in the Supplementary Material.
+F  LLM prompt to generate clinical textual time-series from PMOA case reports
+Prompt You are a physician.
 Extract the clinical events and the related time stamp from the case report.
@@
 DRESS syndrome has no specific time, but it should happen soon after admission to the hospital, so we use our clinical judgment to give the diagnosis of DRESS syndrome the timestamp 0.
-then the output should look like: 18 years old | 0
+then the output should look like:
+18 years old | 0
 male | 0
-admitted to the hospital | 0 fever | -72
+admitted to the hospital | 0
+fever | -72
 rash | -72
@@
 The time is a numeric value in hour unit.
-The two columns are separated by a pipe ‘|’ as a bar-separated file.
+The two columns are separated by a pipe $\mid$ as a bar-separated file.
 Skip the title of the table.
 Reply with the table only.
-Appendix F Prompting strategies for forecasting tasks
-Zero-shot:
-You are an expert physician.
+G  Prompting strategies for forecasting tasks
+F.1  Zero-shot
+Prompt You are an expert physician.
 Reply to the prompt with structured predictions in a k-item, bar-separated row.
-For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: 0 | 1 | 0 would be the correct response.
+For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response.
 Reply with the answer only.
@@
 Do NOT explain.
-Zero-shot ordinal:
-You are an expert physician.
+F.2  Zero-shot (with degree of confidence)
+Prompt You are an expert physician.
 Reply to the prompt with structured predictions in a k-item, bar-separated row.
@@
 Do NOT explain.
-Few-shot:
-You are an expert physician.
+F.3  Few-shot
+Prompt You are an expert physician.
 Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window.
 Example 1.
-If there k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 | 1 | 0 would be the correct response.
-Example 2. Suppose the time series is:
-admitted | 0
+If there k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response.
+Example 2.
+Suppose the time series is: admitted | 0
 history of diabetes | 0
@@
 office visit | 1440
-Then, if the prediction is at 5 hours, the time preceding ‘‘--- PREDICTION WINDOW ---’’, the events admitted, history of diabetes, and elevated troponin have occurred.
+Then, if the prediction is at 5 hours, the time preceding “— PREDICTION WINDOW —”, the events admitted, history of diabetes, and elevated troponin have occurred.
 Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor, and discharge happen.
-So the correct response for the example is: 1 | 1 | 0 | 0 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
-Reply with the answer only. Do NOT provide any other text. Do NOT explain.
-Few-shot ordinal:
-You are an expert physician.
+So the correct response for the example is: 1 $\mid$ 1 $\mid$ 0 $\mid$ 0 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
+Reply with the answer only.
+Do NOT provide any other text.
+Do NOT explain.
+F.4  Few-shot ordinal
+Prompt You are an expert physician.
 Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window, using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen.
-Example 1. If there k=3 events (A, B, C) and only B occurs in the time window, then: 1 | 5 | 1 would be the correct response.
-Example 2. Suppose the time series is:
-admitted | 0
+Example 1.
+If there k=3 events (A, B, C) and only B occurs in the time window, then: 1 $\mid$ 5 $\mid$ 1 would be the correct response.
+Example 2.
+Suppose the time series is: admitted | 0
 history of diabetes | 0
@@
 office visit | 1440
-Then, if the prediction is at 5 hours, the time preceding ‘‘--- PREDICTION WINDOW ---’’, the events admitted, history of diabetes, and elevated troponin have occurred.
+Then, if the prediction is at 5 hours, the time preceding “— PREDICTION WINDOW —”, the events admitted, history of diabetes, and elevated troponin have occurred.
 Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor and discharge happen.
-So the an excellent response for the example is: 5 | 4 | 1 | 1 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
-Reply with the answer only. Do NOT provide any other text. Do NOT explain.
+So the an excellent response for the example is: 5 $\mid$ 4 $\mid$ 1 $\mid$ 1 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
+Reply with the answer only.
+Do NOT provide any other text.
+Do NOT explain.
+Appendix H Open-source models for avoiding potential data leakage
+We recognize that data leakage is a critical concern in evaluating model performance, particularly when using open-access text corpora.
+Below, we outline key factors that mitigate this concern in our study:
+1.
+Temporal annotations—not raw text—drive forecasting tasks.
+The core contribution of this work lies in the temporal structuring of clinical narratives.
+While large language models (LLMs) such as DeepSeek and Llama-3.3 may have encountered PubMed Open Access (PMOA) case reports during pretraining, they were never exposed to the extracted timestamps or temporally ordered and future-masked (event, time) tuples used in our evaluations.
+These derived temporal annotations are essential to all forecasting tasks in our study—including F1-based next-event prediction, temporal ordering, and survival analysis—and constitute novel structured data not present in the original text.
+2.
+Temporal reordering fundamentally alters the learning problem.
+Even if LLMs were pretrained on PMOA narratives, they would have encountered events in their original, text-based order.
+In contrast, our framework chronologically reorders events based on extracted timestamps, creating a new temporal structure.
+This transformation introduces a distinct modeling challenge, as it breaks the sequential narrative flow seen during pretraining.
+As shown in Section 4.4.1, training on time-ordered versus text-ordered annotations leads to measurable differences in performance, underscoring the impact of this restructuring.
+3.
+External validation reveals limited effects of memorization.
+Our experimental design separates internal (T2S2 test set) and external validation (sepsis-10 and sepsis-100 cohorts), using entirely held-out case reports.
+The performance drop observed on external datasets (Tables 1–2) suggests that any potential memorization of PMOA case reports during pretraining yields minimal benefit and reinforces the robustness of our empirical findings.
+4.
+Broader context: standard practices and safeguards.
+Training data overlap is a general challenge in modern machine learning, especially for LLMs trained on web-scale corpora.
+Our work adheres to community standards while incorporating additional safeguards—such as explicit temporal annotation, event reordering, and external dataset validation—that reduce the risk of leakage and differentiate our tasks from standard text modeling.
+To address concerns regarding potential pretraining overlap, we evaluated multiple models with documented training corpora.
+Both OLMO-32B-Instruct (Allen Institute for AI) and RedPajama-INCITE-7B-Instruct represent fully open-source models with transparent training data.
+Notably, RedPajama explicitly excludes PubMed from its training corpus, making it an ideal benchmark for assessing causal leakage.
+The comparable performance of these models to those potentially exposed to PMOA (see point 5 above) provides strong evidence that our temporal forecasting tasks are meaningfully distinct from standard language modeling on clinical narratives.
+5.
+Empirical evidence from PubMed-excluded models.
+Our results provide direct empirical evidence against meaningful causal leakage.
+RedPajama-INCITE-7B-Instruct, which explicitly excludes PubMed from its training corpus, achieves comparable performance to other decoder models (0.352 F1 at 24h, 0.618 concordance) that potentially had access to PMOA during pretraining.
+This similarity in performance across models with and without PubMed exposure demonstrates that our T2S2 prediction task—with its temporal restructuring, event masking, and partial history inputs—is sufficiently distinct from raw case report text that prior exposure provides no meaningful advantage.
+Furthermore, OLMO-32B-Instruct shows competitive performance (0.411 F1 at 24h), providing additional validation from a model with fully documented training data.
+Appendix I Broader Impact
+This work advances equitable, scalable, and transparent medical AI by introducing a framework for forecasting clinical risk in sepsis from narrative documentation.
+Sepsis remains a leading cause of preventable mortality worldwide, with outcomes highly dependent on timely recognition and intervention.
+In many under‑resourced healthcare settings, structured EHR data and real‑time specialist input are unavailable, leaving unstructured narratives as the primary record of patient care.
+Our approach transforms such narratives into temporally structured representations, enabling systematic evaluation of forecasting methods for short‑ and long‑horizon event prediction, event ordering, and survival analysis.
+By releasing the dataset, extraction pipeline, and evaluation framework, we lower barriers for developing and benchmarking temporal forecasting models in sepsis, and promote reproducibility in LLM‑based clinical information extraction.
+However, societal considerations are critical.
+LLM‑assisted extraction can introduce inaccuracies or hallucinations, and unvalidated forecasts risk amplifying bias or error.
+Although timelines are derived from non‑identifiable public case reports, their interpretation could still reinforce historical biases if not carefully contextualized.
+Forecasting results from case reports may also differ from those using routine clinical notes, limiting generalizability.
+Moreover, applying these forecasting models in high‑stakes clinical decision‑making without expert oversight poses safety risks.
+Responsible use requires transparent communication of dataset and model limitations, explicit uncertainty reporting, and safeguards to prevent inappropriate deployment in patient‑facing settings.

### v5

@@
 Clinical case reports encode temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data.
-In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings–extracted via an LLM-assisted annotation pipeline–serve as the primary input for prediction.
+In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings—extracted via an LLM-assisted annotation pipeline—serve as the primary input for prediction.
 We systematically evaluate a diverse suite of models, including fine-tuned decoder-based large language models and encoder-based transformers, on tasks of event occurrence prediction, temporal ordering, and survival analysis.
@@
 Solving this temporal reasoning challenge in case reports establishes the foundation for deployment in live clinical environments where such capabilities could improve patient outcomes.
-Figure 1: Example case report (left) with temporal representations as timelines (middle) and text-ordered, event-time tuples (right).
+Figure 1: Example case report (left) with timeline representations (middle) and text-ordered, (event, time) tuples (right).
 One might ask whether large pretrained language models can already perform forecasting from clinical narratives, given their exposure to biomedical texts (Peng, Chen, and Lu 2020; Weber et al.
@@
 Encoder-based models apply random masking that rarely yields temporally coherent representations, while decoder-style models mask text in linear sequence, modeling events in text order rather than time order.
-Our experiments demonstrate these limitations quantitatively–for example, using the timeline in Figure 1, the probability of recovering a valid time-ordered mask is $0.002\%$, and decoder models achieve low temporal concordance (e.g., $c=0.23$), leading to suboptimal forecasting accuracy without task-specific restructuring.
-To address these temporal reasoning challenges, we develop a forecasting framework that transforms free-text narratives into textual time series—sequences of (event, time) tuples and makes explicit the temporal structure of patient trajectories.
-Building on recent annotation work (Wang and Weiss 2025), we propose a hybrid pipeline combining rule-based heuristics and LLM-assisted extraction to recover temporally localized clinical findings, validated on real clinical notes with expert verification.
+Our experiments make these limitations explicit.
+For instance, on the timeline in Figure 1, the chance of recovering a valid time-ordered mask is below $0.002\%$, and decoder models show low temporal concordance (e.g., $c=0.23$).
+Without a task-specific restructuring, such factors lead to suboptimal forecasting performance.
+To address these temporal reasoning challenges, we develop a forecasting framework that transforms free-text narratives into textual time series—sequences of (event, time) tuples—and makes explicit the temporal structure of patient trajectories.
+Building on recent annotation work (Wang and Weiss 2025; Noroozizadeh et al.
+2025), we propose a hybrid pipeline combining rule-based heuristics and LLM-assisted extraction to recover temporally localized clinical findings, validated on real clinical notes with expert verification.
 Our work builds on extensive research in clinical NLP and temporal reasoning (Kohane 1987; Leeuwenberg and Moens 2020; Cheng and Weiss 2023), but shifts the emphasis from extraction to downstream clinical forecasting.
@@
 narrative) impacts generalization and analyze the sensitivity of models to missing historical context through systematic masking experiments.
-In summary, this paper makes the following contributions: (i) Annotation and Extraction Pipeline: Building on prior methods for extracting information from the PubMed Open Access corpus (Noroozizadeh and Weiss 2025), this work introduces a pipeline that transforms case reports into textual time series via regular expression filtering and LLM-assisted extraction, yielding temporally anchored (event, time) tuples while minimizing causal leakage by restricting input to past events; (ii) Comprehensive Model Comparison: We conduct a comprehensive evaluation of encoder- and instruction-tuned decoder-based models—using both fine-tuned MLP heads and prompting—across event prediction, temporal ordering, and survival analysis tasks.
+In summary, this paper makes the following contributions:
+(i) Annotation and Extraction Pipeline: Building on prior methods for extracting information from the PubMed Open Access corpus (Noroozizadeh and Weiss 2025), this work introduces a pipeline that transforms case reports into textual time series via regular expression filtering and LLM-assisted extraction, yielding temporally anchored (event, time) tuples while minimizing causal leakage by restricting input to past events; (ii) Comprehensive Model Comparison: We conduct a comprehensive evaluation of encoder- and instruction-tuned decoder-based models—using both fine-tuned MLP heads and prompting—across event prediction, temporal ordering, and survival analysis tasks.
 Results show no single model excels universally, underscoring the need to align model choice with the specific forecasting objective; (iii) Temporal- versus Text- Ordering: We investigate the impact of annotation order by comparing training on time-ordered versus text-ordered data.
@@
 While higher masking levels reduce classification performance (F1), the concordance index remains largely stable, highlighting differing sensitivities of binary prediction and ranking tasks to historical context in textual time series; (v) Methodological Framework for Temporal Clinical Forecasting: Our approach demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks, providing a replicable methodology that could be adapted to other clinical and non-clinical text sources.
-Related Works
+Related Work
 Temporal Information Extraction from Clinical Text:
 Extracting timelines from clinical narratives is a challenging biomedical NLP task.
-The i2b2 2012 challenge introduced annotations datasets for temporal relation extraction from discharge summaries (Uzuner et al.
+The i2b2 2012 challenge introduced annotated datasets for temporal relation extraction from discharge summaries (Uzuner et al.
 2011).
 Subsequent methods linked clinical events to timestamps or temporal expressions (Leeuwenberg and Moens 2020; Frattallone-Llado et al.
-2024), typically using pre-defined event spans.
-We adopt the approach of Noroozizadeh and Weiss (2025) to assign timepoints to findings from full-length case reports, enabling finer temporal resolution.
+2024), typically assuming pre-defined event spans.
+We adopt our prior approach (Noroozizadeh and Weiss 2025; Noroozizadeh et al.
+2025) to assign timepoints to findings from full-length case reports, enabling finer temporal resolution.
 Unlike methods using structured EHR data, we focus solely on text—crucial for sources like PubMed that lack structured metadata.
@@
 2023).
-To our knowledge, this is the first work applying LLMs to text-derived time series in clinical domains.
+To our knowledge, this is the first work to forecast from LLM-derived time series in clinical domains.
 While prior sepsis prediction models rely on vitals or scores like SOFA and SAPS (Hou et al.
@@
 2024), reinforcing the need for task-specific evaluations like ours.
-Methods [^1] Our code is available at https://github.com/Shahriarnz14/Textual-Time-Series-Forecasting
+Methods [^1] Our code can be found at: https://github.com/Shahriarnz14/Textual-Time-Series-Forecasting.
+   The extended version of this paper, including the full technical appendix, is available in Noroozizadeh, Kumar, and Weiss (2025).
 We next describe our dataset construction and annotation pipeline, define the forecasting and survival prediction tasks, detail the modeling approaches evaluated, and present the sensitivity analyses used to assess robustness (Figure 2).
@@
 Dataset Extraction
-We utilized the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis.
-Building on the method described in prior work (Wang and Weiss 2025), the body of each case was extracted by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus.
-To maintain relevance, we include only documents containing case-insensitive matches to (case report|case presenta) and year-?
-?old, an approach shown to have higher specificity than PubMed metadata (Noroozizadeh and Weiss 2025).
-As demonstrated in prior work of (Noroozizadeh and Weiss 2025), additional filtering was applied to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care.
+We used the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis.
+Following prior work (Wang and Weiss 2025), we extracted the body of each case by selecting the text between the ==== Body and ==== Ref delimiters in the PMOA corpus.
+To maintain relevance, we included only documents with case-insensitive matches to (case report|case presenta) and year-?
+?old, an approach shown to be more specific than relying on PubMed metadata (Noroozizadeh and Weiss 2025).
+Additional filtering was applied to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care.
 We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender.
@@
 For instance, instead of splitting “pain in chest that radiates substernally” into less informative parts, we retain the full phrase as a single, complete finding to preserve its meaning and context.
-(2) To improve clarity, coordinated phrases are broken into individual findings.
+(2) To improve clarity, compound phrases are broken into individual findings.
 For example, “metastases in the liver and pancreas” becomes “metastasis in the liver” and “metastasis in the pancreas.”
@@
 We refer to this as Textual Time-Series for Sepsis (T2S2).
-The exact LLM extraction prompt can be found in Appendix E.
+Details on the quality evaluation of LLM-extracted temporal annotations and the prompts used for extraction are provided in Appendix A and A, respectively.
 Our dataset was split into training (n = 1,795; 80% train, 20% validation), testing T2S2-test (n = 244), and two external validation sets: sepsis-10 (n = 10) and sepsis-100 (n = 90).
@@
 At each time cut-off $t$, we extract the next $k$ events from the timeline and remove their timestamps.
-The model must correctly output a permutation over these events that matches their true chronological order.
+The model must output a permutation of these events that matches their true chronological order.
 This is framed as a ranking task, evaluated by computing the pairwise concordance between the predicted and true orderings (e.g., proportion of correctly ordered pairs).
@@
 Many case reports specify whether the patient died and, if so, when (e.g., ”the patient died on hospital day 10”), which enables us to define a time-to-event outcome.
-For this task, we evaluate models at predefined cut-off times – specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week) – and use the extracted event history up to each cut-off as input.
+For this task, we evaluate models at predefined cut-off times— specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week)— and use the extracted event history up to each cut-off as input.
 A survival model is trained to predict the probability of survival over time beyond each cut-off.
-We evaluate using the time-dependent concordance index to measure alignment of the predicted survival times with actual outcomes.
+We evaluate using the time-dependent concordance to measure alignment of the predicted survival times with actual outcomes.
 Modeling Approaches
 To evaluate performance on the event forecasting and survival prediction tasks, we implement five modeling paradigms: (i) decoder-only large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), (iii) encoder-only models with task-specific fine-tuned heads, (iv) encoder-masking models with fine-tuning, and (v) encoder-masking models in zero-shot settings.
-In Appendix C and  D, we provide detailed explanations of implementation details.
+In Appendix D and E, we provide technical details of encoder-based forecasting methods and the survival modeling framework respectively.
 (i) Fine-tuned LLMs.
@@
 2024; Liu et al.
-2024).
+2024a).
 We also include open-source models with documented training corpora (OLMO-32B-Instruct (OLMo et al.
 2024), RedPajama-INCITE-7B-Instruct (Weber et al.
-2024)) to address potential data leakage concerns–notably, RedPajama explicitly excludes PubMed from its training data.
+2024)) to address potential data leakage concerns—notably, RedPajama explicitly excludes PubMed from its training data.
 Additionally, we evaluate a medically fine-tuned decoder model (MediPhi-PubMed (Corbeil et al.
@@
 Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format.
-Prompt designs are customized for each task and are detailed in Appendix E.
+Prompt designs are customized for each task and are detailed in Appendix A.
 Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
@@
 We conduct sensitivity analyses to examine how the temporal structure and completeness of input sequences affect forecasting performance.
-These analyses probe the robustness of models to changes in event ordering and to varying levels of missing historical information:
+These analyses probe the robustness of models to changes in event ordering and to varying levels of missing historical information.
 Time-ordering Strategies.
@@
 This comparison isolates the effect of narrative sequencing versus explicit temporal ordering on model performance.
-For text-ordered inputs, to prevent causal leakage, events occurring after the forecast cutoff time $t$ are masked for encoder-based models and omitted for decoder-based models.
+For text-ordered inputs, to prevent causal leakage, events occurring after the forecast cutoff time $t$ are masked for encoder-based models and omitted for decoder models.
 Timestep Dropout.
@@
 This setup enables controlled evaluation of model performance under partial information, probing the extent to which predictive accuracy and event ordering depend on the completeness of historical context.
-Together, these analyses characterize model robustness to disruptions in temporal coherence and input completeness – two key challenges in real-world deployment scenarios.
-Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
-All models are trained/fine-tuned on time-ordered annotations from either DeepSeek-R1 or Llama 3.3-70B.
-Bold values indicate best in category within each column group (refer to Tables A1 and A2 in Appendix A for detailed performance statistics on all models).
-Performance statistics for all models are presented in A1 and A2, respectively.
-Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively.
-c10/100: c-index for sepsis-10 and sepsis-100 respectively.
-T2S2 (DeepSeek-R1) T2S2 (Llama 3.3-70B)
+T2S2 (DeepSeek-R1) T2S2 (Llama-3.3-70B)
 F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100 F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
@@
 Bioclinical ModernBERT-base 0.246 0.458 0.614 0.498 0.301 0.509 0.457 0.498 0.197 0.460 0.533 0.499 0.299 0.520 0.465 0.499
+Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
+All models are trained/fine-tuned on time-ordered annotations from either DeepSeek-R1 or Llama 3.3-70B.
+Bold values indicate best in category within each column group (refer to Tables A3 and A4 in Appendix E for detailed performance statistics on all models).
+Performance statistics for all models are presented in A3 and A4, respectively.
+Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively.
+c10/100: c-index for sepsis-10 and sepsis-100 respectively.
 Results
 In this section, we present results for our three main evaluation settings: event forecast within a subsequent time window, temporal ordering of forecasted events, and survival analysis.
-We then provide sensitivity analyses on the forecasting tasks, examining the effects of temporal ordering and historical context availability.
+We provide sensitivity analyses on the forecasting tasks, examining the effects of temporal ordering and historical context availability in Appendix E and E.
 Forecasting Tasks
@@
 Our results show that encoder-based models outperform decoder-based LLMs in event forecasting across both DeepSeek-R1 and Llama-3.3-70B annotations (Table 1).
-Forecasting performance across all models is shown in Tables A1 and A2 (Appendix A).
+Forecasting performance across all models is shown in Tables A3 and A4 (Appendix E).
 Encoder models, especially those with a fine-tuned MLP head, achieve substantially higher F1 scores than decoder LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
-Among decoder models, we observe varied performance patterns.
-RedPajama-INCITE-7B-Instruct, with documented training that excludes PubMed, achieves 0.352 F1 at 24h—better than some decoder models like MediPhi-PubMed (0.268) and within the range of decoder performance, though below top performers like DeepSeek-Llama-3.3-70B (0.482).
-OLMO-32B-Instruct performs competitively (0.411), suggesting that open models without biomedical pretraining can still achieve reasonable results for temporal forecasting tasks.
+Among decoders, we observe varied performance patterns.
+RedPajama-INCITE-7B-Instruct, whose training excludes PubMed, achieves 0.352 F1 at 24h—better than some decoder models like MediPhi-PubMed (0.268) and within the typical decoder performance range, though below top performers like DeepSeek-Llama-3.3-70B (0.482).
+OLMO-32B-Instruct performs competitively (0.411), suggesting that open models without biomedical pretraining can still achieve strong temporal forecasting results.
 Among encoder models, both general-purpose and biomedically-pretrained variants show strong performance.
@@
 Temporal ordering of forecasted events: concordance.
-Encoder models consistently outperform decoders in correctly ranking the order of upcoming events, as measured by concordance (c-index; Table 1; complete results in Tables A1 and A2 for DeepSeek-R1 and Llama-3.3-70B annotations, respectively, with full details in Appendix A.
+Encoder models consistently outperform decoders in correctly ranking the order of upcoming events, as measured by concordance (c-index; Table 1; complete results in Tables A3 and A4 for DeepSeek-R1 and Llama-3.3-70B annotations, respectively, with full details in Appendix E.
 Decoder models with potential biomedical exposure achieve modest concordance (0.618-0.632), while RedPajama-INCITE-7B-Instruct without PubMed data in its training corpus performs similarly, suggesting domain-specific knowledge is less critical here than for event prediction.
 The comparable performance of RedPajama-INCITE-7B-Instruct to other general models (with access to PubMed in their pretraining) also provides evidence that the T2S2 prediction task is distinct from raw PubMed text and assuages causal leakage concerns.
-Biomedical pretraining in encoder models also demonstrates clear advantages, with BioClinical-ModernBERT-base attaining the highest concordance (0.677 with MLP head, 0.676 fine-tuned masking) and robust performance across datasets.
+Biomedical pretraining in encoders demonstrates clear advantages, with BioClinical-ModernBERT-base achieving the highest concordance (0.677 with an MLP head, 0.676 with fine-tuned masking) and strong performance across datasets.
 MLP-head training yields superior F1 scores, while fine-tuned masking excels in concordance.
-BioClinical-ModernBERT-base maintains strong generalization on external datasets (c-index >0.60), suggesting that biomedical pretraining provides stability in temporal reasoning across dataset shifts.
+BioClinical-ModernBERT-base also generalizes well to external datasets (c-index >0.60), suggesting that biomedical pretraining provides stability in temporal reasoning under dataset shifts.
 Survival analysis: model performance across timepoints and cohorts.
-Table 2 presents time-dependent concordance across evaluation cohorts (T2S2, sepsis-10, sepsis-100) at multiple observation windows (0h, 24h, 168h), selecting models based on validation concordance.
-General models, such as RedPajama-INCITE-7B-Instruct, show strong survival prediction capabilities even without biomedical exposure (0.76 concordance at 168h, sepsis-100).
-Meanwhile, biomedical encoders like BioClinical-ModernBERT-base consistently perform well, particularly early (0h).
+Tables 2 and A5 (Appendix E) show time-dependent concordance results where instruction-tuned decoder models clearly outperform encoder-based models.
+This pattern contrasts with the forecasting tasks (Table 1), where encoders excelled.
+Across both the T2S2 and sepsis-100 cohorts, decoder models consistently achieve the strongest results.
+Notably, RedPajama-INCITE-7B-Instruct attains the highest concordance overall (0.76 at 168h, sepsis-100) despite lacking any biomedical text in pretraining, while larger models such as DeepSeek-R1-Llama-70B and OLMO-32B-Instruct also perform robustly.
+Encoder models—including biomedically fine-tuned variants like BioClinical-ModernBERT-base, which is competitive on sepsis-100 but not on T2S2—are generally surpassed by decoder models.
 Predictive performance does not universally improve with more observation time.
-T2S2 cohort models often peak at earlier windows, while sepsis-10 experiences ceiling effects.
-Instruction-tuned decoders excel overall, yet biomedical encoders remain competitive, particularly at admission (0h).
-Survival prediction thus benefits from both larger decoder models and domain-specific pretraining, each excelling in different temporal contexts.
-Table 2: Survival analyses evaluated using time-dependent concordance index for each model across DeepSeek-R1 test set annotations (T2S2, Sepsis-10, Sepsis-100), evaluated at three observation time thresholds (0h, 24h, and 168h).
-Model T2S2 sepsis-10 sepsis-100
-Time Threshold 0h 24h 168h 0h 24h 168h 0h 24h 168h
-bert-base-uncased 0.60 0.61 0.53 1.00 1.00 1.00 0.58 0.62 0.58
-roberta-base 0.55 0.60 0.60 0.75 1.00 1.00 0.55 0.58 0.57
-deberta-v3-small 0.56 0.57 0.57 0.75 0.00 0.75 0.52 0.49 0.51
-ModernBERT-base 0.52 0.58 0.58 1.00 1.00 1.00 0.54 0.51 0.51
-ModernBERT-large 0.53 0.59 0.58 0.50 1.00 1.00 0.59 0.55 0.49
-Bioclinical ModernBERT-base 0.57 0.53 0.53 1.00 0.75 0.75 0.63 0.59 0.56
-Bioclinical ModernBERT-large 0.54 0.55 0.54 0.75 0.75 0.75 0.61 0.57 0.49
-DeepSeek-R1-Llama-70B 0.64 0.63 0.59 0.50 1.00 1.00 0.61 0.53 0.55
-Llama-3.3-70B-Instruct 0.62 0.63 0.58 0.50 1.00 0.75 0.64 0.57 0.53
-DeepSeek-R1-Llama-8B 0.60 0.58 0.58 1.00 1.00 1.00 0.53 0.66 0.65
-Llama-3.1-8B-Instruct 0.62 0.61 0.61 1.00 1.00 1.00 0.60 0.59 0.61
-OLMO 32B Instruct 0.60 0.62 0.59 0.75 0.75 0.75 0.66 0.58 0.56
-RedPajama-INCITE 7B Instruct 0.56 0.60 0.61 0.50 1.00 1.00 0.61 0.53 0.76
-MediPhi-PubMed 0.54 0.62 0.55 1.00 1.00 1.00 0.62 0.58 0.61
-Sensitivity analysis
-Time-Ordered versus Text-Ordered Training.
-Table 3 shows time-ordering generally improves concordance, particularly for DeepSeek-R1 annotations, while F1 scores show mixed results.
-Text-ordering sometimes yields better external dataset performance with Llama-3.3-70B, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order.
-These results highlight a trade-off between temporal alignment and narrative coherence for different tasks.
-Table 3: Comparison of text-ordered and time-ordered training for DeepSeek-R1 and Llama 3.3-70B annotations on F1 scores and c-index values.
-F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
-T2S2 (DeepSeek-R1)
-Text-ordered 0.290 0.569 0.876 0.659 0.361 0.598 0.615 0.609
-Time-ordered 0.306 0.576 0.877 0.672 0.428 0.576 0.607 0.613
-T2S2 (Llama 3.3-70B)
-Text-ordered 0.311 0.648 0.898 0.586 0.423 0.526 0.662 0.541
-Time-ordered 0.257 0.645 0.903 0.632 0.382 0.595 0.616 0.612
-Impact of timestep dropout on F1 and concordance.
-Tables 4 and A3 (Appendix B) report the effect of timestep dropout on classification and ranking performance.
-F1 scores decrease monotonically with increasing TDR, with relatively mild degradation up to 30–60% dropout and sharper declines thereafter.
-For example, in the DeepSeek-R1 annotations, T2S2/sepsis-10/sepsis-100 drop from 0.576/0.428/0.607 at TDR = 0% to 0.423/0.200/0.510 at TDR = 90%, with sepsis-10 showing the largest relative decline.
-In contrast, concordance (c-index) remains stable across dropout levels (0.661–0.678), indicating that ranking-based evaluations are less sensitive to history truncation.
-These trends are consistent for Llama-3.3 annotations (Appendix B), suggesting that ranking objectives capture temporal relationships that are more robust to missing historical information than binary classification accuracy of predicting whether an events happens within a time window.
-Table 4: Effect of Randomly Masking History (DeepSeek-R1 annotations) on F1 and Concordance.
-Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
-0% 0.576 0.672 0.428 0.576 0.607 0.613
-15% 0.572 0.667 0.415 0.564 0.604 0.616
-30% 0.548 0.672 0.357 0.542 0.591 0.619
-60% 0.542 0.678 0.310 0.578 0.588 0.616
-90% 0.423 0.661 0.200 0.561 0.510 0.614
-Discussion
+Models evaluated on the T2S2 cohort, often peak at earlier windows (0h or 24h) rather than at 168h.
+Meanwhile, the sepsis-10 cohort experiences clear ceiling effects, with many models achieving perfect concordance (1.00).
+These results suggest that, unlike the forecasting tasks, survival prediction in our dataset benefits more from the capabilities inherent in large decoder models than from domain-specific pretrained encoders.
+Model 0h 24h 168h
+bert-base-uncased 0.60 0.61 0.53
+roberta-base 0.55 0.60 0.60
+deberta-v3-small 0.56 0.57 0.57
+ModernBERT-base 0.52 0.58 0.58
+ModernBERT-large 0.53 0.59 0.58
+Bioclinical ModernBERT-base 0.57 0.53 0.53
+Bioclinical ModernBERT-large 0.54 0.55 0.54
+DeepSeek-R1-Llama-70B 0.64 0.63 0.59
+Llama-3.3-70B-Instruct 0.62 0.63 0.58
+DeepSeek-R1-Llama-8B 0.60 0.58 0.58
+Llama-3.1-8B-Instruct 0.62 0.61 0.61
+OLMO 32B Instruct 0.60 0.62 0.59
+RedPajama-INCITE 7B Instruct 0.56 0.60 0.61
+MediPhi-PubMed 0.54 0.62 0.55
+Table 2: Time-dependent concordance index for survival analysis on the DeepSeek-R1 T2S2 annotations evaluated at 0h, 24h, and 168h.
+Extended sepsis-10 and sepsis-100 results are provided in Appendix E.
+Discussion and Conclusion
 Our findings highlight the superiority of encoder-based models over decoder-based LLMs for event forecasting, emphasizing the limitations of autoregressive models in structured prediction tasks.
@@
 Robustness experiments show F1 scores degrade significantly beyond 60% timestep dropout, while concordance remains stable, indicating event ranking is less sensitive to partial history than event classification.
+Additional insights into the forecasting results are presented in Appendix J.
+Specifically, we summarize two observations that clarify model behavior and apparent anomalies: (i) why decoders fare better on survival but not short-horizon forecasting, and (ii) interpreting very low $F_{1}$ scores and external validation drops.
 Methodological Contributions and Potential Impact.
 Our framework demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks.
-The ability to extract temporally structured insights from unstructured clinical text could potentially support clinical decision-making, particularly in settings with limited access to structured data or specialized expertise.
+The ability to extract temporally structured insights from unstructured clinical text could potentially support clinical decision-making, particularly in settings with limited structured data or specialized expertise.
 Importantly, while consumers frequently consult LLMs about health risk via prompting in chats, this study demonstrates that the prompting approach performs substantially worse in risk prediction than alternative approaches, at least with respect to precision/recall/F1, with prompted LLMs achieving at best 0.460 F1 at 24h compared to 0.653 for fine-tuned encoders.
@@
 However, additional work would be required to validate performance on real-time clinical data and demonstrate measurable impact on patient outcomes in deployment.
-Further discussion on societal impact of our work is in Appendix I.
+Further discussion on societal impact of our work is in Appendix M.
 Limitations and Future Directions.
 Our study has important limitations to consider.
-First, the pipeline relies on case reports from the PubMed Open Access (PMOA) corpus, which often reflect rare or atypical presentations and differ from routine clinical notes (e.g., progress notes, discharge summaries), potentially limiting the model’s generalizability to real-world health system texts.
-Compared to routine clinical notes, case reports offer rich, temporally structured narratives with explicit clinical rationale, making them especially useful for evaluating model performance under sparse documentation.
-This study develops a general framework for textual time series forecasting, using case reports as an interpretable, temporally rich foundation for method development with potential extensions to real-world clinical corpora like MIMIC discharge summaries.
+First, our pipeline relies on case reports from the PubMed Open Access (PMOA) corpus.
+Because these reports often highlight rare or atypical presentations and differ from routine clinical notes (e.g., progress notes, discharge summaries), generalizability to real-world health-system text may be limited.
+However, case reports provide rich, temporally structured narratives with explicit clinical rationale, making them well suited for evaluating model performance under sparse documentation.
+Accordingly, we present a general framework for textual time series forecasting that uses case reports as an interpretable, temporally rich foundation for method development, with extensions to real-world corpora such as MIMIC discharge summaries.
+See Appendix L for a detailed discussion on the generalizability of our pipeline beyond case report details, as well as comparisons with other clinical sources.
 Second, while we focused on sepsis due to its clinical relevance and prevalence in case reports, our framework is fundamentally disease-agnostic.
 Preliminary results from over 125K PMOA case reports across various diagnostic conditions show promising generalization, with future work planned on broader diagnostic categories and real-world corpora such as MIMIC-IV discharge summaries.
-Third, despite efforts to mitigate causal leakage during training and evaluation (e.g., timestamp extraction, temporal masking, and use of external test sets), pretrained language models may have been exposed to PMOA content.
-Although these models were not trained on our derived (event, time) sequences or temporally reordered data, it is possible that the original case report narratives were included in their pretraining data.
-Our pipeline applies a series of transformations—including temporal reordering and masking of future events beyond a prediction cutoff—which fundamentally alter the input representation and enforce a causal structure that is not present in the original documents.
-This creates a distinct prediction setup that requires models to reason over partial, temporally grounded inputs rather than complete narratives.
-To further assess the impact of pretraining exposure, we include results from RedPajama-INCITE-7B-Instruct, an open-source model with a fully documented, PubMed-free training corpus.
-Its comparable performance to other decoder models (0.352 F1 at 24h) and similar concordance (0.618) to models potentially exposed to PubMed provides evidence that our T2S2 prediction task is sufficiently distinct from raw PubMed text, helping to mitigate causal leakage concerns.
-A detailed discussion on the scope of using open-source models for avoiding potential data leakage has been presented in Appendix H.
-Finally, our analysis is limited by the modest size of the high-cost manual annotations collected as a gold standard to calibrate and validate our LLM extraction pipeline.
-Expanding gold-standard validation across diverse diagnoses could assess the generalizability of our approach and aid in evaluating model robustness, and is a direction for future work.
+Finally, despite mitigation steps (timestamp extraction, temporal masking, external test sets), pretrained language models may still have been exposed to PMOA content.
+Although they were not trained on our derived (event, time) sequences or temporally reordered data, the original narratives could appear in pretraining.
+Our pipeline applies temporal reordering and masking of future events beyond the prediction cutoff, enforcing a causal structure absent from the originals and yielding a distinct setup that requires reasoning over partial, temporally grounded inputs.
+While we included RedPajama-INCITE-7B-Instruct, an open-source model with a documented PubMed-free corpus for pretraining, and saw its performance (0.352 F1 at 24h; concordance 0.618) was comparable to decoders potentially exposed to PubMed, other strategies supporting the argument that the T2S2 task is sufficiently distinct from raw PMOA text and mitigates causal leakage concerns could be examined.
+See Appendix K for a detailed discussion of how open-source models mitigate potential data leakage, as well as additional forecasting results on 115 case reports published after the pretraining cutoffs of encoder-based models (post-2020).
+Acknowledgments
+This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health.
+This work utilized the computational resources of the NIH HPC Biowulf cluster.
+S.N.
+was supported by Carnegie Mellon University TCS Presidential Fellowship, and Natural Sciences and Engineering Research Council of Canada (NSERC) PGS-D award.
+S.N.
+was also supported in part by an appointment to the National Library of Medicine Research Participation Program administered by the Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S.
+Department of Energy (DOE) and the National Library of Medicine, National Institutes of Health.
+ORISE is managed by ORAU under DOE contract number DE-SC0014664.
+All opinions expressed in this paper are the authors’ and do not necessarily reflect the policies and views of NIH, NLM, DOE, or ORAU/ORISE.
 Supplementary Material
 This supplementary material provides detailed information organized as follows:
-  • Appendix A: Complete forecasting performance tables for all models on DeepSeek-R1 and Llama-3.3-70B annotations
-  • Appendix B: Sensitivity analysis of masking history on Llama-3.3-70B annotations
-  • Appendix C: Technical details of encoder-based forecasting methods, including MLP-head and masking approaches
-  • Appendix D: Survival modeling framework using language model embeddings and the hyperparameters used
-  • Appendix E: Computing infrastructure used for experiments, including hardware specifications, memory requirements, operating system, and software environment
-  • Appendix E: LLM prompt template for extracting clinical time series from case reports
-  • Appendix E: Prompting strategies for zero-shot and few-shot forecasting
-  • Appendix H: Analysis of open-source models to address data leakage concerns
-  • Appendix I: Broader societal impact and ethical considerations
-Appendix A Forecasting performance on DeepSeek-R1 and Llama-3.3-70B textual time-series annotations
-Table A1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
-Models italicized represent the open-source and medically fine-tuned models.
-All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1.
-Bold indicates best in that particular category amongst all models.
-T2S2 (DeepSeek R1) sepsis-10 sepsis-100
-F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
-LLM-MLP head
-DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
-L3.3 70B 0.000 0.295 0.764 0.631 0.312 0.569 0.404 0.597
-DS-L3.1 8B 0.000 0.492 0.650 0.630 0.381 0.566 0.601 0.600
-L3.1 8B 0.000 0.355 0.650 0.630 0.114 0.575 0.442 0.604
-OLMO 32B Instruct 0.000 0.411 0.688 0.621 0.273 0.561 0.432 0.593
-RedPajama-INCITE 7B Instruct 0.000 0.352 0.651 0.618 0.282 0.572 0.471 0.594
-MediPhi-PubMed 0.000 0.268 0.751 0.626 0.169 0.561 0.210 0.597
-Prompting
-L3.3 70B 0-shot 0.085 0.368 0.560 0.236 0.452
-L3.3 70B 0-shot ordinal 0.082 0.352 0.646 0.289 0.483
-L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452
-L3.3 70B few-shot ordinal 0.063 0.356 0.682 0.313 0.480
-OLMO 32B 0-shot 0.018 0.333 0.652 0.322 0.430
-OLMO 32B 0-shot ordinal 0.006 0.323 0.685 0.325 0.443
-OLMO 32B few-shot 0.014 0.316 0.630 0.219 0.423
-OLMO 32B few-shot ordinal 0.013 0.315 0.651 0.247 0.422
-Encoder-MLP head
-BERT 0.264 0.550 0.871 0.501 0.441 0.498 0.614 0.501
-RoBERTa 0.250 0.562 0.866 0.500 0.405 0.500 0.614 0.500
-DeBERTa-small 0.262 0.570 0.871 0.601 0.382 0.498 0.617 0.556
-ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
-ModernBERT-large 0.277 0.568 0.874 0.598 0.391 0.539 0.602 0.553
-Bioclinical ModernBERT-base 0.264 0.559 0.879 0.677 0.395 0.598 0.635 0.610
-Bioclinical ModernBERT-large 0.313 0.562 0.868 0.671 0.399 0.491 0.618 0.607
-Encoder-masking-fine-tuned
-BERT 0.166 0.432 0.690 0.648 0.323 0.554 0.482 0.600
-RoBERTa 0.169 0.416 0.688 0.646 0.313 0.539 0.457 0.609
-DeBERTa-small 0.175 0.435 0.695 0.671 0.325 0.579 0.492 0.617
-ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
-ModernBERT-large 0.177 0.460 0.683 0.663 0.342 0.615 0.496 0.632
-Bioclinical ModernBERT-base 0.166 0.450 0.692 0.676 0.336 0.604 0.499 0.655
-Bioclinical ModernBERT-large 0.165 0.422 0.696 0.652 0.325 0.626 0.425 0.603
-Encoder-masking-zeroshot
-BERT 0.000 0.000 0.000 0.495 0.000 0.513 0.000 0.494
-RoBERTa 0.049 0.207 0.375 0.500 0.137 0.500 0.318 0.500
-DeBERTa-small 0.170 0.460 0.323 0.497 0.268 0.512 0.376 0.509
-ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
-ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
-Bioclinical ModernBERT-base 0.246 0.458 0.614 0.498 0.301 0.509 0.457 0.498
-Bioclinical ModernBERT-large 0.244 0.446 0.596 0.479 0.364 0.502 0.416 0.497
-Table A2: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
-Models italicized represent the open-source and medicallty fine-tuned models.
-All models are trained/fine-tuned on time-ordered annotations from Llama-3.3-70B.
-T2S2 (Llama 3.3) sepsis-10 sepsis-100
-F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
-LLM-MLP head
-DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
-L3.3 70B 0.265 0.466 0.807 0.619 0.160 0.567 0.466 0.597
-DS-L3.1 8B 0.264 0.584 0.837 0.624 0.258 0.576 0.605 0.598
-L3.1 8B 0.338 0.565 0.823 0.619 0.368 0.569 0.536 0.596
-OLMO 32B Instruct 0.190 0.315 0.764 0.604 0.238 0.551 0.315 0.596
-RedPajama-INCITE 7B Instruct 0.127 0.459 0.705 0.620 0.257 0.586 0.491 0.594
-MediPhi-PubMed 0.000 0.507 0.801 0.623 0.277 0.564 0.534 0.600
-Prompting
-L3.3 70B 0-shot 0.085 0.443 0.616 0.236 0.452
-L3.3 70B 0-shot ordinal 0.063 0.454 0.707 0.289 0.483
-L3.3 70B few-shot 0.099 0.457 0.703 0.360 0.452
-L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
-OLMO 32B 0-shot 0.036 0.417 0.708 0.254 0.473
-OLMO 32B 0-shot ordinal 0.034 0.457 0.741 0.167 0.427
-OLMO 32B few-shot 0.013 0.428 0.631 0.267 0.424
-OLMO 32B few-shot ordinal 0.046 0.413 0.647 0.300 0.424
-Encoder-MLP head
-BERT 0.224 0.622 0.899 0.500 0.400 0.500 0.612 0.500
-RoBERTa 0.180 0.608 0.900 0.500 0.364 0.500 0.597 0.500
-DeBERTa-small 0.258 0.641 0.898 0.603 0.427 0.585 0.616 0.592
-ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
-ModernBERT-large 0.312 0.622 0.897 0.500 0.404 0.500 0.623 0.500
-Bioclinical ModernBERT-base 0.290 0.653 0.902 0.650 0.449 0.618 0.662 0.627
-Bioclinical ModernBERT-large 0.258 0.648 0.901 0.626 0.409 0.602 0.634 0.593
-Encoder-masking-fine-tuned
-BERT 0.131 0.471 0.741 0.622 0.320 0.580 0.467 0.622
-RoBERTa 0.129 0.484 0.659 0.625 0.329 0.587 0.472 0.614
-DeBERTa-small 0.135 0.468 0.693 0.648 0.318 0.602 0.462 0.623
-ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
-ModernBERT-large 0.124 0.470 0.712 0.646 0.318 0.586 0.468 0.631
-Bioclinical ModernBERT-base 0.129 0.481 0.733 0.653 0.331 0.615 0.475 0.658
-Bioclinical ModernBERT-large 0.131 0.480 0.840 0.658 0.318 0.587 0.478 0.641
-Encoder-masking-zeroshot
-BERT 0.000 0.000 0.000 0.504 0.000 0.515 0.000 0.506
-RoBERTa 0.074 0.279 0.445 0.500 0.083 0.500 0.319 0.500
-DeBERTa-small 0.008 0.439 0.839 0.562 0.296 0.518 0.515 0.501
-ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
-ModernBERT-large 0.109 0.413 0.622 0.499 0.326 0.521 0.339 0.518
-Bioclinical ModernBERT-base 0.197 0.460 0.533 0.499 0.299 0.520 0.465 0.499
-Bioclinical ModernBERT-large 0.194 0.451 0.525 0.496 0.342 0.489 0.433 0.501
-Tables A1 and A2 present the forecasting performance of our models on textual time-series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively.
-We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
-Appendix B Masking history: Llama-3.3-70B textual time-series annotations
-Table A3 examines the impact of randomly masking portions of the input history on model performance when using Llama-3.3-70B-generated annotations.
-We evaluate performance across three datasets (T2S2-L33, Sepsis-10, and Sepsis-100) using F1 scores for prediction of next $k$ events happening with the next day and concordance index (c-index) correctly ordering the next $k$ events.
-As the proportion of masked history increases, we observe a general decline in F1 performance, particularly at higher masking rates (e.g., 90%), suggesting that access to historical context is important for accurate forecasting.
-Interestingly, the concordance index remains relatively stable, indicating the model’s robustness in ranking the next $k$ events with partial input histories.
-Table A3: Effect of Randomly Masking History (Llama-3.3-70B annotations) on F1 and Concordance.
-Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively.
-c10/100: c-index for sepsis-10 and sepsis-100 respectively.
-Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
-0% 0.645 0.632 0.431 0.595 0.626 0.612
-15% 0.628 0.635 0.422 0.592 0.624 0.610
-30% 0.634 0.634 0.423 0.586 0.621 0.598
-60% 0.604 0.634 0.391 0.577 0.605 0.613
-90% 0.531 0.622 0.328 0.546 0.534 0.607
-Appendix C Textual Time Series Forecasting Framework of Encoder Models
+  • Appendix A: Validity of LLM-extracted temporal annotations (adopted from our prior work (adopted from our prior work (Wang and Weiss 2025)).
+  • Appendix A: LLM prompt template for extracting clinical time series from case reports.
+  • Appendix A: Prompting strategies for zero-shot and few-shot forecasting.
+  • Appendix D: Technical details of encoder-based forecasting methods, including MLP-head and masking approaches.
+  • Appendix E: Survival modeling framework using language model embeddings and the hyperparameters used.
+  • Appendix E: Complete forecasting performance tables for all models on DeepSeek-R1 and Llama-3.3-70B annotations.
+  • Appendix E: Complete survival analysis performance table for all datasets.
+  • Appendix E: Sensitivity analysis of the forecasting performance when using time-ordered versus text-ordered training.
+  • Appendix E: Sensitivity analysis of masking history on Llama-3.3-70B and DeepSeek-R1 annotations.
+  • Appendix J: Additional insights into forecasting results.
+  • Appendix K: Analysis of open-source models to address data leakage concerns.
+  • Appendix L: Generalizability beyond PubMed case reports: comparison with other clinical sources.
+  • Appendix M: Broader societal impact and ethical considerations.
+Appendix A Validity of Temporal Annotations
+We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports.
+We refer to this as Textual Time Series for Sepsis (T2S2).
+We assessed the quality of T2S2 using three key metrics: event match rate, temporal concordance (c-index), and log-time discrepancy on two external validation sets: sepsis-10 (n = 10) and sepsis-100 (n = 90).(adopted from our prior work (Wang and Weiss 2025)).
+Text Event Match
+To align predicted clinical findings with reference annotations, we use a recursive best-match procedure adapted from (Wang and Weiss 2025).
+At each step, we select the reference-prediction pair with the minimal string-based distance; ties are broken by text order (i.e., the order of events in the annotation files).
+Matched items are removed and the process recurses on the remainder, preventing many-to-one mappings.
+We evaluated Levenshtein distance, cosine similarity of BERT (bert-base-uncased) embeddings, and sentence-embedding cosine similarity from PubMedBERT (S-PubMedBert-MS-MARCO).
+Prior validation showed PubMedBERT-based cosine similarity with a threshold of 0.1 performed best, and we used this metric for evaluation.
+Pseudocode is provided in Algorithm 1.
+Temporal Ordering
+Temporal ordering accuracy was quantified with the concordance index (c-index), the probability that a randomly chosen pair of events is correctly ordered.
+For reference times $\{t^{r}_{i}\}$ and predicted times $\{t^{p}_{i}\}$,
+$\text{c-index}=\frac{1}{N}\sum_{i<j}\mathds{1}\{(t^{r}_{i}-t^{r}_{j})(t^{p}_{i}-t^{p}_{j})>0\},$
+where $N$ is the number of comparable pairs (event indices $i,j$ with $t^{r}_{i}\neq t^{r}_{j}$ and $t^{p}_{i}\neq t^{p}_{j}$), and $\mathds{1}\{\cdot\}$ is the indicator function.
+Higher values indicate better agreement with the reference timeline.
+Algorithm 1 Recursive Best Match
+1:Two lists: ref (reference events) and pred (predicted events)
+2:List of best-matching event pairs
+3:function MatchEvents(ref, pred)
+4:  if ref is empty or pred is empty then
+5:   return $[]$
+6:  end if
+7:  $\text{min\_distance}\leftarrow\infty$
+8:  $\text{best\_pair}\leftarrow\text{None}$
+9:  for all $r\in\texttt{ref}$ do
+10:   for all $p\in\texttt{pred}$ do
+11:     $d\leftarrow\text{ComputeDistance}(r,p)$
+12:     if $d<\text{min\_distance}$ then
+13:      $\text{min\_distance}\leftarrow d$
+14:      $\text{best\_pair}\leftarrow(r,p)$
+15:     else if $d=\text{min\_distance}$ then
+16:      Get indices of $r$, $p$, and current best_pair
+17:      if index of $r<$ best $r$ index then
+18:        $\text{best\_pair}\leftarrow(r,p)$
+19:      else if index of $r=$ best $r$ index and index of $p<$ best $p$ index then
+20:        $\text{best\_pair}\leftarrow(r,p)$
+21:      end if
+22:     end if
+23:   end for
+24:  end for
+25:  Remove best_pair.r from ref
+26:  Remove best_pair.p from pred
+27:  $\text{result}\leftarrow[\text{best\_pair}]+\text{MatchEvents}(\texttt{ref},\texttt{pred})$
+28:  return result
+29:end function
+Time Discrepancy Analysis
+Temporal prediction accuracy was assessed via discrepancies between predicted and reference timestamps.
+We computed the log-time discrepancy as $\log(1+\Delta t)$, where $\Delta t$ is the absolute time difference (in hours), and summarized errors with the Area Under the Log-Time Cumulative Distribution Function (AULTC), which reflects how rapidly the error CDF rises—larger AULTC indicates better temporal alignment.
+sepsis-10 Model Event Match Rate Concordance AULTC DeepSeek-R1 0.79 0.876 0.772 Llama-3.3-70B 0.75 0.932 0.759 sepsis-100 Model Event Match Rate Concordance AULTC DeepSeek-R1 0.91 0.875 0.801 Llama-3.3-70B 0.78 0.894 0.791
+Table A1: Comparison of event match rate (Event), median concordance (c), median absolute error (MAE), and area under the log-time curve (AULTC) across DeepSeek-R1 and Llama-3.3-70B for sepsis-10, and sepsis-100 datasets.
+Formally, we compute the AULTC as the area under $F(x)$ from $x=0$ to $x=\log(1+S_{\text{max}})$, normalized by $\log(1+S_{\text{max}})$:
+AULTC $=\frac{1}{\log(1+S_{\max})}\sum_{i=1}^{k}\big(x_{(i)}-x_{(i-1)}\big)\frac{i}{k}$
+$\quad+\frac{\log(1+S_{\max})-x_{(k)}}{\log(1+S_{\max})},$
+where $x_{(0)}=0$.
+With this definition, AULTC = 1 indicates that discrepancies are zero (perfect recovery), resulting in maximum area $\log(1+S_{\text{max}})$, and AULTC = 0 indicates that all discrepancies exceed $S_{\text{max}}$, yielding zero area.
+We also stratified results by reference time intervals (hour, day, week, year) to examine accuracy as a function of time from case presentation.
+Annotation Quality Results
+On the sepsis-10 subset, DeepSeek-R1 achieved an event match rate of 0.79, c-index of 0.876, and AULTC of 0.772, while Llama-3.3-70B obtained 0.75, 0.932, and 0.759, respectively.
+Thus, Llama provided stronger pairwise ordering (higher c-index) but with fewer matched events and a slightly weaker log-time error profile (lower AULTC) than DeepSeek.
+On sepsis-100, DeepSeek-R1 reached 0.91 (event match rate), 0.875 (c-index), and 0.801 (AULTC), whereas Llama-3.3-70B reached 0.78, 0.894, and 0.791.
+Here, DeepSeek matched substantially more events and achieved better AULTC, while Llama again showed a small c-index advantage.
+Overall, DeepSeek tends to optimize event identification and log-time alignment, whereas Llama slightly favors pairwise temporal ordering.
+Overall, these metrics together indicate strong alignment of the generated textual time series with expert annotations, sufficient to support forecasting benchmarks.
+B  LLM Prompt to Generate Clinical Textual Time Series from PMOA Case Reports
+Prompt You are a physician.
+Extract the clinical events and the related time stamp from the case report.
+The admission event has timestamp 0.
+If the event is not available, we treat the event, e.g.
+current main clinical diagnosis or treatment with timestamp 0.
+The events that happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time.
+The timestamps are in hours.
+The unit will be omitted when output the result.
+If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation.
+We want to predict the future events given the events happened in history.
+For example, here is the case report.
+An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash.
+Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks.
+With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome.
+The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present.
+One day later the patient was discharged.
+Let’s find the locations of event in the case report, it shows that four weeks ago of fever and rash, four weeks ago, he was diagnosed with acne and receive treatment.
+So the event of fever and rash happened four weeks ago, 672 hours, it is before admitted to the hospital, so the time stamp is -672.
+diffuse erythematous or maculopapular eruption with pruritus was documented on the admission exam, so the timestamp is 0 hours, since it happens right at admission.
+DRESS syndrome has no specific time, but it should happen soon after admission to the hospital, so we use our clinical judgment to give the diagnosis of DRESS syndrome the timestamp 0.
+then the output should look like:
+18 years old | 0
+male | 0
+admitted to the hospital | 0
+fever | -72
+rash | -72
+acne | -672
+minocycline | -672
+increased WBC count | 0
+eosinophilia | 0
+systemic involvement | 0
+diffuse erythematous or maculopapular eruption | 0
+pruritus | 0
+DRESS syndrome | 0
+fever persisted | 0
+rash persisted | 0
+discharged | 24
+Separate conjunctive phrases into its component events and assign them the same timestamp (for example, separation of ‘fever and rash’ into 2 events: ‘fever’ and ‘rash’).
+If the event has duration, assign the event time as the start of the time interval.
+Attempt to use the text span without modifications except ‘history of’ where applicable.
+Include all patient events, even if they appear in the discussion; do not omit any events; include termination/discontinuation events; include the pertinent negative findings, like ‘no shortness of breath’ and ‘denies chest pain’.
+Show the events and timestamps in rows, each row has two columns: one column for the event, the other column for the timestamp.
+The time is a numeric value in hour unit.
+The two columns are separated by a pipe $\mid$ as a bar-separated file.
+Skip the title of the table.
+Reply with the table only.
+C  Prompting Strategies for Forecasting Tasks
+Zero-Shot
+Prompt You are an expert physician.
+Reply to the prompt with structured predictions in a k-item, bar-separated row.
+For example, if there are k=3 events (A, B, C) and only B occurs in the time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response.
+Reply with the answer only.
+Do NOT provide any other text.
+Do NOT explain.
+Zero-Shot (with degree of confidence)
+Prompt You are an expert physician.
+Reply to the prompt with structured predictions in a k-item, bar-separated row.
+Assess whether the event happens in the time window using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen.
+For example, if there are k=3 events (A, B, C) and only B occurs in the time window, then: $1|5|1$ would be the correct response.
+Reply with the answer only.
+Do NOT provide any other text.
+Do NOT explain.
+Few-Shot
+Prompt You are an expert physician.
+Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window.
+Example 1.
+If there are k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response.
+Example 2.
+Suppose the time series is: admitted | 0
+history of diabetes | 0
+elevated troponin | 5
+--- PREDICTION WINDOW ---
+cardiac catheterization | 6
+admission to icu | 24
+transfer to floor | 72
+discharge | 144
+office visit | 1440
+Then, if the prediction is at 5 hours, the time preceding “— PREDICTION WINDOW —”, the events admitted, history of diabetes, and elevated troponin have occurred.
+Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor, and discharge happen.
+So the correct response for the example is: 1 $\mid$ 1 $\mid$ 0 $\mid$ 0 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
+Reply with the answer only.
+Do NOT provide any other text.
+Do NOT explain.
+Few-Shot Ordinal
+Prompt You are an expert physician.
+Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window, using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen.
+Example 1.
+If there are k=3 events (A, B, C) and only B occurs in the time window, then: 1 $\mid$ 5 $\mid$ 1 would be the correct response.
+Example 2.
+Suppose the time series is: admitted | 0
+history of diabetes | 0
+elevated troponin | 5
+--- PREDICTION WINDOW ---
+cardiac catheterization | 6
+admission to icu | 24
+transfer to floor | 72
+discharge | 144
+office visit | 1440
+Then, if the prediction is at 5 hours, the time preceding “— PREDICTION WINDOW —”, the events admitted, history of diabetes, and elevated troponin have occurred.
+Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor and discharge happen.
+So the an excellent response for the example is: 5 $\mid$ 4 $\mid$ 1 $\mid$ 1 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
+Reply with the answer only.
+Do NOT provide any other text.
+Do NOT explain.
+Appendix D Textual Time Series Forecasting Framework of Encoder Models
 We develop a general framework for predictive modeling over textual time series of the encoder-only models, designed to accommodate the irregular, event-based structure common to clinical case reports.
@@
   • microsoft/deberta-v3-small (He et al., 2021)
-  • answerdotai/ModernBERT-base and answerdotai/ModernBERT-large
-  • thomas-sounack/BioClinical-ModernBERT-base and thomas-sounack/BioClinical-ModernBERT-large
+  • answerdotai/ModernBERT-base, answerdotai/ModernBERT-large
+  • thomas-sounack/BioClinical-ModernBERT-base, thomas-sounack/BioClinical-ModernBERT-large
 Each model is used with its corresponding tokenizer.
@@
 All training is conducted using PyTorch and HuggingFace Transformers.
-Appendix D Survival Modeling with Language Model Embeddings
+Appendix E Survival Modeling with Language Model Embeddings
+Model Hyperparameters RSF n_estimators $\in$ {20,50,100,200,500,1000} min_samples_split $\in$ {2,10,20,50,100} min_samples_leaf $\in$ {1,5,10,20} DeepSurv/DeepHit num_nodes $\in$ {32,64,128,256,512,768,1024,1200, 1536,1800,2048,3072,4096} dropout $\in$ {0.1,0.2,0.4,0.5}
+Table A2: Hyperparameter grid search ranges for our survival models.
 To evaluate the prognostic value of textual information encoded in large language models (LLMs), we adopt a two-stage framework: (1) extraction of fixed-dimensional sequence embeddings from various pre-trained LLMs, and (2) downstream survival modeling using these embeddings as covariates.
@@
 This framework allows for a systematic evaluation of how different pre-trained LLMs, and their associated embedding strategies, contribute to downstream survival prediction tasks.
-Table A4: Hyperparameter search ranges for survival models.
-Model Hyperparameters RSF n_estimators $\in$ {20,50,100,200,500,1000} min_samples_split $\in$ {2,10,20,50,100} min_samples_leaf $\in$ {1,5,10,20} DeepSurv/DeepHit num_nodes $\in$ {32,64,128,256,512,768,1024,1200, 1536,1800,2048,3072,4096} dropout $\in$ {0.1,0.2,0.4,0.5}
-Appendix E Computing Resources
-All experiments were conducted on a high-performance computing cluster running AlmaLinux.
-For the largest language models (DSL33 and L33), training required three NVIDIA A100 GPUs (80 GB each; total 240 GB GPU memory), while inference was performed on two A100 GPUs.
-The encoder models were trained and evaluated on a single NVIDIA A100 GPU (80 GB).
-After getting the embedding vectors from LLMs, survival models were trained and evaluated on a workstation equipped with an NVIDIA RTX 4090 GPU.
-System memory usage was 32 GB RAM.
-Software dependencies, including specific versions of Python libraries and frameworks, are documented in environment_tts_forecasting.yml file of our anonymized repository in the Supplementary Material.
-F  LLM prompt to generate clinical textual time-series from PMOA case reports
-Prompt You are a physician.
-Extract the clinical events and the related time stamp from the case report.
-The admission event has timestamp 0.
-If the event is not available, we treat the event, e.g.
-current main clinical diagnosis or treatment with timestamp 0.
-The events that happened before event with 0 timestamp have negative time, the ones after the event with 0 timestamp have positive time.
-The timestamp are in hours.
-The unit will be omitted when output the result.
-If there is no temporal information of the event, please use your knowledge and events with temporal expression before and after the events to provide an approximation.
-We want to predict the future events given the events happened in history.
-For example, here is the case report.
-An 18-year-old male was admitted to the hospital with a 3-day history of fever and rash.
-Four weeks ago, he was diagnosed with acne and received the treatment with minocycline, 100 mg daily, for 3 weeks.
-With increased WBC count, eosinophilia, and systemic involvement, this patient was diagnosed with DRESS syndrome.
-The fever and rash persisted through admission, and diffuse erythematous or maculopapular eruption with pruritus was present.
-One day later the patient was discharged.
-Let’s find the locations of event in the case report, it shows that four weeks ago of fever and rash, four weeks ago, he was diagnosed with acne and receive treatment.
-So the event of fever and rash happened four weeks ago, 672 hours, it is before admitted to the hospital, so the time stamp is -672.
-diffuse erythematous or maculopapular eruption with pruritus was documented on the admission exam, so the timestamp is 0 hours, since it happens right at admission.
-DRESS syndrome has no specific time, but it should happen soon after admission to the hospital, so we use our clinical judgment to give the diagnosis of DRESS syndrome the timestamp 0.
-then the output should look like:
-18 years old | 0
-male | 0
-admitted to the hospital | 0
-fever | -72
-rash | -72
-acne | -672
-minocycline | -672
-increased WBC count | 0
-eosinophilia | 0
-systemic involvement | 0
-diffuse erythematous or maculopapular eruption | 0
-pruritis | 0
-DRESS syndrome | 0
-fever persisted | 0
-rash persisted | 0
-discharged | 24
-Separate conjunctive phrases into its component events and assign them the same timestamp (for example, separation of ‘fever and rash’ into 2 events: ‘fever’ and ‘rash’).
-If the event has duration, assign the event time as the start of the time interval.
-Attempt to use the text span without modifications except ‘history of’ where applicable.
-Include all patient events, even if they appear in the discussion; do not omit any events; include termination/discontinuation events; include the pertinent negative findings, like ‘no shortness of breath’ and ‘denies chest pain’.
-Show the events and timestamps in rows, each row has two columns: one column for the event, the other column for the timestamp.
-The time is a numeric value in hour unit.
-The two columns are separated by a pipe $\mid$ as a bar-separated file.
-Skip the title of the table.
-Reply with the table only.
-G  Prompting strategies for forecasting tasks
-F.1  Zero-shot
-Prompt You are an expert physician.
-Reply to the prompt with structured predictions in a k-item, bar-separated row.
-For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response.
-Reply with the answer only.
-Do NOT provide any other text.
-Do NOT explain.
-F.2  Zero-shot (with degree of confidence)
-Prompt You are an expert physician.
-Reply to the prompt with structured predictions in a k-item, bar-separated row.
-Assess whether the event happens in the time window using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen.
-For example, if there k=3 events (A, B, C) and only B occurs in the time window, then: $1|5|1$ would be the correct response.
-Reply with the answer only.
-Do NOT provide any other text.
-Do NOT explain.
-F.3  Few-shot
-Prompt You are an expert physician.
-Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window.
-Example 1.
-If there k=3 events (A, B, C) and only B occurs in the requested time window, then: 0 $\mid$ 1 $\mid$ 0 would be the correct response.
-Example 2.
-Suppose the time series is: admitted | 0
-history of diabetes | 0
-elevated troponin | 5
---- PREDICTION WINDOW ---
-cardiac catheterization | 6
-admission to icu | 24
-transfer to floor | 72
-discharge | 144
-office visit | 1440
-Then, if the prediction is at 5 hours, the time preceding “— PREDICTION WINDOW —”, the events admitted, history of diabetes, and elevated troponin have occurred.
-Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor, and discharge happen.
-So the correct response for the example is: 1 $\mid$ 1 $\mid$ 0 $\mid$ 0 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
-Reply with the answer only.
-Do NOT provide any other text.
-Do NOT explain.
-F.4  Few-shot ordinal
-Prompt You are an expert physician.
-Reply to the prompt with predictions in a k-item, bar-separated row to answer if the k events happen within a given forecast time window, using a scale from 1 to 5, where 5 means it definitely happened and 1 means it did not happen.
-Example 1.
-If there k=3 events (A, B, C) and only B occurs in the time window, then: 1 $\mid$ 5 $\mid$ 1 would be the correct response.
-Example 2.
-Suppose the time series is: admitted | 0
-history of diabetes | 0
-elevated troponin | 5
---- PREDICTION WINDOW ---
-cardiac catheterization | 6
-admission to icu | 24
-transfer to floor | 72
-discharge | 144
-office visit | 1440
-Then, if the prediction is at 5 hours, the time preceding “— PREDICTION WINDOW —”, the events admitted, history of diabetes, and elevated troponin have occurred.
-Supposing k=4 and a forecast window of 24 hours, then the task it to determine which of cardiac catheterization, admission to icu, transfer to floor and discharge happen.
-So the an excellent response for the example is: 5 $\mid$ 4 $\mid$ 1 $\mid$ 1 because cardiac catherization (6) and admission to icu (24) happen before 5 + 24 = 29.
-Reply with the answer only.
-Do NOT provide any other text.
-Do NOT explain.
-Appendix H Open-source models for avoiding potential data leakage
-We recognize that data leakage is a critical concern in evaluating model performance, particularly when using open-access text corpora.
-Below, we outline key factors that mitigate this concern in our study:
-1.
-Temporal annotations—not raw text—drive forecasting tasks.
-The core contribution of this work lies in the temporal structuring of clinical narratives.
-While large language models (LLMs) such as DeepSeek and Llama-3.3 may have encountered PubMed Open Access (PMOA) case reports during pretraining, they were never exposed to the extracted timestamps or temporally ordered and future-masked (event, time) tuples used in our evaluations.
-These derived temporal annotations are essential to all forecasting tasks in our study—including F1-based next-event prediction, temporal ordering, and survival analysis—and constitute novel structured data not present in the original text.
-2.
-Temporal reordering fundamentally alters the learning problem.
-Even if LLMs were pretrained on PMOA narratives, they would have encountered events in their original, text-based order.
-In contrast, our framework chronologically reorders events based on extracted timestamps, creating a new temporal structure.
-This transformation introduces a distinct modeling challenge, as it breaks the sequential narrative flow seen during pretraining.
-As shown in Section 4.4.1, training on time-ordered versus text-ordered annotations leads to measurable differences in performance, underscoring the impact of this restructuring.
-3.
-External validation reveals limited effects of memorization.
-Our experimental design separates internal (T2S2 test set) and external validation (sepsis-10 and sepsis-100 cohorts), using entirely held-out case reports.
-The performance drop observed on external datasets (Tables 1–2) suggests that any potential memorization of PMOA case reports during pretraining yields minimal benefit and reinforces the robustness of our empirical findings.
-4.
-Broader context: standard practices and safeguards.
-Training data overlap is a general challenge in modern machine learning, especially for LLMs trained on web-scale corpora.
-Our work adheres to community standards while incorporating additional safeguards—such as explicit temporal annotation, event reordering, and external dataset validation—that reduce the risk of leakage and differentiate our tasks from standard text modeling.
-To address concerns regarding potential pretraining overlap, we evaluated multiple models with documented training corpora.
-Both OLMO-32B-Instruct (Allen Institute for AI) and RedPajama-INCITE-7B-Instruct represent fully open-source models with transparent training data.
-Notably, RedPajama explicitly excludes PubMed from its training corpus, making it an ideal benchmark for assessing causal leakage.
-The comparable performance of these models to those potentially exposed to PMOA (see point 5 above) provides strong evidence that our temporal forecasting tasks are meaningfully distinct from standard language modeling on clinical narratives.
-5.
-Empirical evidence from PubMed-excluded models.
-Our results provide direct empirical evidence against meaningful causal leakage.
-RedPajama-INCITE-7B-Instruct, which explicitly excludes PubMed from its training corpus, achieves comparable performance to other decoder models (0.352 F1 at 24h, 0.618 concordance) that potentially had access to PMOA during pretraining.
-This similarity in performance across models with and without PubMed exposure demonstrates that our T2S2 prediction task—with its temporal restructuring, event masking, and partial history inputs—is sufficiently distinct from raw case report text that prior exposure provides no meaningful advantage.
-Furthermore, OLMO-32B-Instruct shows competitive performance (0.411 F1 at 24h), providing additional validation from a model with fully documented training data.
-Appendix I Broader Impact
+F  Forecasting Performance on DeepSeek-R1 and Llama-3.3 Textual Time Series Annotations
+Tables A3 and A4 present the forecasting performance of our models on textual time series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively.
+We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
+T2S2 (DeepSeek-R1) sepsis-10 sepsis-100
+F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
+LLM-MLP head
+DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
+L3.3 70B 0.000 0.295 0.764 0.631 0.312 0.569 0.404 0.597
+DS-L3.1 8B 0.000 0.492 0.650 0.630 0.381 0.566 0.601 0.600
+L3.1 8B 0.000 0.355 0.650 0.630 0.114 0.575 0.442 0.604
+OLMO 32B Instruct 0.000 0.411 0.688 0.621 0.273 0.561 0.432 0.593
+RedPajama-INCITE 7B Instruct 0.000 0.352 0.651 0.618 0.282 0.572 0.471 0.594
+MediPhi-PubMed 0.000 0.268 0.751 0.626 0.169 0.561 0.210 0.597
+Prompting
+L3.3 70B 0-shot 0.085 0.368 0.560 0.236 0.452
+L3.3 70B 0-shot ordinal 0.082 0.352 0.646 0.289 0.483
+L3.3 70B few-shot 0.095 0.380 0.652 0.360 0.452
+L3.3 70B few-shot ordinal 0.063 0.356 0.682 0.313 0.480
+OLMO 32B 0-shot 0.018 0.333 0.652 0.322 0.430
+OLMO 32B 0-shot ordinal 0.006 0.323 0.685 0.325 0.443
+OLMO 32B few-shot 0.014 0.316 0.630 0.219 0.423
+OLMO 32B few-shot ordinal 0.013 0.315 0.651 0.247 0.422
+Encoder-MLP head
+BERT 0.264 0.550 0.871 0.501 0.441 0.498 0.614 0.501
+RoBERTa 0.250 0.562 0.866 0.500 0.405 0.500 0.614 0.500
+DeBERTa-small 0.262 0.570 0.871 0.601 0.382 0.498 0.617 0.556
+ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
+ModernBERT-large 0.277 0.568 0.874 0.598 0.391 0.539 0.602 0.553
+Bioclinical ModernBERT-base 0.264 0.559 0.879 0.677 0.395 0.598 0.635 0.610
+Bioclinical ModernBERT-large 0.313 0.562 0.868 0.671 0.399 0.491 0.618 0.607
+Encoder-masking-fine-tuned
+BERT 0.166 0.432 0.690 0.648 0.323 0.554 0.482 0.600
+RoBERTa 0.169 0.416 0.688 0.646 0.313 0.539 0.457 0.609
+DeBERTa-small 0.175 0.435 0.695 0.671 0.325 0.579 0.492 0.617
+ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
+ModernBERT-large 0.177 0.460 0.683 0.663 0.342 0.615 0.496 0.632
+Bioclinical ModernBERT-base 0.166 0.450 0.692 0.676 0.336 0.604 0.499 0.655
+Bioclinical ModernBERT-large 0.165 0.422 0.696 0.652 0.325 0.626 0.425 0.603
+Encoder-masking-zeroshot
+BERT 0.000 0.000 0.000 0.495 0.000 0.513 0.000 0.494
+RoBERTa 0.049 0.207 0.375 0.500 0.137 0.500 0.318 0.500
+DeBERTa-small 0.170 0.460 0.323 0.497 0.268 0.512 0.376 0.509
+ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
+ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
+Bioclinical ModernBERT-base 0.246 0.458 0.614 0.498 0.301 0.509 0.457 0.498
+Bioclinical ModernBERT-large 0.244 0.446 0.596 0.479 0.364 0.502 0.416 0.497
+Table A3: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
+Models italicized represent the open-source and medically fine-tuned models.
+All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1.
+Bold indicates best in that particular category amongst all models.
+T2S2 (Llama-3.3) sepsis-10 sepsis-100
+F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
+LLM-MLP head
+DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
+L3.3 70B 0.265 0.466 0.807 0.619 0.160 0.567 0.466 0.597
+DS-L3.1 8B 0.264 0.584 0.837 0.624 0.258 0.576 0.605 0.598
+L3.1 8B 0.338 0.565 0.823 0.619 0.368 0.569 0.536 0.596
+OLMO 32B Instruct 0.190 0.315 0.764 0.604 0.238 0.551 0.315 0.596
+RedPajama-INCITE 7B Instruct 0.127 0.459 0.705 0.620 0.257 0.586 0.491 0.594
+MediPhi-PubMed 0.000 0.507 0.801 0.623 0.277 0.564 0.534 0.600
+Prompting
+L3.3 70B 0-shot 0.085 0.443 0.616 0.236 0.452
+L3.3 70B 0-shot ordinal 0.063 0.454 0.707 0.289 0.483
+L3.3 70B few-shot 0.099 0.457 0.703 0.360 0.452
+L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
+OLMO 32B 0-shot 0.036 0.417 0.708 0.254 0.473
+OLMO 32B 0-shot ordinal 0.034 0.457 0.741 0.167 0.427
+OLMO 32B few-shot 0.013 0.428 0.631 0.267 0.424
+OLMO 32B few-shot ordinal 0.046 0.413 0.647 0.300 0.424
+Encoder-MLP head
+BERT 0.224 0.622 0.899 0.500 0.400 0.500 0.612 0.500
+RoBERTa 0.180 0.608 0.900 0.500 0.364 0.500 0.597 0.500
+DeBERTa-small 0.258 0.641 0.898 0.603 0.427 0.585 0.616 0.592
+ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
+ModernBERT-large 0.312 0.622 0.897 0.500 0.404 0.500 0.623 0.500
+Bioclinical ModernBERT-base 0.290 0.653 0.902 0.650 0.449 0.618 0.662 0.627
+Bioclinical ModernBERT-large 0.258 0.648 0.901 0.626 0.409 0.602 0.634 0.593
+Encoder-masking-fine-tuned
+BERT 0.131 0.471 0.741 0.622 0.320 0.580 0.467 0.622
+RoBERTa 0.129 0.484 0.659 0.625 0.329 0.587 0.472 0.614
+DeBERTa-small 0.135 0.468 0.693 0.648 0.318 0.602 0.462 0.623
+ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
+ModernBERT-large 0.124 0.470 0.712 0.646 0.318 0.586 0.468 0.631
+Bioclinical ModernBERT-base 0.129 0.481 0.733 0.653 0.331 0.615 0.475 0.658
+Bioclinical ModernBERT-large 0.131 0.480 0.840 0.658 0.318 0.587 0.478 0.641
+Encoder-masking-zeroshot
+BERT 0.000 0.000 0.000 0.504 0.000 0.515 0.000 0.506
+RoBERTa 0.074 0.279 0.445 0.500 0.083 0.500 0.319 0.500
+DeBERTa-small 0.008 0.439 0.839 0.562 0.296 0.518 0.515 0.501
+ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
+ModernBERT-large 0.109 0.413 0.622 0.499 0.326 0.521 0.339 0.518
+Bioclinical ModernBERT-base 0.197 0.460 0.533 0.499 0.299 0.520 0.465 0.499
+Bioclinical ModernBERT-large 0.194 0.451 0.525 0.496 0.342 0.489 0.433 0.501
+Table A4: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events.
+Models italicized represent the open-source and medicallty fine-tuned models.
+All models are trained/fine-tuned on time-ordered annotations from Llama-3.3-70B.
+G  Survival Performance on DeepSeek-R1 Annotations for T2S2, sepsis-10, and sepsis-100
+In this section, we report the full concordance analyses for the sepsis-10 and sepsis-100 annotation sets in addition to T2S2 test set, presented in Table A5.
+These results are the extended version of the T2S2-only table (Table 2) included in the main text.
+Model T2S2 sepsis-10 sepsis-100
+Time Threshold 0h 24h 168h 0h 24h 168h 0h 24h 168h
+bert-base-uncased 0.60 0.61 0.53 1.00 1.00 1.00 0.58 0.62 0.58
+roberta-base 0.55 0.60 0.60 0.75 1.00 1.00 0.55 0.58 0.57
+deberta-v3-small 0.56 0.57 0.57 0.75 0.00 0.75 0.52 0.49 0.51
+ModernBERT-base 0.52 0.58 0.58 1.00 1.00 1.00 0.54 0.51 0.51
+ModernBERT-large 0.53 0.59 0.58 0.50 1.00 1.00 0.59 0.55 0.49
+Bioclinical ModernBERT-base 0.57 0.53 0.53 1.00 0.75 0.75 0.63 0.59 0.56
+Bioclinical ModernBERT-large 0.54 0.55 0.54 0.75 0.75 0.75 0.61 0.57 0.49
+DeepSeek-R1-Llama-70B 0.64 0.63 0.59 0.50 1.00 1.00 0.61 0.53 0.55
+Llama-3.3-70B-Instruct 0.62 0.63 0.58 0.50 1.00 0.75 0.64 0.57 0.53
+DeepSeek-R1-Llama-8B 0.60 0.58 0.58 1.00 1.00 1.00 0.53 0.66 0.65
+Llama-3.1-8B-Instruct 0.62 0.61 0.61 1.00 1.00 1.00 0.60 0.59 0.61
+OLMO 32B Instruct 0.60 0.62 0.59 0.75 0.75 0.75 0.66 0.58 0.56
+RedPajama-INCITE 7B Instruct 0.56 0.60 0.61 0.50 1.00 1.00 0.61 0.53 0.76
+MediPhi-PubMed 0.54 0.62 0.55 1.00 1.00 1.00 0.62 0.58 0.61
+Table A5: Survival analyses evaluated using time-dependent concordance index for each model across DeepSeek-R1 test set annotations (T2S2, sepsis-10, sepsis-100), evaluated at three observation time thresholds (0h, 24h, and 168h).
+H  Sensitivity Analysis of Time-Ordered versus Text-Ordered Training
+Table A6 shows time-ordering generally improves concordance, particularly for DeepSeek-R1 annotations, while F1 scores show mixed results.
+Text-ordering sometimes yields better external dataset performance with Llama-3.3-70B, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order.
+These results highlight a trade-off between temporal alignment and narrative coherence for different tasks.
+F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
+T2S2 (DeepSeek-R1)
+Text-ordered 0.290 0.569 0.876 0.659 0.361 0.598 0.615 0.609
+Time-ordered 0.306 0.576 0.877 0.672 0.428 0.576 0.607 0.613
+T2S2 (Llama-3.3-70B)
+Text-ordered 0.311 0.648 0.898 0.586 0.423 0.526 0.662 0.541
+Time-ordered 0.257 0.645 0.903 0.632 0.382 0.595 0.616 0.612
+Table A6: Comparison of text-ordered and time-ordered training for DeepSeek-R1 and Llama 3.3-70B annotations on F1 scores and c-index values.
+I  Masking History: Llama-3.3-70B Textual Time Series Annotations
+Tables A7 and A8 report the effect of timestep dropout on classification and ranking performance.
+We evaluate performance across three datasets (T2S2-L3.3, sepsis-10, and sepsis-100) using F1 scores for prediction of next $k$ events happening within the next day and concordance index (c-index) for correctly ordering the next $k$ events.
+As the proportion of masked history increases, we observe a general decline in F1 performance, particularly at higher masking rates (e.g., 90%), suggesting that access to historical context is important for accurate forecasting.
+Interestingly, the concordance index remains relatively stable, indicating the model’s robustness in ranking the next $k$ events with partial input histories.
+More specifically, we can see that F1 scores decrease monotonically with increasing TDR, with relatively mild degradation up to 30–60% dropout and sharper declines thereafter.
+For example, in the DeepSeek-R1 annotations (Table A7), T2S2/sepsis-10/sepsis-100 drop from 0.576/0.428/0.607 at TDR = 0% to 0.423/0.200/0.510 at TDR = 90%, with sepsis-10 showing the largest relative decline.
+In contrast, concordance (c-index) remains stable across dropout levels (0.661–0.678), indicating that ranking-based evaluations are less sensitive to history truncation.
+These trends are consistent for Llama-3.3 annotations (Table A8), suggesting that ranking objectives capture temporal relationships that are more robust to missing historical information than binary classification accuracy of predicting whether an event happens within a time window.
+Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
+0% 0.576 0.672 0.428 0.576 0.607 0.613
+15% 0.572 0.667 0.415 0.564 0.604 0.616
+30% 0.548 0.672 0.357 0.542 0.591 0.619
+60% 0.542 0.678 0.310 0.578 0.588 0.616
+90% 0.423 0.661 0.200 0.561 0.510 0.614
+Table A7: Effect of Randomly Masking History (DeepSeek-R1 annotations) on F1 and Concordance.
+Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively.
+c10/100: c-index for sepsis-10 and sepsis-100 respectively.
+Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
+0% 0.645 0.632 0.431 0.595 0.626 0.612
+15% 0.628 0.635 0.422 0.592 0.624 0.610
+30% 0.634 0.634 0.423 0.586 0.621 0.598
+60% 0.604 0.634 0.391 0.577 0.605 0.613
+90% 0.531 0.622 0.328 0.546 0.534 0.607
+Table A8: Effect of Randomly Masking History (Llama-3.3-70B annotations) on F1 and Concordance.
+Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively.
+c10/100: c-index for sepsis-10 and sepsis-100 respectively.
+Appendix J Additional Insights into Forecasting Results
+In this section, we provide some additional insights into the forecasting results. Specifically, we summarize two observations that clarify model behavior and apparent anomalies in the results.
+Why decoders fare better on survival but not short-horizon forecasting
+Our findings align with adjacent literature on representation usage over different time horizons.
+Encoder models, typically pooled via a [CLS] token, excel on short-horizon tasks because bidirectional attention aggregates local context effectively (Devlin et al.
+2019; Bai, Kolter, and Koltun 2018).
+In contrast, decoder models trained autoregressively pass information forward through final token states, which is advantageous for long-range aggregation required by survival outcomes (Dai et al.
+2019; Zhou et al.
+2021).
+In our setting, forecasting aligns with near-term prediction, whereas survival analysis emphasizes longer-range temporal dependencies; this mismatch naturally yields the observed performance split.
+While not a one-to-one architectural proof, the pattern is consistent with prior evidence on horizon-specific strengths (Wang and Weiss 2025).
+Interpreting very low $F_{1}$ scores and external validation drops
+Why some $F_{1}$ scores are near zero.
+The very low F1 scores observed for certain models are not random failures but arise from two interacting factors.
+First, F1 is highly sensitive to threshold choice when the positive class is rare; small shifts in threshold calibration can yield 0.000 even when the model produces meaningful probabilities.
+Second, **architectural limitations** amplify this effect: **autoregressive decoders** tend to reproduce narrative rather than temporal order, and **instruction-tuned models** often default to generic or high-probability completions rather than selecting the correct event token—a behavior documented in prior analyses of instruction-following LLMs ((Liu et al.
+2024b; Wei et al.
+2021)).
+These failures do not undermine model comparisons; rather, they highlight systematic architectural differences.
+**Encoder-based models**, trained with a **bidirectional masking objective** ((Devlin et al.
+2019)), are better aligned with time-structured prediction tasks, as their contextual representation naturally integrates information from both past and future tokens.
+In contrast, **autoregressive decoders**, optimized for left-to-right generation ((Brown et al.
+2020)), are disadvantaged for temporal reasoning without explicit sequence restructuring or supervision.
+The contrast underscores that forecasting from clinical narratives is fundamentally different from standard text generation.
+Why external validation is lower.
+Performance drops on external subsets (e.g., sepsis-10 vs.
+sepsis-100) reflect shifts in writing style and case mix, not random overfitting.
+Crucially, the relative ranking of models remains stable across datasets, so comparative conclusions hold even when absolute scores vary.
+Appendix K Open-Source Models for Avoiding Potential Data Leakage
+We evaluated potential data leakage arising from large language models (LLMs) pretrained on web-scale corpora that may include PubMed Open Access (PMOA) case reports.
+Our safeguards target two avenues of leakage: (i) content overlap with PMOA text, and (ii) task leakage from using representations too close to raw prose.
+Below we summarize the controls embedded in our methods and the additional empirical evidence gathered for this concern.
+Task Design Minimizes Leakage
+Temporal annotations—not raw text—drive all evaluations.
+We transform narratives into chronological lists of (event, time) tuples and operate on these derived annotations (future-masked where applicable).
+Models are never scored on lexical similarity to the raw prose.
+Temporal reordering changes the learning problem.
+Events are reordered by extracted timestamps, breaking the original textual sequence seen during pretraining.
+This creates a distinct modeling challenge: correct temporal ordering and causal dependencies must be inferred from structured tuples rather than reproduced from prose.
+As shown in the main text, training/evaluating on time-ordered versus text-ordered sequences yields measurably different performance.
+Temporal metrics reduce any advantage from memorization.
+Evaluation emphasizes temporal reasoning—forecasting F1 at fixed horizons, pairwise concordance (c-index), and AULTC—rather than surface-form similarity.
+Memorizing text order can even hurt time-alignment metrics because clinical narratives often interleave retrospective commentary where textual order $\neq$ chronological order.
+Open-Source, PubMed-Free Baselines
+Decoder baseline.
+RedPajama-INCITE-7B-Instruct is pretrained on English CommonCrawl (67%), C4 (15%), GitHub (4.5%), Wikipedia (4.5%), Books (4.5%), arXiv (2.5%), and StackExchange (2%), explicitly excluding PubMed.
+In our experiments it performs comparably to decoders that may have seen PMOA during pretraining.
+Encoder baselines.
+We include BERT (BooksCorpus + Wikipedia), RoBERTa (CommonCrawl News, OpenWebText, Stories, Wikipedia), and DeBERTa (RoBERTa-style corpora).
+None of these encoders include PubMed in their pretraining data.
+Post-Cutoff External Validation (Additional Evidence)
+To probe encoder leakage further, we evaluated BERT, RoBERTa, and DeBERTa on 115 case reports published after their pretraining cutoffs (post-2020).
+Results mirror submission trends without degradation, which is inconsistent with a memorization hypothesis.
+Model 0h 24h 168h
+BERT 0.63 0.62 0.51
+RoBERTa 0.54 0.61 0.65
+DeBERTa 0.58 0.56 0.53
+Table A9: Post-cutoff evaluation on 115 case reports (forecasting F1 at 0h/24h/168h).
+We also report competitive performance from OLMO-32B-Instruct (fully documented training data) and parity from the PubMed-excluded RedPajama-INCITE-7B-Instruct relative to decoders potentially exposed to PMOA, further indicating that prior PubMed exposure does not meaningfully advantage our temporally structured tasks.
+The key takeaways are
+  • Using derived, time-ordered (event, time) tuples and temporal metrics directly targets temporal reasoning and reduces susceptibility to content memorization.
+  • PubMed-free baselines and post-cutoff tests provide external checks
+  • Empirically, models without PubMed exposure achieve comparable performance, arguing against causal leakage into our forecasting, concordance, and AULTC evaluations.
+Appendix L Generalizability Beyond Case Reports: Comparison with other Clinical Sources
+Scope and Rationale for using Case Reports
+We acknowledge that published case reports differ from routine clinical documentation: they emphasize atypical presentations and are written for publication rather than point-of-care use, so they cannot substitute for validation on EHR notes.
+Our choice is methodological and pragmatic.
+Case reports are (i) publicly available, (ii) rich in temporally explicit clinical reasoning, and (iii) sufficiently structured to support extraction of textual time series.
+This makes them a practical, reproducible testbed for assessing whether forecasting models operate on temporally structured narratives without IRB barriers or restricted data access.
+We therefore use case reports to establish methodological feasibility and transparency, while recognizing that generalization to clinical operations must be demonstrated on EHR notes.
+Position Relative to Structured Clinical Scores
+Direct comparison to structured risk scores (e.g., SOFA, SAPS) is infeasible because required variables are absent from case reports.
+Our contribution is complementary to prior work that integrates text with structured data (Kline et al.
+2022; Wang and Weiss 2025): we show that free-text narratives, once transformed into timelines, carry prognostic signal usable for forecasting.
+Whereas Kline et al.
+(2022) demonstrated that text can augment structured models, our experiments isolate the incremental value of text when converted into time series, independent of structured covariates.
+In contexts where structured variables are available, these representations can be fused with established scores to assess additive value; here we focus on the text-only signal because the necessary structured inputs are not present in case reports.
+Appendix M Broader Societal Impact
 This work advances equitable, scalable, and transparent medical AI by introducing a framework for forecasting clinical risk in sepsis from narrative documentation.